### v1

# disposition

Abstract
It is well-known that scalarization techniques (e.g., in the sense of Gerstewitz; Kasimbeyli; Pascoletti and Serafini; Zaffaroni) are useful for generating (weakly, properly) efficient solutions of vector optimization problems. One recognized approach is the conic scalarization method in vector optimization in real normed spaces proposed by Kasimbeyli (2010, SIAM J Optim 20), which is based on augmented dual cones and Bishop-Phelps type (norm-linear) scalarizing functions. In this paper, we present new results on cone separation in real topological-linear spaces by using Bishop-Phelps type separating cones / separating seminorm-linear functions. Moreover, we study some extensions of known scalarization results in vector optimization (in the sense of Eichfelder; Gerstewitz; Jahn; Kasimbeyli; Pascoletti and Serafini). On this basis, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s conic scalarization method in real normed spaces. Within this framework, we derive new Bishop-Phelps type scalarization results for the concepts of weak efficiency and different types of proper efficiency.
Abstract
Keywords: Vector optimization; efficiency; weak efficiency; proper efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
Mathematics subject classifications (MSC 2010): 90C48, 90C29, 06F20, 52A05
1 Introduction
The field of vector optimization in real topological-linear spaces is concerned with the minimization of a vector-valued function $f:X\to Y$ acting between real topological-linear spaces $X$ and $Y$ with respect to a convex cone $C\subseteq Y$ over a nonempty feasible set $\Omega\subseteq X$. A review of the existing literature reveals that many papers focus on the (finite dimensional) normed setting (i.e., $Y$ is a normed space), see e.g. [16], [8, 9], [12, 13], [17], [32], [45, 46], [53], [58], [57, Ch. 7 and Sec. 15.2], and references therein. In particular, the analysis includes real normed spaces $Y$ where the convex cone $C$ is of Bishop-Phelps type (named after Bishop and Phelps [2], see also [16], [15], [28, 29], [36, 37], [35],[41, 43], [45, 46]). The importance of the class of Bishop-Phelps cones can be emphasized by the fact that a nontrivial, closed, convex cone in a real normed space is representable as a Bishop-Phelps cone if and only if it is well-based (see Example 2.4 for details). We are primarily interested on the more general setting of a real (possibly infinite dimensional and not normed) topological-linear space $Y$ motivated by optimization problems under uncertainty and problems from mathematical finance. Especially, robust counterparts to optimization problems under uncertainty can be formulated as vector optimization problems with an infinite dimensional topological-linear image space (see Göpfert et al. [23, Ch. 6] and references therein). Furthermore, in arbitrage theory (see the class of financial market models in Marohn and Tammer [51, 52], and Riedel [56]), problems in the setting of linear spaces are considered. Some further works in vector optimization that focus on this setting are [5], [20, 21], [22], [23], [30], [34], [25], [42, 43], [57].
The main focus in our paper is on the computation of efficient / weakly efficient / properly efficient solutions of the given vector optimization problem via scalarization (based on Bishop-Phelps type cones), where the proper efficiency concepts considered in this paper cover certain known concepts of proper efficiency from the literature (e.g., in the sense of Benson [1]; Borwein [3, 4]; Henig [38]; Hurwicz [40]). It is well established that scalarization methods are useful for solving vector optimization problems. A notable feature of such methods is the ability to compute weakly efficient solutions by solving scalar optimization problems, which are often parameter-dependent, instead of the original vector optimization problem. Moreover, using scalarization, one can derive necessary and sufficient optimality conditions for the initial vector optimization problem, using generalized differentiation concepts for the new, scalarized objective function (see, e.g., [24], [32], [10], [33], [8], [34], [9]). The application of the scalarization methods for Henig dilating cones (enlarging convex cones) of the convex cone $C$ yields solutions that are properly efficient in the sense of Henig [38]. The relationships of the Henig proper efficiency concept to other known proper efficiency concepts have been well studied in the literature (see, e.g., [5, Sec. 2.4.3], [25, Ch. 3], [26], [49, Sec. 2.4] and references therein).
We will now recall some well-known scalarization methods in vector optimization. Of course, the linear scalarization method in vector optimization (respectively, the weighted sum method in multiobjective optimization) is the best known technique for generating weakly / properly efficient solutions of vector (multiobjective) optimization problems. It is known that this method only works well under cone-convexity type assumptions (see, e.g., [27]). For nonconvex vector optimization problems, there are some other scalarization methods frequently used in the literature that we recall below. In the framework of real (possibly infinite dimensional) normed spaces, the conic scalarization method by Kasimbeyli [45, 46] (see also [16]), the Drummond–Svaiter scalarization method [7] (see also [6]) and the Zaffaroni scalarization method [58] (see also [14, Sec. 5.1]) based on the Hiriart-Urruty function [39] (also known as oriented/signed distance function) are known to be useful, where the scalarizing function of the first two methods involves the underlying norm of the space $Y$. Considering the more general framework of real topological-linear spaces, only a few methods (besides the linear scalarization method) are known to be valid, such as the Gerstewitz scalarization method [20, 21] (see also [22], [49], [57], [23]) and the Pascoletti–Serafini scalarization method [53] (see also [12, 13], [57, Sec. 6.2.2]). Certain well-known scalarization methods in vector and multiobjective optimization (e.g. linear scalarization method / weighted sum method, $\varepsilon$-constraint method, hybrid method, Benson method, weighted Chebyshev norm method, and other methods; see [57, Sec. 7.3]) are special cases of these general scalarization methods. Some further comparison studies of scalarization methods in vector optimization are given in [6], [48], [12, Ch. 2], while generalizations of scalarization methods to vector and set optimization involving variable domination structures are also studied in the recent literature (see, e.g., [35], [15], [14], [43], [57, Ch. 10 and 11] and references therein).
As mentioned above, one recognized approach in vector optimization in a real normed space $(Y,||\cdot||)$ is the conic scalarization method in vector optimization proposed by Kasimbeyli [45], which is based on the following parameterized scalar optimization problem:
$\begin{cases}x^{*}(f(x)-a)+\alpha||f(x)-a||\to\min\\ x\in\Omega\end{cases}$ (1.1)
with a pair $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ (respectively, $(x^{*},\alpha)$ belongs to the so-called augmented dual cone of $C$; see Section 2.3 for details) and a given reference point $a\in Y$. Several relationships between (weakly, properly) efficient solutions of the original vector problem and solutions of scalar problems of type (1.1) are known (see [16], [17], [31], [45, 46]). In this paper, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s scalarization method in real normed spaces. By taking a look on (1.1) one can see that Kasimbeyli’s method is based on Bishop-Phelps type norm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of the norm |||||$\cdot$||| | ⋅ | | with a linear continuous functional $x^{*}$, see also Zaffaroni [59]), where the involving norm |||||$\cdot$||| | ⋅ | | is exactly the underlying norm in the real normed space $Y$. In our Bishop-Phelps type scalarization method, we consider Bishop-Phelps type seminorm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of a seminorm $\psi$ with a linear continuous functional $x^{*}$), where the seminorm involved (namely $\psi$) must not have anything to do with the underlying topology of the real (locally convex) topological-linear space (e.g., the seminorm must not be part of a family of seminorms that generates this topology).
The rest of the paper is organized as follows:
In Section 2, after presenting some basics in real topological-linear spaces including properties of seminorms, we study properties of cones and their bases, focusing in particular on augmented dual cones, Bishop-Phelps type cones and zero level sets of Bishop-Phelps type (seminorm-linear) functions. We then introduce the solution concepts of vector optimization necessary for our work.
Section 3 is devoted to the study of scalarization methods in vector optimization which are based on cone monotone and cone representing scalarizing functions. Beside classical scalarization results (in the sense of Eichfelder [12, 13]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]), we establish some new scalarization results (namely Propositions 3.11, 3.14 and 3.17) for vector optimization problems in real topological-linear spaces, which extend the classical ones.
In Section 4, we present the key tool to prove our main Bishop-Phelps type scalarization results in vector optimization, namely the technique of nonlinear separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. Besides presenting some classical nonlinear cone separation results, we also reflect on the separation conditions involved in both real topological-linear spaces and real normed spaces. Propositions 4.1 and 4.11 give some new insights into the relationships between the classical separation condition (the intersection of two cones is the singleton set $\{0\}$) and the separation conditions used in the nonlinear separation approach based on Bishop-Phelps type cones / functions discussed in [16], [28, 29], [45, 46]. We are able to deduce some new variants of nonlinear cone separation theorems involving the classical separation condition (see Corollaries 4.16 and 4.17).
The main Section 5 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces. The parameterized scalarization problems (in the sense of Kasimbeyli [45]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]) are formulated using Bishop-Phelps type cone representing (seminorm-linear) functions. In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector optimization problem (based on the convex cone $C$), an updated vector optimization problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions). Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorem 5.25 for proper efficiency w.r.t. a cone-valued map, Theorems 5.38, 5.40 for Henig proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
Finally, in Section 6 we present a conclusion and an outlook for further research.
2 Preliminaries
2.1 Basics in topological-linear spaces
Throughout the paper, we assume that $Y$ is a real topological-linear space. As usual, $Y^{*}$ denotes the topological dual space of $Y$. In certain situations we will also consider a more specific setting where $(Y,||\cdot||)$ is a real normed space. Consistently, the set of nonnegative real numbers is denoted by $\mathbb{R}_{+}$, while $\mathbb{P}:=\mathbb{R}_{++}$ denotes the set of positive real numbers. Given a set $\Omega\subseteq Y$, we denote by ${\mathrm{c}onv}\,\Omega$ the convex hull of $\Omega$. Moreover, we denote by ${\mathrm{c}l}\,\Omega$, ${\mathrm{b}d}\,\Omega$, and ${\mathrm{i}nt}\,\Omega$ the closure, the boundary, and the interior of $\Omega$ with respect to the underlying topology in $Y$. The algebraic interior (the core) of $\Omega$ is defined by $\operatorname{cor}\,\Omega:=\{x\in\Omega\mid\forall\,v\in Y\;\exists\,\varepsilon>0:\;x+[0,\varepsilon]\cdot v\subseteq\Omega\}.$ It is known that $\operatorname{int}\,\Omega\subseteq\operatorname{cor}\,\Omega\subseteq\Omega\subseteq\operatorname{cl}\,\Omega=(\operatorname{int}\,\Omega)\cup\operatorname{bd}\,\Omega,$ and if $\operatorname{int}\,\Omega\neq\emptyset$ (i.e., $\Omega$ is solid) and convex, then $\operatorname{cor}\,\Omega=\operatorname{int}\,\Omega$.
Given a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$, we define the $\psi$-unit ball by $\operatorname{\mathbb{B}}_{\psi}:=\{y\in Y\mid\psi(y)\leq 1\}$ and the $\psi$-unit sphere by $\operatorname{\mathbb{S}}_{\psi}:=\{y\in Y\mid\psi(y)=1\}.$
Remark 2.1
Suppose that $Y$ is a real topological-linear space and $\psi:Y\to\mathbb{R}$ is a seminorm. The following assertions hold:
  • $1^{\circ}$.
    $\psi$ is upper semicontinuous $\iff$ the set $\{y\in Y\mid\psi(y)<1\}$ is open $\iff$ $\psi$ is continuous at the origin $\iff$ $\psi$ is continuous.
  • $2^{\circ}$.
    $\psi$ is lower semicontinuous $\iff$ $\operatorname{\mathbb{B}}_{\psi}$ is closed.
Remark that ${\mathrm{k}er}\,\psi=\{0\}$ if and only if $\psi$ is a norm, and the singleton set $\{0\}$ is closed if and only if the topological-linear space $Y$ is Hausdorff. Moreover, it is known (see, e.g., Gelfand [18, 19], Eberlein [11]) that if $Y$ is a Banach space and $\psi:Y\to\mathbb{R}$ is a lower semicontinuous seminorm (alternatively, the seminorm $\psi$ has the Baire property), then $\psi$ is continuous.
2.2 Cones and their bases
A set $K\subseteq Y$ is called a cone if $0\in K=\operatorname{\mathbb{R}}_{+}\cdot K$. A cone $K\subseteq Y$ is said to be convex if $K+K=K$ (or equivalently, $K$ is a convex set); proper or nontrivial if $\{0\}\neq K\neq Y$; pointed if $\ell(K):=K\cap(-K)=\{0\}$. Given a convex cone $K\subseteq Y$ in the real topological-linear space $Y$, the set $\ell(K)$ is known as the lineality space of $K$. In this case, $K$ is a linear subspace of $Y$ if and only if $K=\ell(K)$, and moreover, $\operatorname{int}\,K=K+\operatorname{int}\,K$ and $K\setminus\ell(K)=K+(K\setminus\ell(K))$. If $K\subseteq Y$ is a cone, then $K\neq Y\iff 0\notin\operatorname{cor}\,K\iff 0\notin\operatorname{int}\,K,$ and if further $K$ is convex, then (see also [27, Lem. 2.10]) $K\neq Y\iff\ell(K)\cap\operatorname{int}\,K=\emptyset.$
The topological dual cone of a cone $K\subseteq Y$ is given by $K^{+}:=\{y^{*}\in Y^{*}\mid\forall\,k\in K:\;y^{*}(k)\geq 0\}.$ Furthermore, the subset $K^{\#}:=\{y^{*}\in Y^{*}\mid\forall\,k\in K\setminus\{0\}:\;y^{*}(k)>0\}$ of $K^{+}$ is of special interest. It is important to mention that both sets $K^{+}$ and $K^{\#}$ are convex for any (not necessarily convex) cone $K\subseteq Y$. Moreover, one has
$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\quad\text{and}\quad(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
but the inclusion $(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}$ can be strict (see Göpfert et al. [23, p. 55]).
In the following definition, we recall a general base concept for cones in real topological-linear spaces (cf. Göpfert et al. [23, Def. 2.1.42]).
Definition 2.2
Consider a nontrivial cone $K\subseteq Y$. A set $B\subseteq K$ is called a base for $K$, if $B$ is a nonempty set, and $K=\operatorname{\mathbb{R}}_{+}\cdot B$ with $0\notin\operatorname{cl}\,B$. Moreover, $K$ is said to be well-based if there exists a bounded, convex base of $K$.
Remark 2.3
Taking into account (2.1), for any nontrivial cone $K\subseteq Y$ in a real normed space $Y$, it is well-known (see [23, Prop. 2.2.23 and 2.2.32]) that
$\operatorname{conv}\,K\text{ has a convex base}\iff K^{\#}\neq\emptyset,$ (2.2)
and
$\operatorname{conv}\,K\text{ is well-based}\iff\operatorname{cl}(\operatorname{conv}\,K)\text{ is well-based}\iff\operatorname{int}\,K^{+}\neq\emptyset.$ (2.3)
Moreover, if $Y$ has finite dimension and $K$ is closed, then $K^{\#}=\operatorname{int}\,K^{+}$ (see [29, Th. 2.1(4), Rem. 2.6]), hence all the conditions involved in (2.2) and (2.3) are equivalent.
Example 2.4
Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$. For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by
$C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}.$
It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [36, 37], Phelps [55]). Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone. If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and
$C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset.$
Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$. It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 4.11). Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 2.3), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]).
Moreover, we recall another (algebraic) base concept for cones in linear spaces (cf. Göpfert et al. [23, Def. 2.1.14], Jahn [42, Def. 1.10 (d)]).
Definition 2.5
Consider a nontrivial cone $K\subseteq Y$. A set $B\subseteq K\setminus\{0\}$ is called an algebraic base for $K$, if $B$ is a nonempty set, and every $x\in K\setminus\{0\}$ has a unique representation of the form $x=\lambda b$ for some $\lambda>0$ and some $b\in B$.
Remark 2.6
If $B$ is base in the sense of Definition 2.2 or Definition 2.5 for the nontrivial cone $K\subseteq Y$, then $K=\operatorname{\mathbb{R}}_{+}\cdot B$ and $K\setminus\{0\}=\operatorname{\mathbb{P}}\cdot B$.
Our focus will be on a specific type of base in real topological-linear spaces, which we will recall in the following definition.
Definition 2.7
Consider a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ and a cone $K\subseteq Y$. Define a set
$B_{K}:=B_{K}(\psi):=\{x\in K\mid\psi(x)=1\}=K\cap\operatorname{\mathbb{S}}_{\psi}.$
If $K$ is nontrivial and $\psi$ is positive on $K\setminus\{0\}$, then $B_{K}$ is called a normlike-base for $K$. If $\psi$ is a norm $||\cdot||:Y\to\operatorname{\mathbb{R}}$, then $B_{K}(||\cdot||)$ is called a norm-base for $K$. Note that $B_{K}(\psi)$ is a base in the sense of Definition 2.5, and if $\psi$ is continuous and $K$ is closed, then it is a base also in the sense of Definition 2.2.
2.3 Augmented dual cones and Bishop-Phelps type cones
Assume that $Y$ is a topological-linear space, $K\subseteq Y$ is a nontrivial cone, and $\psi:Y\to\mathbb{R}$ is a seminorm. For any $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, we consider the Bishop-Phelps type (seminorm-linear) function $\varphi_{x^{*},\alpha}:Y\to\mathbb{R}$ defined by
$\varphi_{x^{*},\alpha}(y):=x^{*}(y)+\alpha\psi(y)\quad\text{for all }y\in Y,$
as used in [28]. Note that Bishop-Phelps type (norm-linear) functions in a normed setting were studied for instance by García-Castaño, Melguizo-Padial and Parzanese [16], Gasimov (Kasimbeyli) [17], Günther, Khazayel and Tammer [29], Guo and Hang [31], Ha [35], Jahn [43], Kasimbeyli [45, 46], and Zaffaroni [59].
Following the definitions of so-called augmented dual cones by Gasimov [17], Kasimbeyli [45] (see also [28] for some generalizations), we consider a (generalized) augmented dual cone by
$K^{a+}(\psi):=\{(x^{*},\alpha)\in K^{+}\times\mathbb{R}_{+}\mid\forall\,y\in K:\;x^{*}(y)-\alpha\psi(y)\geq 0\}$
as well the following sets
$K^{a\circ}(\psi)$ :=$:=\{(x^{*},\alpha)\in(K^{+}\setminus\{0\})\times\mathbb{R}_{+}\mid\forall\,y\in\operatorname{int}\,K:\;x^{*}(y)-\alpha\psi(y)>0\},$
$K^{a\#}(\psi)$ :=$:=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid\forall\,y\in K\setminus\{0\}:\;x^{*}(y)-\alpha\psi(y)>0\}.$
If no danger of confusion arises, we will use the notations $K^{a+}$, $K^{a\circ}$ and $K^{a\#}$.
Remark 2.8
Note that $(\operatorname{conv}\,K)^{a+}=K^{a+}$ and $(\operatorname{conv}\,K)^{a\#}=K^{a\#}$, and if $\psi$ is lower semicontinuous, then $(\operatorname{cl}(\operatorname{conv}\,K))^{a+}=K^{a+}$. It is a simple observation that for a normlike-base $B_{K}=B_{K}(\psi)$ of $K$ we have
$K^{a+}$ =$=\{(x^{*},\alpha)\in K^{+}\times\mathbb{R}_{+}\mid\forall\,y\in B_{K}:\;x^{*}(y)\geq\alpha\},$
$K^{a\#}$ =$=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid\forall\,y\in B_{K}:\;x^{*}(y)>\alpha\}.$
Furthermore, it is easy to observe that $K^{a\#}\subseteq K^{a\circ}\cup K^{a+}$, and if $K$ is solid and $\psi$ is lower semicontinuous, then $K^{a\circ}\subseteq K^{a+}$.
In the following definition, we introduce two classes of functions that satisfy cone (interior) representation properties (cf. Ha [35, Sec. 3.1], Jahn [43, Th. 3.1]).
Definition 2.9
The functional $\varphi:Y\to\mathbb{R}$ is
  • $-K$ representing if $-K=\{y\in Y\mid\varphi(y)\leq 0\}.$
  • strictly $-K$ representing if $\varphi$ is $-K$ representing and $-\operatorname{int}\,K=\{y\in Y\mid\varphi(y)<0\}.$
Remark 2.10
It is known that several scalarizing functions in vector optimization (e.g., in the sense of Gerstewitz [20, 21]; Hiriart-Urruty [39] / Zaffaroni [58, Prop. 3.2]) admit cone representation properties. Moreover, for certain topological-linear spaces $Y$, which are preordered by a (solid) convex cone $K\subseteq Y$, one can state $-K$ representing (strictly $-K$ representing) functionals (see, e.g., Günther et al. [30, Sec. 4.1], Ha [35], and Jahn [43]).
For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, the sublevel set of the Bishop-Phelps type function $\varphi_{x^{*},\alpha}$ at level zero is defined by
$S^{\leq}_{\psi}(x^{*},\alpha)$ :=$:=\{y\in Y\mid\varphi_{x^{*},\alpha}(y)=x^{*}(y)+\alpha\psi(y)\leq 0\},$
while a Bishop-Phelps type cone is defined by
$C_{\psi}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha\psi(y)\}.$
Remark that, in case $\psi=||\cdot||$ and $\alpha>0$, this reduces to the usual Bishop-Phelps cone (see Example 2.4). Moreover, the sets
$S^{<}_{\psi}(x^{*},\alpha)$ :=$:=\{y\in Y\mid\varphi_{x^{*},\alpha}(y)=x^{*}(y)+\alpha\psi(y)<0\},$
$C^{>}_{\psi}(x^{*},\alpha)$ :=$:=\{y\in Y\mid x^{*}(y)>\alpha\psi(y)\}$
are of interest. In our work we will focus in particular on Bishop-Phelps type functions, since they are (strictly) $-K$ representing functions if $K$ is a Bishop-Phelps type cone. Let us first note that the sets $K^{a+}$, $K^{a\circ}$ and $K^{a\#}$ can be written as
$K^{a+}$ =$=\{(x^{*},\alpha)\in K^{+}\times\mathbb{R}_{+}\mid K\subseteq C_{\psi}(x^{*},\alpha)\},$
$K^{a\circ}$ =$=\{(x^{*},\alpha)\in(K^{+}\setminus\{0\})\times\mathbb{R}_{+}\mid\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)\},$
$K^{a\#}$ =$=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)\}.$
In the following lemmata, we collect some important properties of Bishop-Phelps type cones and functions. For the first lemma, the proof is straightforward and we omit it.
Lemma 2.11
Consider some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. Then, we have:
  • $1^{\circ}$.
    Cone property: $0\in C_{\psi}(x^{*},\alpha),\operatorname{\mathbb{P}}\cdot C_{\psi}(x^{*},\alpha)=C_{\psi}(x^{*},\alpha)$ and $\operatorname{\mathbb{P}}\cdot C^{>}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)$.
  • $2^{\circ}$.
    Convexity: $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$ are convex.
  • $3^{\circ}$.
    Closedness: If $\varphi_{x^{*},\alpha}$ (respectively, $\psi$) is lower semicontinuous, then $C_{\psi}(x^{*},\alpha)$ is closed.
  • $4^{\circ}$.
    Pointedness: $\ell(C_{\psi}(x^{*},\alpha))=\{y\in Y\mid x^{*}(y)=0,\alpha\psi(y)=0\}$. If $\psi$ is a norm and $\alpha>0$, then $C_{\psi}(x^{*},\alpha)$ is pointed. If $(x^{*},\alpha)\in K^{a\#}$, then $K\cap\ell(C_{\psi}(x^{*},\alpha))=\{0\}$.
  • $5^{\circ}$.
    Nontrivialness: If $(x^{*},\alpha)\in K^{a\#}$, then $C_{\psi}(x^{*},\alpha)\neq\{0\}$. If $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)\neq Y$.
  • $6^{\circ}$.
    Solidness: Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ (e.g. if either $(x^{*},\alpha)\in K^{a\#}$ or $(x^{*},\alpha)\in K^{a\circ}$ and $K$ is solid). Then, $\operatorname{cor}\,C_{\psi}(x^{*},\alpha)\neq\emptyset,$ and if either $(x^{*},\alpha)\in K^{a\circ}$ and $K$ is solid or $\varphi_{x^{*},\alpha}$ (respectively, $\psi$) is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset.$
  • $7^{\circ}$.
    Lineality: If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ (e.g. if $(x^{*},\alpha)\in K^{a\#}$), and either $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)$ is not a linear subspace (or equivalently, $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$).
  • $8^{\circ}$.
    Dilating cone for $K$: If $(x^{*},\alpha)\in K^{a\#}$, then $C_{\psi}(x^{*},\alpha)$ is a dilating cone for $K$, i.e., $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$, and if further $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $K\setminus\{0\}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
Lemma 2.12
For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, the following assertions are valid:
  • $1^{\circ}$.
    $-C_{\psi}(x^{*},\alpha)=S^{\leq}_{\psi}(x^{*},\alpha)$ and $-C^{>}_{\psi}(x^{*},\alpha)=S^{<}_{\psi}(x^{*},\alpha)$.
  • $2^{\circ}$.
    $\varphi_{x^{*},\alpha}$ is sublinear (convex) and $-C_{\psi}(x^{*},\alpha)$ representing.
  • $3^{\circ}$.
    $C^{>}_{\psi}(x^{*},\alpha)\subseteq\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$, and if one has $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $C^{>}_{\psi}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$.
  • $4^{\circ}$.
    If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
Proof 2.13.
$1^{\circ}$ Both equalities follow from the definitions.
$2^{\circ}$ The sublinearity (convexity) of $\varphi_{x^{*},\alpha}$ follows from the sublinearity (convexity) of both $x^{*}$ and $\psi$. In view of $1^{\circ}$, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing.
$3^{\circ}$ This assertion is a direct consequence of Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
$4^{\circ}$ Of course, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing by $2^{\circ}$. Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. By $1^{\circ}$ and $3^{\circ}$ we get $-S^{<}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)=$. Since $C_{\psi}(x^{*},\alpha)$ is a convex set with $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ it follows that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=$. Thus, we conclude $-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=-$ which shows that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
Remark 2.14.
Note that the properties of $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$, and the relationships between $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$ follow immediately from the one related to $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$. Furthermore, note that the conditions $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ (as given in the assumptions of Lemma 2.12 ($4^{\circ}$)) do not imply each other, as can be seen below:
  • Let us first give an example with $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. According to [37, Ex. 2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has
    $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$
    and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
  • Next, we give an example with $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$. Recall that, for a set $\Omega\subseteq Y$ in a real normed space $(Y,||\cdot||)$, one denotes the weak interior and weak closure of $\Omega$ by ${\mathrm{i}nt}_{w}\,\Omega$ and ${\mathrm{c}l}_{w}\,\Omega$, respectively. We have the next result:
    Proposition 2.15.
    Consider a closed, pointed cone $C\subseteq Y$ in a real infinite dimensional normed space $(Y,||\cdot||$). Then, the weak interior of $C$ is empty, i.e., $\operatorname{int}_{w}\,C=\emptyset$.
    Proof 2.16.
    On the contrary, assume that $\operatorname{int}_{w}\,C\neq\emptyset$. Take some $c\in\operatorname{int}_{w}\,C$, i.e., there is a weakly open neighbourhood $U(c)\subseteq Y$ around $c$ which satisfies $U(c)\subseteq C$. Since $Y$ has infinite dimension, the set $U(c)$ contains a line $c+\mathbb{R}\cdot v$ for some $v\in Y\setminus\{0\}$. This implies, for any $n\in\mathbb{N}$,
    $\frac{c}{n}\pm v=\frac{1}{n}(c\pm nv)\in C.$
    Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction.
    As a desired second example, we endow a real infinite dimensional (normed) space $Y$ with the weak topology, we put $\psi(\cdot):=||\cdot||$ and we assume $||x^{*}||^{*}>\alpha>0$. In this case, it is known that $C^{>}_{||\cdot||}(x^{*},\alpha)=\operatorname{int}_{||\cdot||}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Since $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed cone in $(Y,||\cdot||)$ by Lemma 2.11 ($1^{\circ}$, $3^{\circ}$, $4^{\circ}$), we conclude that $\operatorname{int}_{w}\,C_{||\cdot||}(x^{*},\alpha)=\emptyset$ by Proposition 2.15.
Remark also that, if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\psi$ is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ by Lemma 2.11 ($6^{\circ}$). In the second example from above, the weak upper semicontinuity of |||||$\cdot$||| | ⋅ | | is not satisfied, since $Y$ has infinite dimension (cf. [29, Rem. 2.2]).
2.4 Solution concepts in vector optimization
Given two real topological-linear spaces $X$ and $Y$, a seminorm $\psi:Y\to\mathbb{R}$, a nonempty feasible set $\Omega\subseteq X$, and a vector-valued objective function $f:X\rightarrow Y$, we consider the following vector optimization problem:
$\begin{cases}f(x)\to\min\mbox{ with respect to }K\\ x\in\Omega,\end{cases}$ ($P$)
where we assume throughout the paper that $K\subseteq Y$ is a cone. It is well-known that $K$ induces on $Y$ a reflexive binary relation $\leq_{K}$ defined, for any two points $y,\overline{y}\in Y$, by $y\leq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K.$ Note that if $K$ is convex, then $\leq_{K}$ is a preorder; if $K$ is pointed and convex, then $\leq_{K}$ is a partial order. For notational convenience, we consider further the binary relations $\lneq_{K}$ and $<_{K}$ that are defined, for any two points $y,\overline{y}\in Y$, by $y\lneq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K\setminus\{0\},y<_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-\operatorname{int}\,K$, respectively. Two well-known types of solutions to the vector problem ($P$) can be defined according to the next two definitions (see, e.g., Boţ, Grad and Wanka [5], Göpfert et al. [23], Jahn [42], and Khan, Tammer and Zălinescu [49]).
Definition 2.17 (Efficiency).
A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$. The set of all Pareto efficient solutions of ($P$) is denoted by
${\mathrm{E}ff}(\Omega\mid f,K):=\{\overline{x}\in\Omega\mid\not\exists\,x\in\Omega:\;f(x)\lneq_{K}f(\overline{x})\}.$
Definition 2.18 (Weak efficiency).
A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$. The set of all weakly Pareto efficient solutions of ($P$) is denoted by
${\mathrm{W}Eff}(\Omega\mid f,K):=\{\overline{x}\in\Omega\mid\not\exists\,x\in\Omega:\;f(x)<_{K}f(\overline{x})\}.$
In addition to the efficiency concepts above, we also want to consider a proper efficiency with respect to a cone-valued map.
Definition 2.19 (Proper efficiency with respect to a cone-valued map).
Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$. A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and
$\mathbb{A}(\bar{x})\cap(-K)=\{0\}.$
The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
Remark 2.20.
In Definition 2.19 one can remove the assumption $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ if $\mathbb{A}(\bar{x})$ contains $f[\Omega]-f(\bar{x})$. Notice that ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)={\mathrm{E}ff}(\Omega\mid f,K)$ and ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,\{0\}\cup\operatorname{int}\,K)={\mathrm{W}Eff}(\Omega\mid f,K)$ with $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]-f(\cdot))$ or $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot)).$
Remark 2.21.
For the particular choice of the cone-valued map $\operatorname{\mathbb{A}}:\Omega\to 2^{Y}$, one can see that several well-known concepts of proper efficiency are special cases of Definition 2.19: Hurwicz proper efficiency ([40], 1958), Borwein proper efficiency ([3], 1977), Benson proper efficiency ([1], 1979), extended Borwein proper efficiency ([4], 1980). For more details see, e.g., [5, Sec. 2.4.3], [25, Ch. 3], [26], [49, Sec. 2.4] and references therein.
Beside the proper efficiency concept given in Definition 2.19, we like also to consider a proper efficiency concept in the sense of Henig ([38], 1982). To formulate this concept in the next definition, consider the following family of dilating cones:
$\mathcal{D}(K):=\{C\subseteq Y\mid C\text{ is a nontrivial, convex cone with }K\setminus\{0\}\subseteq\operatorname{int}\,C\}.$
Definition 2.22 (Proper efficiency in the sense of Henig).
A point $\bar{x}\in\Omega$ is said to be a Henig properly efficient solution of ($P$) if there is $C\in\mathcal{D}(K)$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C)$. The set of all Henig properly efficient solutions of ($P$) is denoted by $\operatorname{PEff}_{He}(\Omega\mid f,K)$.
The following relationships between the different solution concepts are valid. The proof is straightforward.
Proposition 2.23.
Consider a cone $K\subseteq Y$ with $0\notin\operatorname{int}\,K$ (or equivalently, $K\neq Y$). Then, the following assertions hold:
  • $1^{\circ}$.
    $\operatorname{PEff}_{He}(\Omega\mid f,K)\subseteq\operatorname{Eff}(\Omega\mid f,K)\subseteq\operatorname{WEff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    $\operatorname{Eff}(\Omega\mid f,C)=\operatorname{WEff}(\Omega\mid f,K)$ for $C:=(\operatorname{int}\,K)\cup\{0\}$.
  • $3^{\circ}$.
    $\operatorname{PEff}_{He}(\Omega\mid f,K)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{Eff}(\Omega\mid f,C)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{WEff}(\Omega\mid f,C)$.
3 Scalarization in Vector Optimization
Under the assumptions of Section 2.4, consider the following scalarization of the vector optimization problem ($P$):
$\begin{cases}(\varphi\circ f)(x)\to\min\\ x\in\Omega,\end{cases}$ ($P_{\varphi}$)
where $\varphi:Y\to\mathbb{R}$ is a (frequently parameter-dependent) function. Clearly, to get useful relationships between the problems ($P$) and ($P_{\varphi}$) one needs to impose certain properties on $\varphi$. By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get some knowledge about the original vector problem ($P$). Applying such a strategy is called scalarization method in the literature of vector optimization. Within such methods, the function $\varphi$ is called scalarizing function. For more details, we also refer the reader to the standard books of vector and set optimization by Boţ, Grad and Wanka [5], Göpfert et al. [23], Jahn [42], and Khan, Tammer and Zălinescu [49].
In this paper, we like to analyze the relationships between solutions of ($P_{\varphi}$) and the (weakly, properly) efficient solutions of ($P$) for the case that $\varphi$ satisfies certain monotonicity properties. For doing this, we recall monotonicity concepts for the function $\varphi$ (c.f. Jahn [42, Def. 5.1]).
Definition 3.1.
A scalarizing function $\varphi:Y\to\mathbb{R}$ is said to be
  • $K$-increasing if $\forall\,y,\overline{y}\in Y,\,y\leq_{K}\overline{y}:\;\varphi(y)\leq\varphi(\overline{y}).$
  • strictly $K$-increasing if $\forall\,y,\overline{y}\in Y,\,y<_{K}\overline{y}:\;\varphi(y)<\varphi(\overline{y}).$
  • strongly $K$-increasing if $\forall\,y,\overline{y}\in Y,\,y\lneq_{K}\overline{y}:\;\varphi(y)<\varphi(\overline{y}).$
As an immediate consequence, any strongly $K$-increasing function is $K$-increasing, and if $K\neq Y$, strictly $K$-increasing as well.
Lemma 3.2.
Consider the subadditive functional $\varphi:Y\to\mathbb{R}$ and let $K,C$ be two cones in $Y$. The following assertions hold:
  • $1^{\circ}$.
    If $\varphi$ is (strictly) $-K$ representing, then $\varphi$ is (strictly) $K$-increasing.
  • $2^{\circ}$.
    If $\operatorname{int}K\subseteq\operatorname{int}C$ and $\varphi$ is strictly $C$-increasing, then $\varphi$ is strictly $K$-increasing.
  • $3^{\circ}$.
    If $K\subseteq C$ and $\varphi$ is (strongly) $C$-increasing, then $\varphi$ is (strongly) $K$-increasing.
Proof 3.3.
$1^{\circ}$ We only prove the strict case, since the other case is similar. The case $\operatorname{int}\,K=\emptyset$ is clear. Now assume that $\operatorname{int}\,K\neq\emptyset$. Consider $y^{1},y^{2}\in Y$ with $y^{2}<_{K}y^{1}$, i.e., $y^{2}-y^{1}\in-\operatorname{int}\,K$. Since $\varphi$ is strictly $-K$ representing, this is equivalent to $\varphi(y^{2}-y^{1})<0$. Thus, we derive $\varphi(y^{2})=\varphi(y^{1}+y^{2}-y^{1})\leq\varphi(y^{1})+\varphi(y^{2}-y^{1})<\varphi(y^{1}).$ This shows that $\varphi$ is strictly $K$-increasing. The proofs of $2^{\circ}$ and $3^{\circ}$ are straightforward.
The following scalarization results involving cone-monotone functions are well-known.
Lemma 3.4 (c.f. [50, Lem. 5.6], [27, Lem. 5.2]).
Consider a real-valued function $\varphi:Y\to\mathbb{R}$ and a cone $K\subseteq Y$. Then, the following assertions hold:
  • $1^{\circ}$
    If $\varphi$ is strictly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    If $\varphi$ is strongly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $3^{\circ}$.
    If $\varphi$ is $K$-increasing, and ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)=\{\overline{x}\}$ for some $\overline{x}\in\Omega$, then $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $\varphi$ is strictly $C$-increasing or strongly $C$-increasing for some $C\in\mathcal{D}(K)$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
For a (strictly) $-K$ representing function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$ we have the following result.
Lemma 3.5 (cf. [43, Th. 4.1]).
Suppose that $K\subseteq Y$ is a cone. Assume that $\varphi:Y\to\mathbb{R}$ is a function with $\varphi(0)=0$. Then, for any $\bar{x}\in\Omega$, the following assertions hold:
  • $1^{\circ}$.
    If $\varphi$ is $-K$ representing, then
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)\iff\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
  • $2^{\circ}$.
    If $\varphi$ is strictly $-K$ representing, then
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)\iff\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
Next, we recall two well-known nonlinear scalarization methods in vector optimization which are due to Pascoletti and Serafini [53], and Gerstewitz [20, 21]. For a-priori given parameters $a\in Y$ and $k\in Y\setminus\{0\}$, in [53] the following scalar optimization problem is used
$\begin{cases}\lambda\to\min\\ f(x)\in\lambda k+(a-K),\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ (3.1)
while in [20, 21] the scalar optimization problem is given as
$\begin{cases}\varphi^{G}_{a-K,k}(f(x))\to\min\\ x\in\Omega,\end{cases}$ (3.2)
where the nonlinear scalarizing function $\varphi^{G}_{a-K,k}:Y\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty,+\infty\}$ is defined (for $k\in K\setminus\ell(K)$ and $K$ closed) by
$\varphi^{G}_{a-K,k}(y):=\inf\{t\in\mathbb{R}\mid y\in tk+(a-K)\}\quad\text{forall }y\in Y.$ (3.3)
Many facts about the problems (3.1) and (3.2), as well as about the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$, are known (see, e.g., [20, 21], [53], [22], [12, 13], [49], [57], [23]). For the sake of completeness, we recall some important facts below.
Lemma 3.6.
Suppose that $K\subseteq Y$ is a closed, convex cone and take $k\in K\setminus\ell(K)$. Then, the scalar problems (3.1) and (3.2) are equivalent (i.e., each problem has an optimal solution if and only if the other one has an optimal solution, and the optimal solutions $x$ as well as the optimal value of both problems coincide).
Proof 3.7.
Under the given assumptions, the function $\varphi^{G}_{a-K,k}$ has the level set property $f(x)\in\lambda k+(a-K)\iff\varphi^{G}_{a-K,k}(f(x))\leq\lambda$ (see [23, Th. 2.4.1]). Thus, (3.1) is equivalent to
$\begin{cases}\lambda\to\min\\ \varphi^{G}_{a-K,k}(f(x))\leq\lambda,\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$
respectively, to (3.2).
In the next lemma, we state classical relationships between the vector problem ($P$) and the scalar problem (3.1) (see, e.g., Eichfelder [12, Th. 2.1 and Cor. 2.9]).
Lemma 3.8.
Suppose that $K\subseteq Y$ is a convex cone. Then, the following assertions are valid:
  • $1^{\circ}$.
    If $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in K\setminus\ell(K)$. More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1).
  • $2^{\circ}$.
    If $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,K$.
  • $3^{\circ}$.
    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$. If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$ solves the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$ and $f(\bar{x})\in\bar{\lambda}k+(a-\operatorname{bd}\,K)$.
Alternatively, one can immediately obtain scalarization results by using the relationships between (3.1) and (3.2), and properties of the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$ defined in (3.3) (see [57, Ch. 6] for an extended overview on this topic).
Lemma 3.9.
Suppose that $K\subseteq Y$ is a closed, convex cone, and take $k\in K\setminus\ell(K)$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$
  • $2^{\circ}$.
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
  • $3^{\circ}$.
    $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the solution set of the problem (3.1) for $a=f(\bar{x})$.
Lemma 3.10.
Suppose that $K\subseteq Y$ is a nontrivial, closed, solid, convex cone, and take $k\in\operatorname{int}\,K$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$
  • $2^{\circ}$.
    $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
  • $3^{\circ}$.
    $(\bar{x},0)$ is a solution of the problem (3.1) for $a=f(\bar{x})$.
The scalar optimization problems of type (3.1) and (3.2) involved in the classical scalarization results given in Lemmata 3.8, 3.9 and 3.10 have the drawback that the parameter $a$ depends on the solution $\bar{x}$ (more precisely, we have $a=f(\bar{x})$). In order to avoid this disadvantage, Jahn [43, Th. 4.2] stated a scalarization result where the parameter $a$ does not depend on the solution $\bar{x}$.
In the following proposition, we present a new scalarization result that contains the result of Jahn [43, Th. 4.2] as a special case (see Remark 3.13 for details).
Proposition 3.11.
Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the scalar optimization problem (3.1).
If, in addition, $K$ is closed, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $3^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof 3.12.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 3.8 ($3^{\circ}$).
Now, we show the implication $1^{\circ}\Longrightarrow 2^{\circ}$. Let $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$. Take any $k\in\mathcal{P}_{k}\,(\subseteq K\setminus\ell(K))$. By Lemma 3.8 ($1^{\circ}$), the set of solutions of the problem (3.1) (with $f(\bar{x})$ in the role of $a$) is given by $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$. Note that, for any $x\in\Omega$, the pair ($x,0$) solves (3.1) (with $f(\bar{x})$ in the role of $a$) if and only if ($x,t$) solves (3.1) with $f(\bar{x})-tk$ in the role of $a$ for all $t\in\mathbb{R}$. By our assumption $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, there are $\bar{a}\in\mathcal{P}_{a}$, $\bar{k}\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ such that $f(\bar{x})=\bar{a}+s\bar{k}$. Thus, $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) with $\bar{a}=f(\bar{x})-s\bar{k}\in\mathcal{P}_{a}$ and $\bar{k}\in\mathcal{P}_{k}$ in the role of $a$ and $k$.
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
Remark 3.13.
[…]
For the concept of weak efficiency we have the following result.
Proposition 3.14.
Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq\operatorname{int}\,K$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of the scalar optimization problem (3.1).
If, in addition, $K$ is closed and nontrivial, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $3^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that
    $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof 3.15.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 3.8 ($4^{\circ}$) taking into account that $\operatorname{int}\,K\subseteq K\setminus\{0\}$, or equivalently $K\neq Y$ (since (3.1) has a solution), while the proof of the implication $1^{\circ}\Longrightarrow 2^{\circ}$ is similar to the proof of the corresponding implication in Proposition 3.11.
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
Remark 3.16.
Assuming that $K$ has a normlike-base $B_{K}(\psi)$ (which means that the seminorm $\psi:Y\to\mathbb{R}$ is positive on $K\setminus\{0\}$) and $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq a+K\setminus\ell(K)$ (respectively, $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq a+\operatorname{int}\,K$) for some $a\in Y$, in Proposition 3.11 (respectively, Proposition 3.14), one may take $\mathcal{P}_{a}:=\{a\}$, $\mathcal{P}_{k}:=B_{\{0\}\cup(K\setminus\ell(K))}(\psi)$ (respectively, $\mathcal{P}_{k}:=B_{\{0\}\cup\operatorname{int}\,K}(\psi)$ if $0\notin\operatorname{int}\,K$, or equivalently, $K\neq Y$) and $\mathcal{R}:=\mathbb{P}$.
The following proposition provides possible choices for the sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ to ensure $f[D]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ with $D={\mathrm{E}ff}(\Omega\mid f,K)$, $D={\mathrm{W}Eff}(\Omega\mid f,K)$ and $D=\Omega$, respectively, in Propositions 3.11 and 3.14.
Proposition 3.17.
Let $K\subseteq Y$ be a nontrivial, convex cone and let $y^{*}\in Y^{*}\setminus\{0\}$, $a\in Y$ and $k\in K\setminus\{0\}$ be given. Suppose that $y^{*}(k)>0$ (e.g. if $y^{*}\in K^{\#}$). Define
$H_{y^{*},a}^{\sim}:=\{y\in Y\mid y^{*}(y)\sim y^{*}(a)\}\quad\text{for all }\sim\in\{=,<,>\}.$
Then, the following assertions are valid:
  • $1^{\circ}$.
    $H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $H_{y^{*},a}^{<}=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
  • $2^{\circ}$.
    If $y^{*}\in K^{\#}$, then $a+K\setminus\{0\}\subseteq H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $a-K\setminus\{0\}\subseteq H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
  • $3^{\circ}$.
    If $y^{*}\in K^{+}$ and $k\in\operatorname{int}\,K$, then $a+\operatorname{int}\,K\subseteq H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $a-\operatorname{int}\,K\subseteq H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
  • $4^{\circ}$.
    $Y=\bigcup_{\sim\in\{<,=,>\}}\,H_{y^{*},a}^{\sim}=H_{y^{*},a}^{=}+\mathbb{R}\cdot k$.
Proof 3.18.
$1^{\circ}$. Let us verify the equality $H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$. Since $a\in H_{y^{*},a}^{=}$, $y^{*}(k)>0$ and $\mathbb{P}\cdot k\subseteq K\setminus\{0\}$, we conclude $H_{y^{*},a}^{=}+\mathbb{P}\cdot k\subseteq H_{y^{*},a}^{>}$. Take some $y\in H_{y^{*},a}^{>}$. Noting that $y-tk\in H_{y^{*},a}^{=}$ for some $t>0$, we derive $y=y-tk+tk\in H_{y^{*},a}^{=}+\mathbb{P}\cdot k$, which shows the reverse inclusion. Applying the result for the cone $-K$, the functional $-y^{*}$ and the direction $-k$ (in the role of $K$, $y^{*}$ and $k$), we get immediately $H_{y^{*},a}^{<}=H_{(-y^{*}),a}^{>}=H_{(-y^{*}),a}^{=}+\mathbb{P}\cdot(-k)=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
$2^{\circ}$. Assume that $y^{*}\in K^{\#}$, which means that $y^{*}(x)>0$ for all $x\in K\setminus\{0\}$. Then, using $1^{\circ}$ we easily derive $a+K\setminus\{0\}\subseteq H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$. Applying the result for $-K$, $-y^{*}$ and $-k$ (in the role of $K$, $y^{*}$ and $k$), we obtain $a-K\setminus\{0\}\subseteq H_{-y^{*},a}^{=}-\mathbb{P}\cdot k=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
$3^{\circ}$. Define $C:=({\mathrm{i}nt}\,K)\cup\{0\}$. Note that ${\mathrm{i}nt}\,K=C\setminus\{0\}$ and $K^{+}\setminus\{0\}=\{x^{*}\in Y^{*}\mid\forall x\in{\mathrm{i}nt}\,K:x^{*}(x)>0\}=C^{\#}$ by [50, Th. 4.8, Cor. 4.9] (if $K$ is a solid, convex cone). Applying assertion $2^{\circ}$ for $C$ in the role of $K$ we get the result.
$4^{\circ}$. The first equality is obvious, while the second equality follows immediately by $1^{\circ}$.
Remark 3.19.
Let the assumptions of Proposition 3.17 be fulfilled. Eichfelder’s adaptive scalarization method in multiobjective optimization proposed in [12, 13] is based on parameterized scalar optimization problems of type (3.1) and (3.2), respectively, with different choices of the reference point $a\in H_{y^{*},a}^{=}$ and an a-priori fixed direction $k\in K\setminus\{0\}$. Note that Proposition 3.11 generalizes a result by Eichfelder [13, Th. 3.2] taking into account that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq Y=\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ for $\mathcal{P}_{a}:=H_{y^{*},a}^{=}$, $\mathcal{P}_{k}:=\{k\}$ and $\mathcal{R}:=\mathbb{R}$ by Proposition 3.17 ($4^{\circ}$).
4 Nonlinear Cone Separation Conditions
The key tool to prove our main Bishop-Phelps type scalarization results in vector optimization in Section 5.3 will be the separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. In this section we present all the necessary basics for nonlinear cone separation.
Consider a real topological-linear space $Y$ and a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$. For a cone $C\subseteq Y$, we use the notations:
$S_{C}(\psi):=\operatorname{conv}(B_{C}(\psi))\quad\text{and}\quad S_{C}^{0}(\psi):=\operatorname{conv}(\{0\}\cup B_{C}(\psi)).$
If the underlying seminorm (norm) $\psi$ is clear in the context, we will simply use $S_{C}$ and $S_{C}^{0}$ instead of $S_{C}(\psi)$ and $S_{C}^{0}(\psi)$.
Throughout the next section, we consider a nontrivial cone $K\subseteq Y$ with the normlike-base $B_{K}=B_{K}(\psi)$, and a nontrivial cone $A\subseteq Y$ with the normlike-base $B_{A}=B_{A}(\psi)$. The following conditions will be of interest in our upcoming auxiliary results:
$(\operatorname{cl}\,S_{A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (4.1)
$A\cap(\operatorname{cl}(\operatorname{conv}(-K))=\{0\},$ (4.2)
$0\notin\operatorname{cl}\,S_{-K}(\psi).$ (4.3)
The following proposition will later play an important role in proving a Bishop-Phelps type scalarization result (namely Corollary 5.28) for vector optimization problems in real locally convex spaces.
Proposition 4.1.
Assume that (4.2) and (4.3) are valid, and one of the following statements is satisfied:
  1. a)
    $A$ is closed and convex.
  2. b)
    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that
    $\operatorname{cl}\,S_{-K}(\psi)$ =$=\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ (4.4)
    is valid, and one of the following conditions is satisfied:
    $B_{A}\text{ is compact},$ (4.5)
    $S_{A}^{0}\text{ is closed},$ (4.6)
    $A\text{ is closed and }\operatorname{cl}\,B_{A}\text{ is compact},$ (4.7)
    $A\text{ is closed and }\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\text{ is closed}.$ (4.8)
Then, (4.1) is valid.
Proof 4.2.
a) Assuming $A$ is closed and convex, by (4.2) we get $\operatorname{cl}\,S_{A}^{0}\cap\operatorname{cl}\,S_{-K}\subseteq A\cap\operatorname{cl}(\operatorname{conv}(-K))=\{0\}$, and so by (4.3) we conclude (4.1).
b) Obviously, (4.2) and (4.3) imply $A\cap\operatorname{cl}\,S_{-K}=\emptyset$. Since there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with (4.4), we get that $x^{*}(x)<\alpha$ for all $x\in A\cap(\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi})$. Note that $B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$, and if $A$ is closed, then $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$. Consider cases:
Case 1: (4.5) is valid. By the continuity of $x^{*}$ and the compactness of $B_{A}\cup\{0\}$, we get some $\beta\in(0,\alpha)$ such that $B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Hence, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
Case 2: (4.6) is valid. Then, $\operatorname{cl}\,S_{A}^{0}=S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
Case 3: (4.7) is valid. By the continuity of $x^{*}$ and the compactness of $\operatorname{cl}\,B_{A}\cup\{0\}$ we get some $\beta\in(0,\alpha)$ such that $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Thus, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
Case 4: (4.8) is valid. Then, $\operatorname{cl}\,S_{A}^{0}\subseteq\operatorname{cl}(\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\}))=\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
Recalling that (4.4) is valid, we conclude (4.1).
Remark 4.3.
The result in Proposition 4.1 (regarding assumption b)) is a generalization of a result by Kasimbeyli [45, Lem. 3], where the author studied the Euclidean space $Y=\operatorname{\mathbb{R}}^{n}$, a closed cone $A\subseteq Y$ and a closed convex cone $K\subseteq Y$ with $A\cap(-K)=\{0\}$. Moreover, it is assumed that there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid (hence $0\notin\operatorname{cl}\,S_{-K}$ is valid). The conclusion in [45, Lem. 3] is given as the condition (4.1) with $\operatorname{bd}\,A$ in the role of $A$.
The condition (4.4) involved in b) of Proposition 4.1 can be fulfilled by certain Bishop-Phelps type cones, as the following lemma shows.
Lemma 4.4.
Consider a real topological-linear space $Y$, a continuous seminorm $\psi:Y\to\mathbb{R}$, and a nontrivial Bishop-Phelps type cone $C=C_{\psi}(x^{*},\alpha)$ for some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$. Then,
$\operatorname{cl}\,B_{C}(\psi)=B_{C}(\psi)=\{x\in\operatorname{\mathbb{S}}_{\psi}\mid x^{*}(x)\geq\alpha\},$
and if
$\{e\in Y\setminus\{0\}\mid\psi(e)=1,x^{*}(e)=0\}\neq\emptyset$ (4.9)
(e.g., if $Y$ has dimension two or higher, and $\psi$ is a norm), then
$\operatorname{cl}\,S_{C}(\psi)=S_{C}(\psi)$ =$=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}.$
Proof 4.5.
The equality $B_{C}=\{x\in\operatorname{\mathbb{S}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ is obvious. Since $\psi$ is continuous, the sets $C$ and $\operatorname{\mathbb{S}}_{\psi}$ are closed, hence $B_{C}$ is closed. Define $D:=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$. Since $D$ is a convex set with $B_{C}\subseteq D$, the inclusion $S_{C}\subseteq D$ is clear.
Let us show that $S_{C}\supseteq D$. Take some $d\in D$. We consider two cases:
Case 1: Assume that $\psi(d)=1$. Then, $d\in B_{C}\subseteq S_{C}$.
Case 2: Assume that $\psi(d)<1$. Put $\gamma:=x^{*}(d)\,(\geq\alpha)$. By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$. Then,
$x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma\quad\text{ for all }t\geq 0$
and
$\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty\quad\text{ for }t\to\infty.$
Taking into account the continuity of $t\mapsto h_{+}(t):=\psi(d+te)$ and $t\mapsto h_{-}(t):=\psi(d-te)$, as well as the conditions $h_{+}(0)=h_{-}(0)=\psi(d)<1$, $\lim_{t\to\infty}\,h_{+}(t)=\lim_{t\to\infty}\,h_{-}(t)=\infty$, by the well-known intermediate value theorem there are $t_{+},t_{-}\in\operatorname{\mathbb{P}}$ with $h_{+}(t_{+})=h_{-}(t_{-})=1$. Since $x^{*}(d+t_{+}e)=x^{*}(d-t_{-}e)=\gamma\geq\alpha$ we infer $d-t_{-}e\in B_{C}$ and $d+t_{+}e\in B_{C}$, hence $d\in\operatorname{conv}(\{d-t_{-}e,\,d+t_{+}e\})\subseteq S_{C}.$
Of course, $S_{C}=D$ and the closedness of $D$ imply $\operatorname{cl}\,S_{C}=S_{C}$.
Remark 4.6.
Kasimbeyli and Kasimbeyli [44, Th. 4.5] stated a similar result as in Lemma 4.4 in a reflexive Banach space setting (with $\psi(\cdot):=||\cdot||$), but note that the assumption given in (4.9), which in this setting means that $Y$ has dimension two or higher, was not considered in [44, Th. 4.5]. In the one dimensional space ($Y:=\mathbb{R}$, ||$\cdot$|| ⋅ |), the condition (4.9) fails and the conclusion of Lemma 4.4 may fail. To see this, consider $Y:=\mathbb{R}$, $x^{*}:=1$ and $\alpha\in(0,1]$. Then, $C=C(x^{*},\alpha)=\mathbb{R}_{+}$ and $B_{C}(|\cdot|)=\{1\}$ as well as $\operatorname{cl}\,S_{C}(|\cdot|)=\{1\}$ while $\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}=[\alpha,1]$.
To prove some more specific results in the normed space setting in the upcoming Proposition 4.11 and Corollary 5.30 (see Section 5.3), we need additional lemmata.
Lemma 4.7.
Consider a real normed space $(Y,||\cdot||)$, a (weakly) compact set $\Omega_{1}\subseteq Y$ with $0\notin\Omega_{1}$, and a (weakly) closed, bounded set $\Omega_{2}\subseteq Y$ with $\Omega_{2}\subseteq\operatorname{\mathbb{R}}_{+}\cdot\Omega_{1}$. Then, $\Omega_{2}$ is (weakly) compact.
Proof 4.8.
Note that a set $\Omega_{2}$ in a normed space is weakly bounded if and only if it is (norm) bounded. First, we prove that there exists $t>0$ such that
$\Omega_{2}\subseteq[0,t]\cdot\Omega_{1}.$ (4.10)
On the contrary, assume that there are sequences $(x_{n}^{1})$, $(x_{n}^{2})$ and $(t_{n})$ with $x_{n}^{1}\in\Omega_{1}$, $x_{n}^{2}\in\Omega_{2}$, $t_{n}\in\mathbb{P}$ such that $x_{n}^{2}=t_{n}\cdot x_{n}^{1}$ with $t_{n}\to\infty$. By the boundedness of $\Omega_{2}$ there is $\lambda>0$ such that $||x||\leq\lambda$ for all $x\in\Omega_{2}$. Consequently, we have
$||x_{n}^{1}||=||t_{n}^{-1}x_{n}^{2}||=t_{n}^{-1}||x_{n}^{2}||\leq t_{n}^{-1}\lambda\to 0\quad\text{for }t_{n}\to\infty.$
This means that $0\in\operatorname{cl}\,\Omega_{1}\subseteq\operatorname{cl}_{w}\,\Omega_{1}$. Noting that the compact (respectively, weakly compact) set $\Omega_{1}$ is closed (respectively, weakly closed), we get a contradiction.
Define the function $h:\operatorname{\mathbb{R}}\times Y\to Y$ by $h(s,\omega):=s\cdot\omega$ for all $(s,\omega)\in\operatorname{\mathbb{R}}\times Y$. Since $[0,t]\times\Omega_{1}$ is (weakly) compact in ×$\operatorname{\mathbb{R}}\times Y$, and the function $h$ is (weakly) continuous by the topological vector space properties, we get that $[0,t]\cdot\Omega_{1}$ is (weakly) compact. From (4.10) we derive that the (weakly) closed subset $\Omega_{2}$ of $[0,t]\cdot\Omega_{1}$ is (weakly) compact as well.
Lemma 4.9.
Consider a real normed space $(Y,||\cdot||)$ and a nontrivial cone $C\subseteq Y$. Then, the following assertions are valid:
  • $1^{\circ}$.
    If $B_{C}(||\cdot||)$ is compact, then $B_{C}(||\cdot||^{\prime})$ is compact for all $||\cdot||^{\prime}\sim||\cdot||$.
  • $2^{\circ}$.
    If $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, then $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is weakly compact for all $||\cdot||^{\prime}\sim||\cdot||$.
  • $3^{\circ}$.
    If $\operatorname{cl}\,S_{C}(||\cdot||)$ is (weakly) compact with $0\notin\operatorname{cl}\,S_{C}(||\cdot||)$, then $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is (weakly) compact for all $||\cdot||^{\prime}\sim||\cdot||$.
  • $4^{\circ}$.
    If $Y$ is reflexive (or equivalently, $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact), then $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact for all $||\cdot||^{\prime}\sim||\cdot||$.
Proof 4.10.
Consider any $||\cdot||^{\prime}\sim||\cdot||$.
$1^{\circ}$. Obviously, $B_{C}(||\cdot||^{\prime})\subseteq C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $B_{C}(||\cdot||^{\prime})$ is bounded. Moreover, according to [23, Lemma 2.1.43 (iii)], for the bounded and closed norm-base $B_{C}(||\cdot||)$ with $0\notin B_{C}(||\cdot||)$ we have $C=\mathbb{R}_{+}\cdot B_{C}(||\cdot||)$ is closed, hence $B_{C}(||\cdot||^{\prime})$ is closed as well. Then, the compactness of $B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the compact set $\Omega_{1}:=B_{C}(||\cdot||)$ and the closed, bounded set $\Omega_{2}:=B_{C}(||\cdot||^{\prime})$. Note that $0\notin B_{C}(||\cdot||)=\Omega_{1}$.
$2^{\circ}$. Since $C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)\subseteq\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)\subseteq\operatorname{cl}_{w}\,C$, we have $\operatorname{cl}_{w}\,C=\operatorname{cl}_{w}(\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||))$. Taking into account that $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is a weakly closed, bounded set with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, applying [23, Lemma 2.1.43 (iii)] we conclude that $\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly closed (remark that the weak topology of the normed space $Y$ is Hausdorff). Consequently, $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}_{w}\,C=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is bounded. Then, the weak compactness of $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the weakly compact set $\Omega_{1}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ and the weakly closed, bounded set $\Omega_{2}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$.
$3^{\circ}$. We have $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}(\operatorname{conv}\,C)=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}\,S_{C}(||\cdot||)$, where the latter equality is discussed in [29, Rem. 2.5]. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence the set $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is bounded. Finally, the (weak) compactness of $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the (weakly) compact set $\Omega_{1}:=\operatorname{cl}\,S_{C}(||\cdot||)$ and the (weakly) closed, bounded set $\Omega_{2}:=\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$.
$4^{\circ}$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence both weakly closed sets $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact if $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact.
To later prove a Bishop-Phelps type scalarization result (namely Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
Proposition 4.11.
Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$. Suppose that (4.2) and (4.3) with $\psi(\cdot):=||\cdot||$ are valid. Moreover, assume that one of the following statements is valid:
  1. a)
    $B_{A}(||\cdot||)$ is compact.
  2. b)
    $A$ is weakly closed, and either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
Then, there is $||\cdot||^{\prime}\sim||\cdot||$ such that (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
Proof 4.12.
Suppose that (4.2) and (4.3) (with $\psi(\cdot):=||\cdot||$) are valid. First, note that the following assertions are known to be equivalent (see [14, Lem. 1.25 and 1.26], [29, Lem. 3.4] and [54, Th. 3.2]):
  • $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ (i.e., (4.3) with $\psi(\cdot):=||\cdot||$ is valid);
  • $0\notin\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||)$;
  • $(-\operatorname{cl}(\operatorname{conv}\,K))^{a+}(||\cdot||)\cap(Y^{*}\times\mathbb{P})\neq\emptyset$;
  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is well-based;
  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is representable as a Bishop-Phelps cone (cf. Example 2.4).
Thus, under assumption (4.3) (with $\psi(\cdot):=||\cdot||$), there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $-\operatorname{cl}(\operatorname{conv}\,K)=C_{||\cdot||^{\prime}}(x^{*},\alpha)$. By Lemma 4.4, in the real normed space $(Y,||\cdot||^{\prime})$ with dimension two or higher we have $\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\{x\in\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\mid x^{*}(x)\geq\alpha\}$, hence (4.3) and (4.4) with $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ are valid. Let us finish the proof by assuming a) and b), respectively:
a) In view of Lemma 4.9 ($1^{\circ}$), $B_{A}(||\cdot||^{\prime})$ is compact if $B_{A}(||\cdot||)$ is compact. Proposition 4.1 b) applied for the space $(Y,||\cdot||^{\prime})$, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures $(\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap(\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime}))=\emptyset$, hence (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid, taking into account that $\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
b) In view of Lemma 4.9 ($2^{\circ}$, $4^{\circ}$), $\operatorname{cl}_{w}\,B_{A}(||\cdot||^{\prime})$ is weakly compact if either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$. Proposition 4.1 b) applied for the space $Y$ endowed with the weak topology, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$, taking into account that $\operatorname{cl}_{w}\,S_{A}^{0}(||\cdot||^{\prime})=\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime})$ and $\operatorname{cl}_{w}\,S_{-\operatorname{cl}_{w}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
Furthermore, we recall some recent results on cone separation in real topological-linear (locally convex) spaces established in [28].
Proposition 4.13 ([28, Th. 5.4]).
Assume that $Y$ is a real topological-linear space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$, and $K$ is solid. Suppose that $S_{-K}$ is solid and $S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ Then, there exists $(x^{*},\alpha)\in K^{a\circ}$ such that
$x^{*}(a)+\alpha\psi(a)$ $\geq 0\geq x^{*}(k)+\alpha\psi(k)\quad\mbox{for all }a\in A\mbox{and }k\in-K,$
$x^{*}(a)+\alpha\psi(a)$ $\geq 0>x^{*}(k)+\alpha\psi(k)\quad\mbox{for all }a\in A\mbox{ and}k\in\operatorname{int}(-K),$
or equivalently,
$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad K\subseteq C_{\psi}(x^{*},\alpha)\quad$ (4.11)
$\text{and}\quad\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha).$
Proposition 4.14 ([28, Th. 5.8]).
Assume that $Y$ is a real topological-linear space, $K,A\subseteq Y$ are nontrivial, convex cones with normlike-bases $B_{K}$ and $B_{A}$, and $K$ is solid. Suppose that $S_{-K}$ is solid. Then, the following assertions are equivalent:
  • $1^{\circ}$.
    $S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset$.
  • $2^{\circ}$.
    $A\cap\operatorname{int}(-K)=\emptyset$.
  • $3^{\circ}$.
    There exists $(x^{*},\alpha)\in K^{a\circ}$ such that (4.11) is valid.
Proposition 4.15 ([28, Th. 5.9]).
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.1) is valid. Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$x^{*}(a)+\alpha\psi(a)>0>x^{*}(k)+\alpha\psi(k)\quad\mbox{for all }a\in A\setminus\{0\}\mbox{ and }k\in-K\setminus\{0\},$
or equivalently,
$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$ (4.12)
From the general strict cone separation result in Proposition 4.15 and our Propositions 4.1 and 4.11 we immediately derive some new variants of strict cone separation results involving the classical separation condition (4.2).
Corollary 4.16.
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.2) and (4.3) are valid. Moreover, assume that the assumptions in a) or in b) of Proposition 4.1 are satisfied. Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that (LABEL:sep_conditions_strict) is valid.
Corollary 4.17.
Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ and ${\mathrm{c}l}\,S_{A}^{0}(||\cdot||)$ is weakly compact (e.g., if $Y$ is reflexive), and (4.2) and (4.3) with $\psi(\cdot):=||\cdot||$ are valid. Moreover, assume that the assumptions in a) or in b) of Proposition 4.11 are satisfied. Then, there are $||\cdot||^{\prime}\sim||\cdot||$ and $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that (LABEL:sep_conditions_strict) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
5 Bishop-Phelps Type Scalarization in Vector Optimization
Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$. Then, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, and any given reference point $a\in Y$, one defines $\varphi_{x^{*},\alpha}^{a}:Y\to\operatorname{\mathbb{R}}$ as
$\varphi_{x^{*},\alpha}^{a}(y):=\varphi_{x^{*},\alpha}(y-a)\quad\text{for all }y\in Y.$
The following parameterized scalar optimization problem of the vector problem ($P$) (in the sense of Kasimbeyli [45]) will be of interest:
$\begin{cases}(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\varphi_{x^{*},\alpha}(f(x)-a)=x^{*}(f(x)-a)+\alpha\psi(f(x)-a)\to\min\\ x\in\Omega.\end{cases}$ ($P_{\varphi}^{a}$)
In addition, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$ and any given reference point $a\in Y$ and direction $k\in Y\setminus\{0\}$, we are interested in the parameterized scalar problem (in the sense of Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]):
$\begin{cases}\lambda\to\min\\ \varphi_{x^{*},\alpha}(f(x)-a-\lambda k)\leq 0,\\ (x,\lambda)\in\Omega\times\mathbb{R}.\end{cases}$ ($P_{\varphi}^{a,k}$)
5.1 Basic scalarization results
First, we present monotonicity properties of the seminorm-linear function $\varphi_{x^{*},\alpha}$ with respect to a Bishop-Phelps type cone $C_{\psi}(x^{*},\alpha)$.
Lemma 5.1.
The following assertions hold:
  • $1^{\circ}$.
    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, then $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing.
  • $2^{\circ}$.
    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
  • $3^{\circ}$.
    If $K$ is solid and $(x^{*},\alpha)\in K^{a\circ}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
  • $4^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
  • $5^{\circ}$.
    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, then $\varphi_{x^{*},\alpha}$ is strongly $(C_{\psi}^{>}(x^{*},\alpha)\cup\{0\})$-increasing.
Proof 5.2.
$1^{\circ}$ Noting that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing, the assertion follows from Lemma 3.2 ($1^{\circ}$).
$2^{\circ}$ If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$, there is nothing to prove. If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, using the assumption $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ we have that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing, and the result follows from Lemma 3.2 ($1^{\circ}$).
$3^{\circ}$ and $4^{\circ}$ Using the assumptions it follows that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, so by assertion $2^{\circ}$ we get the conclusion.
$5^{\circ}$ It follows from the definitions.
The next lemma, which studies the monotonicity properties of $\varphi_{x^{*},\alpha}$ with respect to the original cone $K$, extends a corresponding result derived by Kasimbeyli [45, Th. 3.5] in a normed setting.
Lemma 5.3.
Consider some $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
  • $1^{\circ}$.
    $(x^{*},\alpha)\in K^{a+}$ $\iff$ $\varphi_{x^{*},\alpha}$ is $K$-increasing.
  • $2^{\circ}$.
    $(x^{*},\alpha)\in K^{a\circ}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing.
  • $3^{\circ}$.
    $(x^{*},\alpha)\in K^{a\#}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing.
Proof 5.4.
$1^{\circ}$ By Lemma 5.1 ($1^{\circ}$), the function $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing. If $(x^{*},\alpha)\in K^{a+}$, then $K\subseteq C_{\psi}(x^{*},\alpha)$, hence $\varphi_{x^{*},\alpha}$ is $K$-increasing by Lemma 3.2 ($3^{\circ}$). Conversely, assuming $\varphi_{x^{*},\alpha}$ is $K$-increasing, for $k\in K$ we have $-k\leq_{K}0$, hence $\varphi_{x^{*},\alpha}(-k)\leq\varphi_{x^{*},\alpha}(0)=0$, which means that $x^{*}(k)-\alpha\psi(k)\geq 0$. Noting that $x^{*}\in K^{+}$, we conclude $(x^{*},\alpha)\in K^{a+}$.
$2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid. Now assume that $\operatorname{int}\,K\neq\emptyset$. If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so the function $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 5.1 ($2^{\circ}$) and 3.2 ($2^{\circ}$). The proof of the reverse implication is similar to $1^{\circ}$.
$3^{\circ}$ If $(x^{*},\alpha)\in K^{a\#}$, then $K\subseteq C_{\psi}^{>}(x^{*},\alpha)\cup\{0\}$, so $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing by Lemmata 5.1 ($5^{\circ}$) and 3.2 ($3^{\circ}$). Also here the proof of the reverse implication is similar to $1^{\circ}$.
Now, we are able to present first relationships between the parameterized scalar problem ($P_{\varphi}^{a}$) and the vector problem ($P$).
Theorem 5.5.
For any $a\in Y$, the following assertions hold:
  • $1^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\circ}$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $3^{\circ}$.
    Assume that $\overline{x}\in\Omega$. If $(x^{*},\alpha)\in K^{a+}$ and ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\{\overline{x}\}$, then $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
Proof 5.6.
$1^{\circ}$ Consider $(x^{*},\alpha)\in K^{a\circ}$. By Lemma 5.3 ($2^{\circ}$) $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing. It is easy to check that $\varphi_{x^{*},\alpha}^{a}$ is strictly $K$-increasing for any $a\in Y$ as well. Finally, Lemma 3.4 ($1^{\circ}$) yields the conclusion in assertion $1^{\circ}$.
$2^{\circ}$ and $3^{\circ}$ The proof of the assertions is similar to the proof of assertion $1^{\circ}$ by using Lemmata 5.3 and 3.4.
$4^{\circ}$ Take some $(x^{*},\alpha)\in K^{a\#}$. By Lemma 5.1 ($4^{\circ}$) $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing, and so $\varphi_{x^{*},\alpha}^{a}$ as well. Because $(x^{*},\alpha)\in K^{a\#}$ we get $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Taking into account that $C_{\psi}(x^{*},\alpha)$ is a nontrivial convex cone, we conclude that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$. Finally, with the aid of Lemma 3.4 ($4^{\circ}$) we get the desired inclusion.
Remark 5.7.
Consider $\bar{x}\in\Omega$. In some upcoming results, the specific choice $a:=f(\bar{x})$ will be of interest. In this case, the scalarized objective function $\varphi_{x^{*},\alpha}^{a}\circ f$ is given by $(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$ for all $x\in\Omega$. According to Theorem 5.5, for a specific choice of the pair $(x^{*},\alpha)$ one can ensure the (weak, proper) efficiency of $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
Let us shift our focus to relationships between the parameterized scalar problem ($P_{\varphi}^{a,k}$) and the vector problem ($P$) involving a Bishop-Phelps type cone.
Lemma 5.8.
Consider some $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions are valid:
  • $1^{\circ}$.
    If $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$. More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$).
  • $2^{\circ}$.
    If $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
  • $3^{\circ}$.
    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$. If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $4^{\circ}$.
    If $(\bar{x},\bar{\lambda})$ solves the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
Proof 5.9.
Follows immediately from Lemma 3.8 taking into account that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing.
Theorem 5.10.
Consider $\bar{x}\in\Omega$. Take some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, $a\in Y,k\in Y\setminus\{0\}$ and assume that $(\bar{x},\bar{\lambda})$ is a solution of the optimization problem ($P_{\varphi}^{a,k}$). Then:
  • $1^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$ and $C_{\psi}(x^{*},\alpha)$ is solid, then $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\circ}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
Proof 5.11.
$1^{\circ}$. Remark that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in view of Lemma 5.8 ($4^{\circ}$). The latter set is contained in ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)$ (since $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$).
$2^{\circ}$. For the case $\operatorname{int}\,K=\emptyset$ the statement is clear. Now assume that $\operatorname{int}\,K\neq\emptyset$. Again we have $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Here the latter set is contained in ${\mathrm{W}Eff}(\Omega\mid f,K)$ (since $\emptyset\neq\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
Let us summarize our findings related to vector problems involving a Bishop-Phelps type cone in the next two propositions.
Proposition 5.12.
Let $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $2^{\circ}$.
    For any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$, the set $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
  • $3^{\circ}$.
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then the assertions $1^{\circ}-3^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed, then the assertions $1^{\circ}-4^{\circ}$ are equivalent to:
  • $5^{\circ}$.
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
Proof 5.13.
Note that $C_{\psi}(x^{*},\alpha)$ is a convex cone, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi_{x^{*},\alpha}(0)=0$ and $C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))\neq\emptyset$.
$1^{\circ}\iff 2^{\circ}$. The equivalence follows immediately from Lemma 5.8 ($1^{\circ}$, $3^{\circ}$).
$1^{\circ}\iff 3^{\circ}$. The result follows directly from Lemma 3.5 ($1^{\circ}$).
$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$. The result follows from Proposition 3.11 taking into account that ($P_{\varphi}^{a,k}$) is the same problem as (3.1) with $C_{\psi}(x^{*},\alpha)$ in the role of $K$.
Proposition 5.14.
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $2^{\circ}$.
    For any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, the point $(\bar{x},0)$ is a solution of the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $3^{\circ}$.
    $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then the assertions $1^{\circ}$ and $2^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then the assertions $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
  • $5^{\circ}$.
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that
    $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
Proof 5.15.
First, note that $C_{\psi}(x^{*},\alpha)$ is a solid, convex cone and $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi_{x^{*},\alpha}(0)=0$.
$1^{\circ}\iff 2^{\circ}$. The equivalence follows immediately from Lemma 5.8 ($2^{\circ}$, $4^{\circ}$).
$1^{\circ}\iff 3^{\circ}$. Since $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, the function $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$-representing by Lemma 2.12 ($4^{\circ}$) with $\varphi_{x^{*},\alpha}(0)=0$. Then, the result follows from Lemma 3.5 ($2^{\circ}$).
$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$. The result follows directly from the Proposition 3.14.
Remark 5.16.
The results derived in Theorem 5.5 and Propositions 5.12 and 5.14 extend some known results in the literature, which are given in a normed space $(Y,\psi)$, $\psi(\cdot):=||\cdot||$, preordered by a convex cone $K\subseteq Y$:
  • Gasimov [17, Th. 1] proved a corresponding statement as in Theorem 5.5 ($4^{\circ}$) for Benson proper efficiency (instead of Henig proper efficiency).
  • Kasimbeyli [46, Th. 5.4] stated similar assertions as in Theorem 5.5 for a closed, pointed, convex cone $K$ and Benson proper efficiency (instead of Henig proper efficiency).
  • Gasimov [17, Th. 2] (see also Guo and Zhang [31, Th. 4.5]) stated
    $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$
    $\Longrightarrow\{x\in\Omega$ $\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$
    with the cone-valued map $\mathbb{A}$ considered in Benson proper efficiency.
Remark 5.17.
It is worth noting that in Proposition 5.12 (respectively, in Proposition 5.14) the function $\varphi_{x^{*},\alpha}$, which appears in the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$), can be replaced by any $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$. For instance, the nonlinear scalarizing function $\varphi^{G}_{-C_{\psi}(x^{*},\alpha),k}$ (i.e., the function defined in (3.3) with $a=0$ and with $C_{\psi}(x^{*},\alpha)$ in the role of $K$) is $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) if $C_{\psi}(x^{*},\alpha)$ is nontrivial and closed, and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha)$ (respectively, $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
Remark 5.18.
Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$. Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence
$\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{E}ff}(\Omega\mid f,K).$
If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g. if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so (see Proposition 2.23)
$\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K).$
If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence
$\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{W}Eff}(\Omega\mid f,K).$
5.2 Scalarization results for the concept of weak efficiency
In this section, we focus on the solution concept of weak efficiency. More precisely, we will present conditions that ensure that a weakly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
Theorem 5.19.
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid cone, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively. Suppose $S_{-K}$ is solid and
$S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ (5.1)
Then, there is $(x^{*},\alpha)\in K^{a\circ}$ such that $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
Proof 5.20.
Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that
$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha).$
Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have
$(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$
hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 5.14 yields the rest of the conclusion.
Remark 5.21.
In view of Remark 5.18, the point $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.19 belongs to ${\mathrm{W}Eff}(\Omega\mid f,K)$.
In the convex case (i.e., $K$ and $\mathbb{A}(\bar{x})$ are convex cones), using Proposition 4.14 we are able to replace the separation condition (5.1) by $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, as shown in the following result.
Corollary 5.22 (convex case).
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid, convex cone, $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial and convex. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively, and $S_{K}$ is solid.
Then, the conclusion of Theorem 5.19 holds.
Proof 5.23.
Follows from Theorem 5.19 and Proposition 4.14.
Remark 5.24.
Recall that, for any set $\Omega\subseteq Y,$ the quasi-interior of $\Omega$ is the set ${\mathrm{q}i}\,\Omega:=\{x\in\Omega\mid\operatorname{cl}(\mathbb{R}_{+}\cdot(\Omega-x))=Y\}$. In order to ensure that the assumption $A$ is nontrivial and convex from Corollary 5.22 is satisfied it is sufficient to impose that $Y$ is a real Hausdorff locally convex space and $f$ is $K$-convexlike (i.e., $f[\Omega]+K$ is convex). Indeed, assume that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$. Since $f$ is $K$-convexlike, we infer that $A=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is a convex cone in $Y$. By [27, Lem. 3.3 (4)] we know that $f(\bar{x})\notin\operatorname{int}(f[\Omega]+K)$. Since the interior and the quasi interior of a solid, convex set $f[\Omega]+K$ in a real Hausdorff locally convex space coincide (see [23, p. 66-67]), we have $A\subseteq{\mathrm{c}l}(\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x})))\neq Y$. Moreover, since $K\neq\{0\}$ we have $A\neq\{0\}$ as well, hence $A$ is nontrivial.
5.3 Scalarization results for the concept of proper efficiency with respect to a cone-valued map
We are ready to present some further main scalarization results which ensure that an $\mathbb{A}$-properly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
Theorem 5.25.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})$ is a nontrivial cone in $Y$ with $f[\Omega]-f(\bar{x})\subseteq A$. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively. Suppose that one of the sets $\operatorname{cl}\,S_{-K}$ and $\operatorname{cl}\,S_{A}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{A}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$ (5.2)
Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.12 is valid.
Proof 5.26.
Under our assumptions, by Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that
$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$
Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have
$(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$
hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
Remark 5.27.
In view of Remark 5.18, if $K$ is solid or $\psi$ is continuous, then the point $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.25 belongs to ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
Using Proposition 4.1 we are able to replace the separation condition (5.2) by $0\notin\operatorname{cl}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ as well as some more specific conditions that we formalize below.
Corollary 5.28.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, and that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact. Moreover, assume that one of the following statements is valid:
  1. a)
    $A$ is closed and convex.
  2. b)
    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid, and one of the conditions (4.5)–(4.8) holds.
Then, the conclusion of Theorem 5.25 is valid.
Proof 5.29.
Using Proposition 4.1 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (5.2). Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$. Then, the result is a direct consequence of Theorem 5.25.
In the remaining part of the subsection, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.
Corollary 5.30.
Assume that ($Y,||\cdot||$) is a real normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Moreover, assume that one of the following statements is valid:
  1. a)
    $B_{A}(||\cdot||)$ and ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ are compact.
  2. b)
    $A$ is weakly closed, and either $Y$ is reflexive or ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ and $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ are weakly compact and $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
Then, there is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid.
Proof 5.31.
Note that $\varphi_{x^{*},\alpha}$ from the conclusion is constructed with respect to the equivalent norm $||\cdot||^{\prime}$. Since $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ and $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$, using Proposition 4.11 a) and b) respectively, under our standing assumptions a) and b), one gets that $({\mathrm{c}l}\,S_{A}^{0}(||\cdot||^{\prime}))\cap({\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime}))=\emptyset$ for some norm $||\cdot||^{\prime}\sim||\cdot||$. Applying Lemma 4.9 ($3^{\circ}$, $4^{\circ}$), it follows that ${\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime})$ is compact under a), respectively, weakly compact under b). Consequently, the conclusion follows by applying Theorem 5.25 for $Y$ endowed with the norm $\psi(\cdot):=||\cdot||^{\prime}$ under a), respectively, endowed with the weak topology and $\psi(\cdot):=||\cdot||^{\prime}$ under b).
For the finite dimensional case, we obtain from Corollary 5.30 the following result.
Corollary 5.32.
Assume that ($Y,||\cdot||$) is a real finite dimensional normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, pointed, convex cone, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a closed cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
Then, the conclusion of Corollary 5.30 holds.
Proof 5.33.
By the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]) we know that $K^{\#}\neq\emptyset$, and by [29, Th. 3.1 ($4^{\circ}$)] it follows that $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$. So, the assumptions of Corollary 5.30 are satisfied.
Remark 5.34.
Besides the proper efficiency concepts in the senses of Benson, Borwein and Hurwicz, we gain also scalarization results for the proper efficiency concept in the sense of Henig taking into account the known relationships between proper efficiency concepts in vector optimization. In order to emphasize this fact, consider a nontrivial cone $K\subseteq Y$ with normlike-base $B_{K}(\psi)$ and suppose that either $K$ is solid or $\psi$ is continuous. Assume that ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)\subseteq{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$. Taking into account Remark 5.27, in this subsection, we have provided sufficient conditions for the equivalence of the following assertions:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
  • $3^{\circ}$.
    There is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12 is valid),
where in a normed setting the third assertion is replaced by
  • $3^{\circ}$’.
    There is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid).
Remark 5.35.
Under the assumptions in Corollary 5.28 a) (in particular the convexity of $A$), we were able to replace the separation condition (5.2) by $0\notin{\mathrm{c}l}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ (which implies the condition $A\cap(-K)=\{0\}$) in our scalarization results. In contrast, the scalarization result (for Henig and Benson proper efficiency concepts) by García-Castaño, Melguizo-Padial and Parzanese [16, Cor. 3.10] in real normed spaces still contains a separation condition (SSP), also under convexity of $A$ (ensured by a convexlikness assumption on $f$).
5.4 Scalarization results for the concept of proper efficiency in the sense of Henig
Let us end the paper with stating some further scalarization results for the concept of proper efficiency in the sense of Henig by making special use of the structure of the dilating cones $D\in\mathcal{D}(K)$.
Lemma 5.36.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, and $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$ for some $D\in\mathcal{D}(K)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D.$
Proof 5.37.
First, note that $\emptyset\neq K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ for all $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ taking into account the continuity of $\psi$. By Proposition 4.15, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$x^{*}(d)+\alpha\psi(d)\geq 0>x^{*}(-k)+\alpha\psi(k)\quad\mbox{for all }d\in\bar{D}\mbox{ and all }k\in K\setminus\{0\}.$ (5.3)
Any $x\in C^{>}_{\psi}(x^{*},\alpha)$ satisfies $0>x^{*}(-x)+\alpha\psi(x)$, hence $-x\notin\bar{D}$ in view of (5.3), and so $x\in\operatorname{int}\,D$.
Theorem 5.38.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)$. Suppose that $D$ is closed. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, the conclusion of Theorem 5.25 holds.
Proof 5.39.
By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D.$
Since $D$ is closed, we have $C_{\psi}(x^{*},\alpha)\subseteq D$, hence $C_{\psi}(x^{*},\alpha)\setminus\{0\}\subseteq D\setminus\{0\}$. We conclude that
$\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)\subseteq{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha)).$
Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
Theorem 5.40.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)$. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
Proof 5.41.
By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D,$
hence $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)\subseteq{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Proposition 5.14 finishes the proof.
Remark 5.42.
Proposition 4.1 gives sufficient conditions for the equality $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset$, which is given as an assumption in Theorems 5.38 and 5.40. Suppose that there is $(x^{*},\alpha)\in Y^{*}\times\mathbb{P}$ such that (4.4) is valid. Notice that the set $\bar{D}$ is a nontrivial, closed cone in $Y$ with $\bar{D}\cap(-K)=\{0\}$. Assume that the nontrivial, convex cone $K$ is closed and that either ${\mathrm{c}onv}({\mathrm{c}l}(B_{\bar{D}})\cup\{0\})$ is closed or ${\mathrm{c}l}\,B_{\bar{D}}$ is compact. Then, Proposition 4.1 ensures $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Remark 5.43.
The approach in this Section 5.4 follows basically some ideas by Kasimbeyli [45, Th. 4.4 and 5.7], who deals (in a normed setting) with $\varepsilon$-conic neighborhoods for $K$ (as dilating cones) and supposes a separation property for a whole family of dilating cones.
In Section 5.3 one has a separation condition for $A=\mathbb{A}(\bar{x})$ and $-K$ (for a given $\bar{x}$), and one can further ensure that $A$ is convex. Here in Section 5.4 we have a separation condition for $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ and $-K$. One has to notice that the complement cone $\bar{D}$ is usually not convex if $D$ is convex. Moreover, the cone $\mathbb{A}(\bar{x})$, which has a particular structure for specific choices of $\mathbb{A}$, is more easy to handle than the cone $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ based on a general dilating cone $D\in\mathcal{D}(K)$.
6 Conclusions
It is well-known that several scalarizing functions used in scalarization methods in vector optimization and in the treatment of optimization problems under uncertainty admit certain cone representation and cone monotonicity properties (e.g., methods in the sense of Gerstewitz [20, 21]; Kasimbeyli [45, 46]; Pascoletti and Serafini [53]; Zaffaroni [58]). In the general framework of real topological-linear spaces, our paper provides new insights into the relationships between a vector optimization problem based on an original (convex) cone, an updated vector optimization problem based on an enlarging/dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions. Thus, our new scalarization results with respect to the concepts of weak efficiency and different types of proper efficiency (e.g., in the sense of Benson; Borwein; Henig; Hurwicz) are also useful for several established scalarization methods in vector optimization (see also Remark 5.17).
In future research, we like to derive Bishop-Phelps type scalarization results for approximate solutions of vector optimization problems in real topological-linear spaces, which would extend recent scalarization results obtained by García-Castaño, Melguizo-Padial and Parzanese [16]. Moreover, the derivation of duality statements for such vector optimization problems based on Bishop-Phelps type scalarization is of interest (see also Kasimbeyli and Karimi [47]).

### v2

@@
 Abstract
-Keywords: Vector optimization; efficiency; weak efficiency; proper efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
+Keywords: Vector optimization; (proper, weak) efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
 Mathematics subject classifications (MSC 2010): 90C48, 90C29, 06F20, 52A05
@@
 The parameterized scalarization problems (in the sense of Kasimbeyli [45]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]) are formulated using Bishop-Phelps type cone representing (seminorm-linear) functions.
-In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector optimization problem (based on the convex cone $C$), an updated vector optimization problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions).
-Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorem 5.25 for proper efficiency w.r.t.
-a cone-valued map, Theorems 5.38, 5.40 for Henig proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
+In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector problem (based on the convex cone $C$), an updated vector problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions).
+Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorems 5.25, 5.38 and 5.40 for proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
 Finally, in Section 6 we present a conclusion and an outlook for further research.
@@
 Moreover, one has
-$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\quad\text{and}\quad(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
+$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\text{ and }(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
 but the inclusion $(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}$ can be strict (see Göpfert et al. [23, p. 55]).
@@
 Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$.
-For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by
-$C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}.$
+For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by $C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}$.
 It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [36, 37], Phelps [55]).
 Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone.
-If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and
-$C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset.$
+If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and $C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$.
 Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$.
@@
 Definition 2.7
-Consider a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ and a cone $K\subseteq Y$. Define a set
+Consider a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ and a cone $K\subseteq Y$. Define
 $B_{K}:=B_{K}(\psi):=\{x\in K\mid\psi(x)=1\}=K\cap\operatorname{\mathbb{S}}_{\psi}$.
@@
 Lemma 2.11
-Consider some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. Then, we have:
+Consider $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. The following assertions hold:
   • $1^{\circ}$.
@@
   • $5^{\circ}$.
-    Nontrivialness: If $(x^{*},\alpha)\in K^{a\#}$, then $C_{\psi}(x^{*},\alpha)\neq\{0\}$. If $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)\neq Y$.
+    Nontrivialness: If $(x^{*},\alpha)\in K^{a+}$, then $C_{\psi}(x^{*},\alpha)\neq\{0\}$. If $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)\neq Y$.
   • $6^{\circ}$.
@@
 In view of $1^{\circ}$, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing.
-$3^{\circ}$ This assertion is a direct consequence of Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
+$3^{\circ}$ This assertion follows from Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
 $4^{\circ}$ Of course, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing by $2^{\circ}$.
@@
 According to [37, Ex.
-2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has
-    $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$
-    and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
+2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$ and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
 Next, we give an example with $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$.
@@
 Proposition 2.15.
-    Consider a closed, pointed cone $C\subseteq Y$ in a real infinite dimensional normed space $(Y,||\cdot||$). Then, the weak interior of $C$ is empty, i.e., $\operatorname{int}_{w}\,C=\emptyset$.
+Consider a closed, pointed cone $C\subseteq Y$ in a real infinite dimensional normed space $(Y,||\cdot||$). Then, we have $\operatorname{int}_{w}\,C=\emptyset$.
 Proof 2.16.
@@
 Since $Y$ has infinite dimension, the set $U(c)$ contains a line $c+\mathbb{R}\cdot v$ for some $v\in Y\setminus\{0\}$.
-This implies, for any $n\in\mathbb{N}$,
-    $\frac{c}{n}\pm v=\frac{1}{n}(c\pm nv)\in C.$
+This implies, for any $n\in\mathbb{N}$, we have $cn^{-1}\pm v=n^{-1}(c\pm nv)\in C$.
 Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction.
@@
 For notational convenience, we consider further the binary relations $\lneq_{K}$ and $<_{K}$ that are defined, for any two points $y,\overline{y}\in Y$, by $y\lneq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K\setminus\{0\},y<_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-\operatorname{int}\,K$, respectively.
-Two well-known types of solutions to the vector problem ($P$) can be defined according to the next two definitions (see, e.g., Boţ, Grad and Wanka [5], Göpfert et al.
-[23], Jahn [42], and Khan, Tammer and Zălinescu [49]).
+Two well-known types of solutions to the vector problem ($P$) can be defined according to the next two definitions (see, e.g., [5], [23], [42], [49]).
 Definition 2.17 (Efficiency).
 A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$.
-The set of all Pareto efficient solutions of ($P$) is denoted by
-${\mathrm{E}ff}(\Omega\mid f,K):=\{\overline{x}\in\Omega\mid\not\exists\,x\in\Omega:\;f(x)\lneq_{K}f(\overline{x})\}.$
+The set of all Pareto efficient solutions of ($P$) is denoted by ${\mathrm{E}ff}(\Omega\mid f,K)$.
 Definition 2.18 (Weak efficiency).
 A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$.
-The set of all weakly Pareto efficient solutions of ($P$) is denoted by
-${\mathrm{W}Eff}(\Omega\mid f,K):=\{\overline{x}\in\Omega\mid\not\exists\,x\in\Omega:\;f(x)<_{K}f(\overline{x})\}.$
+The set of all weakly Pareto efficient solutions of ($P$) is denoted by ${\mathrm{W}Eff}(\Omega\mid f,K)$.
 In addition to the efficiency concepts above, we also want to consider a proper efficiency with respect to a cone-valued map.
@@
 Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$.
-A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and
-$\mathbb{A}(\bar{x})\cap(-K)=\{0\}.$
+A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and $\mathbb{A}(\bar{x})\cap(-K)=\{0\}$.
 The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
@@
 By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get some knowledge about the original vector problem ($P$).
-Applying such a strategy is called scalarization method in the literature of vector optimization.
-Within such methods, the function $\varphi$ is called scalarizing function.
-For more details, we also refer the reader to the standard books of vector and set optimization by Boţ, Grad and Wanka [5], Göpfert et al.
-[23], Jahn [42], and Khan, Tammer and Zălinescu [49].
+The application of such a strategy is called scalarization method, while the function $\varphi$ is called scalarizing function (see e.g.
+[5], [23], [42], [49]).
 In this paper, we like to analyze the relationships between solutions of ($P_{\varphi}$) and the (weakly, properly) efficient solutions of ($P$) for the case that $\varphi$ satisfies certain monotonicity properties.
@@
 The proofs of $2^{\circ}$ and $3^{\circ}$ are straightforward.
-The following scalarization results involving cone-monotone functions are well-known.
+The next scalarization results involving cone-monotone functions are well-known.
 Lemma 3.4 (c.f. [50, Lem. 5.6], [27, Lem. 5.2]).
@@
 Lemma 3.5 (cf. [43, Th. 4.1]).
-Suppose that $K\subseteq Y$ is a cone. Assume that $\varphi:Y\to\mathbb{R}$ is a function with $\varphi(0)=0$. Then, for any $\bar{x}\in\Omega$, the following assertions hold:
+Suppose that $K\subseteq Y$ is a cone. Assume that $\varphi:Y\to\mathbb{R}$ is a function with $\varphi(0)=0$. Then, for any $\bar{x}\in\Omega$, we have:
   • $1^{\circ}$.
@@
 Lemma 3.8.
-Suppose that $K\subseteq Y$ is a convex cone. Then, the following assertions are valid:
+Suppose that $K\subseteq Y$ is a convex cone. Then, we have:
   • $1^{\circ}$.
@@
   • $2^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the scalar optimization problem (3.1).
+    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of (3.1).
 If, in addition, $K$ is closed, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
   • $3^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and
-    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
+    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
 Proof 3.12.
-The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 3.8 ($3^{\circ}$).
+The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is a direct consequence of Lemma 3.8 ($3^{\circ}$).
 Now, we show the implication $1^{\circ}\Longrightarrow 2^{\circ}$.
@@
   • $2^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of the scalar optimization problem (3.1).
+    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of (3.1).
 If, in addition, $K$ is closed and nontrivial, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
   • $3^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that
-    $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
+    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
 Proof 3.15.
@@
 Applying assertion $2^{\circ}$ for $C$ in the role of $K$ we get the result.
-$4^{\circ}$. The first equality is obvious, while the second equality follows immediately by $1^{\circ}$.
+$4^{\circ}$. The first equality is obvious, while the second equality follows by $1^{\circ}$.
 Remark 3.19.
@@
 (e.g., if $Y$ has dimension two or higher, and $\psi$ is a norm), then
-$\operatorname{cl}\,S_{C}(\psi)=S_{C}(\psi)$ =$=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}.$
+$\operatorname{cl}\,S_{C}(\psi)=S_{C}(\psi)=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}.$
 Proof 4.5.
@@
 Case 1: Assume that $\psi(d)=1$. Then, $d\in B_{C}\subseteq S_{C}$.
-Case 2: Assume that $\psi(d)<1$. Put $\gamma:=x^{*}(d)\,(\geq\alpha)$. By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$. Then,
-$x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma\quad\text{ for all }t\geq 0$
-and
-$\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty\quad\text{ for }t\to\infty.$
+Case 2: Assume that $\psi(d)<1$.
+Put $\gamma:=x^{*}(d)\,(\geq\alpha)$.
+By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$.
+Then, $x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma$ for all $t\geq 0$, and $\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty$ for $t\to\infty$.
 Taking into account the continuity of $t\mapsto h_{+}(t):=\psi(d+te)$ and $t\mapsto h_{-}(t):=\psi(d-te)$, as well as the conditions $h_{+}(0)=h_{-}(0)=\psi(d)<1$, $\lim_{t\to\infty}\,h_{+}(t)=\lim_{t\to\infty}\,h_{-}(t)=\infty$, by the well-known intermediate value theorem there are $t_{+},t_{-}\in\operatorname{\mathbb{P}}$ with $h_{+}(t_{+})=h_{-}(t_{-})=1$.
@@
 $\Omega_{2}\subseteq[0,t]\cdot\Omega_{1}.$ (4.10)
-On the contrary, assume that there are sequences $(x_{n}^{1})$, $(x_{n}^{2})$ and $(t_{n})$ with $x_{n}^{1}\in\Omega_{1}$, $x_{n}^{2}\in\Omega_{2}$, $t_{n}\in\mathbb{P}$ such that $x_{n}^{2}=t_{n}\cdot x_{n}^{1}$ with $t_{n}\to\infty$.
+On the contrary, assume that there are sequences $(x_{n}^{1})$, $(x_{n}^{2})$ and $(t_{n})$ with $x_{n}^{1}\in\Omega_{1}$, $x_{n}^{2}\in\Omega_{2}$, $t_{n}\in\mathbb{P}$ such that $x_{n}^{2}=t_{n}\cdot x_{n}^{1}$ with $t_{n}\to\infty$ for $n\to\infty$.
 By the boundedness of $\Omega_{2}$ there is $\lambda>0$ such that $||x||\leq\lambda$ for all $x\in\Omega_{2}$.
-Consequently, we have
-$||x_{n}^{1}||=||t_{n}^{-1}x_{n}^{2}||=t_{n}^{-1}||x_{n}^{2}||\leq t_{n}^{-1}\lambda\to 0\quad\text{for }t_{n}\to\infty.$
+Consequently, we have $||x_{n}^{1}||=||t_{n}^{-1}x_{n}^{2}||=t_{n}^{-1}||x_{n}^{2}||\leq t_{n}^{-1}\lambda\to 0$ for $n\to\infty$.
 This means that $0\in\operatorname{cl}\,\Omega_{1}\subseteq\operatorname{cl}_{w}\,\Omega_{1}$.
@@
 Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence both weakly closed sets $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact if $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact.
-To later prove a Bishop-Phelps type scalarization result (namely Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
+To later prove a Bishop-Phelps type scalarization result (Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
 Proposition 4.11.
@@
 Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$.
-Then, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, and any given reference point $a\in Y$, one defines $\varphi_{x^{*},\alpha}^{a}:Y\to\operatorname{\mathbb{R}}$ as
-$\varphi_{x^{*},\alpha}^{a}(y):=\varphi_{x^{*},\alpha}(y-a)\quad\text{for all }y\in Y.$
+Then, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, and any given reference point $a\in Y$, one defines $\varphi_{x^{*},\alpha}^{a}:Y\to\operatorname{\mathbb{R}}$ as $\varphi_{x^{*},\alpha}^{a}(y):=\varphi_{x^{*},\alpha}(y-a)$ for all $y\in Y$.
 The following parameterized scalar optimization problem of the vector problem ($P$) (in the sense of Kasimbeyli [45]) will be of interest:
@@
 Lemma 5.3.
-Consider some $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
+Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
   • $1^{\circ}$.
@@
 $2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid.
-Now assume that $\operatorname{int}\,K\neq\emptyset$.
-If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so the function $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 5.1 ($2^{\circ}$) and 3.2 ($2^{\circ}$).
+Assume that $\operatorname{int}\,K\neq\emptyset$.
+If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 5.1 ($2^{\circ}$) and 3.2 ($2^{\circ}$).
 The proof of the reverse implication is similar to $1^{\circ}$.
@@
 Lemma 5.8.
-Consider some $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions are valid:
+Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions are valid:
   • $1^{\circ}$.
@@
 For the case $\operatorname{int}\,K=\emptyset$ the statement is clear.
-Now assume that $\operatorname{int}\,K\neq\emptyset$.
+Assume that $\operatorname{int}\,K\neq\emptyset$.
 Again we have $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
@@
     $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
-If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then the assertions $1^{\circ}-3^{\circ}$ are equivalent to:
-If further $C_{\psi}(x^{*},\alpha)$ is closed, then the assertions $1^{\circ}-4^{\circ}$ are equivalent to:
+If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}-3^{\circ}$ are equivalent to:
+If further $C_{\psi}(x^{*},\alpha)$ is closed, then $1^{\circ}-4^{\circ}$ are equivalent to:
   • $5^{\circ}$.
-    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and
-    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
+    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
 Proof 5.13.
@@
     $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
-If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then the assertions $1^{\circ}$ and $2^{\circ}$ are equivalent to:
-If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then the assertions $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
+If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
+If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
   • $5^{\circ}$.
-    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that
-    $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
+    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
 Proof 5.15.
@@
   • Kasimbeyli [46, Th. 5.4] stated similar assertions as in Theorem 5.5 for a closed, pointed, convex cone $K$ and Benson proper efficiency (instead of Henig proper efficiency).
-  • Gasimov [17, Th. 2] (see also Guo and Zhang [31, Th. 4.5]) stated
-    $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$
-    $\Longrightarrow\{x\in\Omega$ $\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$
-    with the cone-valued map $\mathbb{A}$ considered in Benson proper efficiency.
+  • Gasimov [17, Th.
+2] (see also Guo and Zhang [31, Th.
+4.5]) stated that $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$ (with $\mathbb{A}$ considered in Benson proper efficiency) implies $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
 Remark 5.17.
@@
 Remark 5.18.
-Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$. Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence
-$\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{E}ff}(\Omega\mid f,K).$
+Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$.
+Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
 If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g.
-if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so (see Proposition 2.23)
-$\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K).$
-If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence
-$\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{W}Eff}(\Omega\mid f,K).$
+if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so (see Proposition 2.23) $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
+If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence $\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
 5.2 Scalarization results for the concept of weak efficiency
@@
 Proof 5.20.
-Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that
-$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha).$
-Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have
-$(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$
-hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that $A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
+Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$ hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
 Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 5.14 yields the rest of the conclusion.
@@
 Proof 5.26.
-Under our assumptions, by Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that
-$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$
-Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have
-$(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$
-hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
+By Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $A\cap-C_{\psi}(x^{*},\alpha)=\{0\}$ and $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$.
+Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$ hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+Finally, observing that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
 Remark 5.27.
@@
 Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$.
-Then, the result is a direct consequence of Theorem 5.25.
-In the remaining part of the subsection, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.
+Then, the result follows from Theorem 5.25.
+In the remaining part of the section, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.
 Corollary 5.30.

### v3

@@
-disposition
+
 Abstract
@@
 Moreover, we study some extensions of known scalarization results in vector optimization (in the sense of Eichfelder; Gerstewitz; Jahn; Kasimbeyli; Pascoletti and Serafini).
-On this basis, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s conic scalarization method in real normed spaces.
+On this basis, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which is based on Bishop-Phelps type cone representing and cone monotone scalarizing functions (e.g., Gerstewitz scalarizing functions or seminorm-linear scalarizing functions).
+Thus, our method also extends Kasimbeyli’s conic scalarization method from real normed spaces to real topological-linear spaces.
 Within this framework, we derive new Bishop-Phelps type scalarization results for the concepts of weak efficiency and different types of proper efficiency.
-Abstract
-Keywords: Vector optimization; (proper, weak) efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
-Mathematics subject classifications (MSC 2010): 90C48, 90C29, 06F20, 52A05
+keywords:
+Vector optimization, (proper, weak) efficiency, Bishop-Phelps cone, augmented dual cone, Bishop-Phelps type scalarization, conic scalarization
+pacs:
+MSC Classification]90C48, 90C29, 06F20, 52A05
 1 Introduction
 The field of vector optimization in real topological-linear spaces is concerned with the minimization of a vector-valued function $f:X\to Y$ acting between real topological-linear spaces $X$ and $Y$ with respect to a convex cone $C\subseteq Y$ over a nonempty feasible set $\Omega\subseteq X$.
-A review of the existing literature reveals that many papers focus on the (finite dimensional) normed setting (i.e., $Y$ is a normed space), see e.g.
-[16], [8, 9], [12, 13], [17], [32], [45, 46], [53], [58], [57, Ch.
+A review of the existing literature reveals that many papers focus on the (finite dimensional) normed setting and, in particular, the analysis includes real normed spaces $Y$ where the convex cone $C$ is of Bishop-Phelps type, see e.g.
+[1], [2, 3], [4, 5], [6], [7], [8, 9], [10], [11], [12, Ch.
 7 and Sec.
-15.2], and references therein.
-In particular, the analysis includes real normed spaces $Y$ where the convex cone $C$ is of Bishop-Phelps type (named after Bishop and Phelps [2], see also [16], [15], [28, 29], [36, 37], [35],[41, 43], [45, 46]).
-The importance of the class of Bishop-Phelps cones can be emphasized by the fact that a nontrivial, closed, convex cone in a real normed space is representable as a Bishop-Phelps cone if and only if it is well-based (see Example 2.4 for details).
+15.2], [13], [14, 15], [16, 17], [18],[19, 20], and references therein.
+The importance of the class of Bishop-Phelps cones can be emphasized by the fact that a nontrivial, closed, convex cone in a real normed space is representable as a Bishop-Phelps cone if and only if it is well-based.
 We are primarily interested on the more general setting of a real (possibly infinite dimensional and not normed) topological-linear space $Y$ motivated by optimization problems under uncertainty and problems from mathematical finance.
 Especially, robust counterparts to optimization problems under uncertainty can be formulated as vector optimization problems with an infinite dimensional topological-linear image space (see Göpfert et al.
-[23, Ch.
+[21, Ch.
 6] and references therein).
@@
 Some further works in vector optimization that focus on this setting are [25], [26, 27], [28], [21], [29], [30], [31], [32, 20], [12].
-The main focus in our paper is on the computation of efficient / weakly efficient / properly efficient solutions of the given vector optimization problem via scalarization (based on Bishop-Phelps type cones), where the proper efficiency concepts considered in this paper cover certain known concepts of proper efficiency from the literature (e.g., in the sense of Benson [1]; Borwein [3, 4]; Henig [38]; Hurwicz [40]).
+The main focus in our paper is on the derivation of scalarization results for (weakly, properly) efficient solutions of the given vector optimization problem, where the proper efficiency concepts considered in this paper cover certain known concepts of proper efficiency from the literature (e.g., in the sense of Benson [33]; Borwein [34, 35]; Henig [36]; Hurwicz [37]).
 It is well established that scalarization methods are useful for solving vector optimization problems.
@@
 The application of the scalarization methods for Henig dilating cones (enlarging convex cones) of the convex cone $C$ yields solutions that are properly efficient in the sense of Henig [36].
-The relationships of the Henig proper efficiency concept to other known proper efficiency concepts have been well studied in the literature (see, e.g., [5, Sec.
-2.4.3], [25, Ch.
-3], [26], [49, Sec.
+The relationships of the Henig proper efficiency concept to other known proper efficiency concepts have been well studied in the literature (see, e.g., [25, Sec.
+2.4.3], [31, Ch.
+3], [41], [42, Sec.
 2.4] and references therein).
-We will now recall some well-known scalarization methods in vector optimization.
-Of course, the linear scalarization method in vector optimization (respectively, the weighted sum method in multiobjective optimization) is the best known technique for generating weakly / properly efficient solutions of vector (multiobjective) optimization problems.
-It is known that this method only works well under cone-convexity type assumptions (see, e.g., [27]).
-For nonconvex vector optimization problems, there are some other scalarization methods frequently used in the literature that we recall below.
-In the framework of real (possibly infinite dimensional) normed spaces, the conic scalarization method by Kasimbeyli [45, 46] (see also [16]), the Drummond–Svaiter scalarization method [7] (see also [6]) and the Zaffaroni scalarization method [58] (see also [14, Sec.
-5.1]) based on the Hiriart-Urruty function [39] (also known as oriented/signed distance function) are known to be useful, where the scalarizing function of the first two methods involves the underlying norm of the space $Y$.
-Considering the more general framework of real topological-linear spaces, only a few methods (besides the linear scalarization method) are known to be valid, such as the Gerstewitz scalarization method [20, 21] (see also [22], [49], [57], [23]) and the Pascoletti–Serafini scalarization method [53] (see also [12, 13], [57, Sec.
-6.2.2]).
-Certain well-known scalarization methods in vector and multiobjective optimization (e.g.
-linear scalarization method / weighted sum method, $\varepsilon$-constraint method, hybrid method, Benson method, weighted Chebyshev norm method, and other methods; see [57, Sec.
-7.3]) are special cases of these general scalarization methods.
-Some further comparison studies of scalarization methods in vector optimization are given in [6], [48], [12, Ch.
-2], while generalizations of scalarization methods to vector and set optimization involving variable domination structures are also studied in the recent literature (see, e.g., [35], [15], [14], [43], [57, Ch.
-10 and 11] and references therein).
-As mentioned above, one recognized approach in vector optimization in a real normed space $(Y,||\cdot||)$ is the conic scalarization method in vector optimization proposed by Kasimbeyli [45], which is based on the following parameterized scalar optimization problem:
-$\begin{cases}x^{*}(f(x)-a)+\alpha||f(x)-a||\to\min\\ x\in\Omega\end{cases}$ (1.1)
-with a pair $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ (respectively, $(x^{*},\alpha)$ belongs to the so-called augmented dual cone of $C$; see Section 2.3 for details) and a given reference point $a\in Y$.
-Several relationships between (weakly, properly) efficient solutions of the original vector problem and solutions of scalar problems of type (1.1) are known (see [16], [17], [31], [45, 46]).
-In this paper, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s scalarization method in real normed spaces.
-By taking a look on (1.1) one can see that Kasimbeyli’s method is based on Bishop-Phelps type norm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of the norm |||||$\cdot$||| | ⋅ | | with a linear continuous functional $x^{*}$, see also Zaffaroni [59]), where the involving norm |||||$\cdot$||| | ⋅ | | is exactly the underlying norm in the real normed space $Y$.
-In our Bishop-Phelps type scalarization method, we consider Bishop-Phelps type seminorm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of a seminorm $\psi$ with a linear continuous functional $x^{*}$), where the seminorm involved (namely $\psi$) must not have anything to do with the underlying topology of the real (locally convex) topological-linear space (e.g., the seminorm must not be part of a family of seminorms that generates this topology).
-The rest of the paper is organized as follows:
-In Section 2, after presenting some basics in real topological-linear spaces including properties of seminorms, we study properties of cones and their bases, focusing in particular on augmented dual cones, Bishop-Phelps type cones and zero level sets of Bishop-Phelps type (seminorm-linear) functions.
+One recognized approach in vector optimization in a real normed space $(Y,||\cdot||)$ is the conic scalarization method in vector optimization proposed by Kasimbeyli [8], which is based on the following parameterized scalar optimization problem:
+$\begin{cases}x^{*}(f(x)-a)+\alpha||f(x)-a||\to\min\\ x\in\Omega\end{cases}$ (1)
+with a pair $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ (respectively, $(x^{*},\alpha)$ belongs to the so-called augmented dual cone of $C$) and a given point $a\in Y$.
+By taking a look on (1) one can see that Kasimbeyli’s method is based on Bishop-Phelps type norm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of the norm |||||$\cdot$||| | ⋅ | | with a linear continuous functional $x^{*}$, see also Zaffaroni [43]), where the involving norm |||||$\cdot$||| | ⋅ | | is exactly the underlying norm in the real normed space $Y$.
+Several relationships between (weakly, properly) efficient solutions of the original vector problem and solutions of scalar problems of type (1) are known (see [1], [6], [44], [8, 9]).
+In this paper, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which is based on Bishop-Phelps type cone representing and cone monotone scalarizing functions (e.g., Gerstewitz scalarizing functions or Bishop-Phelps type seminorm-linear scalarizing functions, i.e., functions that are a nonnegative weighted sum of a seminorm with a linear continuous functional).
+The underlying scalarizations of the vector problem will be of the types (in the sense of Gerstewitz [26, 27]; Jahn [20]; Pascoletti and Serafini [10]):
+$\begin{cases}\varphi(f(x)-a)\to\min\\ x\in\Omega\end{cases}$ ($P_{\varphi}^{a}$)
+and
+$\begin{cases}\lambda\to\min\\ \varphi(f(x)-a-\lambda k)\leq 0,\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ ($P_{\varphi}^{a,k}$)
+where $a\in Y$, $k\in Y\setminus\{0\}$, and $\varphi:Y\to\mathbb{R}$ is a Bishop-Phelps (type) cone representing function, i.e., the level set
+$C_{\rm BP}:=-\{y\in Y\mid\varphi(y)\leq 0\}$
+is a Bishop-Phelps (type) cone in $Y$. In Kasimbeyli’s conic scalarization method based on the scalar problem (1), the set
+$C_{\rm BP}=-\{y\in Y\mid x^{*}(y)+\alpha||y||\leq 0\}$
+with $(x^{*},\alpha)\in(Y^{*}\setminus\{0\})\times(\mathbb{R}_{+}\setminus\{0\})$ is indeed a Bishop-Phelps cone in the real normed space $Y$.
+Hence, our method can be seen as an extension of Kasimbeyli’s scalarization method from real normed spaces to real topological-linear spaces.
+The rest of the paper is organized as follows: In Section 2, after presenting some basics in real topological-linear spaces including properties of seminorms, we study properties of cones and their bases, focusing in particular on augmented dual cones, Bishop-Phelps type cones and zero level sets of Bishop-Phelps type (seminorm-linear) functions.
 We then introduce the solution concepts of vector optimization necessary for our work.
 Section 3 is devoted to the study of scalarization methods in vector optimization which are based on cone monotone and cone representing scalarizing functions.
-Beside classical scalarization results (in the sense of Eichfelder [12, 13]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]), we establish some new scalarization results (namely Propositions 3.11, 3.14 and 3.17) for vector optimization problems in real topological-linear spaces, which extend the classical ones.
-In Section 4, we present the key tool to prove our main Bishop-Phelps type scalarization results in vector optimization, namely the technique of nonlinear separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions.
+Beside classical scalarization results (in the sense of Eichfelder [4, 5]; Gerstewitz [26, 27]; Jahn [20]; Pascoletti and Serafini [10]), we establish some new scalarization results (namely Propositions 9, 10 and 11) for vector optimization problems in real topological-linear spaces, which extend the classical ones.
+The main Section 4 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces.
+The parameterized scalarization problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) are formulated using Bishop-Phelps type cone representing and cone monotone scalarizing functions (e.g., Gerstewitz [26, 27] scalarizing functions or seminorm-linear scalarizing functions).
+In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 17 and for the concept of weak efficiency in Proposition 18) the relationships between the original vector problem (based on the convex cone $C$), an updated vector problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar problems involving Bishop-Phelps type cone representing and cone monotone scalarizing functions.
+In Section 4.2, we present the key tool to prove our main Bishop-Phelps type scalarization results in vector optimization, namely the technique of nonlinear separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions.
 Besides presenting some classical nonlinear cone separation results, we also reflect on the separation conditions involved in both real topological-linear spaces and real normed spaces.
-Propositions 4.1 and 4.11 give some new insights into the relationships between the classical separation condition (the intersection of two cones is the singleton set $\{0\}$) and the separation conditions used in the nonlinear separation approach based on Bishop-Phelps type cones / functions discussed in [16], [28, 29], [45, 46].
-We are able to deduce some new variants of nonlinear cone separation theorems involving the classical separation condition (see Corollaries 4.16 and 4.17).
-The main Section 5 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces.
-The parameterized scalarization problems (in the sense of Kasimbeyli [45]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]) are formulated using Bishop-Phelps type cone representing (seminorm-linear) functions.
-In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector problem (based on the convex cone $C$), an updated vector problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions).
-Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorems 5.25, 5.38 and 5.40 for proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
+Propositions 19 and 23 give some new insights into the relationships between the classical separation condition (the intersection of two cones is the singleton set $\{0\}$) and the separation conditions used in the nonlinear separation approach based on Bishop-Phelps type cones / functions discussed in [1], [14, 15], [8, 9].
+Then, using nonlinear cone separation results (from Section 4.2) and Bishop-Phelps type scalarization results (Propositions 17 and 18), we present new scalarization theorems (Theorem 29 for weak efficiency, Theorems 31, 36 and 37 for proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
 Finally, in Section 5 we present a conclusion and an outlook for further research.
 2 Preliminaries
-2.1 Basics in topological-linear spaces
 Throughout the paper, we assume that $Y$ is a real topological-linear space.
@@
 In certain situations we will also consider a more specific setting where $(Y,||\cdot||)$ is a real normed space.
-Consistently, the set of nonnegative real numbers is denoted by $\mathbb{R}_{+}$, while $\mathbb{P}:=\mathbb{R}_{++}$ denotes the set of positive real numbers.
-Given a set $\Omega\subseteq Y$, we denote by ${\mathrm{c}onv}\,\Omega$ the convex hull of $\Omega$.
-Moreover, we denote by ${\mathrm{c}l}\,\Omega$, ${\mathrm{b}d}\,\Omega$, and ${\mathrm{i}nt}\,\Omega$ the closure, the boundary, and the interior of $\Omega$ with respect to the underlying topology in $Y$.
-The algebraic interior (the core) of $\Omega$ is defined by $\operatorname{cor}\,\Omega:=\{x\in\Omega\mid\forall\,v\in Y\;\exists\,\varepsilon>0:\;x+[0,\varepsilon]\cdot v\subseteq\Omega\}.$ It is known that $\operatorname{int}\,\Omega\subseteq\operatorname{cor}\,\Omega\subseteq\Omega\subseteq\operatorname{cl}\,\Omega=(\operatorname{int}\,\Omega)\cup\operatorname{bd}\,\Omega,$ and if $\operatorname{int}\,\Omega\neq\emptyset$ (i.e., $\Omega$ is solid) and convex, then $\operatorname{cor}\,\Omega=\operatorname{int}\,\Omega$.
-Given a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$, we define the $\psi$-unit ball by $\operatorname{\mathbb{B}}_{\psi}:=\{y\in Y\mid\psi(y)\leq 1\}$ and the $\psi$-unit sphere by $\operatorname{\mathbb{S}}_{\psi}:=\{y\in Y\mid\psi(y)=1\}.$
-Remark 2.1
-Suppose that $Y$ is a real topological-linear space and $\psi:Y\to\mathbb{R}$ is a seminorm. The following assertions hold:
-  • $1^{\circ}$.
-    $\psi$ is upper semicontinuous $\iff$ the set $\{y\in Y\mid\psi(y)<1\}$ is open $\iff$ $\psi$ is continuous at the origin $\iff$ $\psi$ is continuous.
-  • $2^{\circ}$.
-    $\psi$ is lower semicontinuous $\iff$ $\operatorname{\mathbb{B}}_{\psi}$ is closed.
-Remark that ${\mathrm{k}er}\,\psi=\{0\}$ if and only if $\psi$ is a norm, and the singleton set $\{0\}$ is closed if and only if the topological-linear space $Y$ is Hausdorff.
-Moreover, it is known (see, e.g., Gelfand [18, 19], Eberlein [11]) that if $Y$ is a Banach space and $\psi:Y\to\mathbb{R}$ is a lower semicontinuous seminorm (alternatively, the seminorm $\psi$ has the Baire property), then $\psi$ is continuous.
-2.2 Cones and their bases
+Consistently, the set of nonnegative real numbers and the set of positive real numbers is denoted by $\mathbb{R}_{+}$ and $\mathbb{P}$, respectively.
+Given a set $\Omega\subseteq Y$, we denote by ${\rm conv}\,\Omega$ the convex hull of $\Omega$.
+Moreover, we denote by ${\rm cl}\,\Omega$, ${\rm bd}\,\Omega$, and ${\rm int}\,\Omega$ the closure, the boundary, and the interior of $\Omega$ with respect to the underlying topology in $Y$.
+The algebraic interior (the core) of $\Omega$ is defined by $\operatorname{cor}\,\Omega:=\{x\in\Omega\mid\forall\,v\in Y\;\exists\,\varepsilon>0:\;x+[0,\varepsilon]\cdot v\subseteq\Omega\}.$
+Given a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ (i.e., $\psi$ is sublinear and symmetric), we define the $\psi$-unit ball by $\operatorname{\mathbb{B}}_{\psi}:=\{y\in Y\mid\psi(y)\leq 1\}$ and the $\psi$-unit sphere by $\operatorname{\mathbb{S}}_{\psi}:=\{y\in Y\mid\psi(y)=1\}.$ In what follows, we will suppose that $\psi$ is nontrivial (i.e., $\psi\not\equiv 0$).
+It is well-known that if $Y$ is a real topological-linear space and $\psi:Y\to\mathbb{R}$ is a seminorm, then $\psi$ is upper semicontinuous $\Longleftrightarrow$ $\{y\in Y\mid\psi(y)<1\}$ is open $\Longleftrightarrow$ $\psi$ is continuous at the origin $\Longleftrightarrow$ $\psi$ is continuous.
+Moreover, $\psi$ is lower semicontinuous $\Longleftrightarrow$ $\operatorname{\mathbb{B}}_{\psi}$ is closed.
+Remark also that $\{y\in Y\mid\psi(y)=0\}=\{0\}$Longleftrightarrow $\psi$ is a norm.
+The singleton set $\{0\}$ is closed $\Longleftrightarrow$ the topological-linear space $Y$ is Hausdorff.
+Moreover, it is known (see, e.g., Gelfand [45, 46], Eberlein [47]) that if $Y$ is a Banach space and $\psi:Y\to\mathbb{R}$ is a lower semicontinuous seminorm, then $\psi$ is continuous.
 A set $K\subseteq Y$ is called a cone if $0\in K=\operatorname{\mathbb{R}}_{+}\cdot K$.
@@
 Given a convex cone $K\subseteq Y$ in the real topological-linear space $Y$, the set $\ell(K)$ is known as the lineality space of $K$.
-In this case, $K$ is a linear subspace of $Y$ if and only if $K=\ell(K)$, and moreover, $\operatorname{int}\,K=K+\operatorname{int}\,K$ and $K\setminus\ell(K)=K+(K\setminus\ell(K))$.
-If $K\subseteq Y$ is a cone, then $K\neq Y\iff 0\notin\operatorname{cor}\,K\iff 0\notin\operatorname{int}\,K,$ and if further $K$ is convex, then (see also [27, Lem.
+In this case, $K$ is a linear subspace of $Y$Longleftrightarrow $K=\ell(K)$, and moreover, $\operatorname{int}\,K=K+\operatorname{int}\,K$ and $K\setminus\ell(K)=K+(K\setminus\ell(K))$.
+If $K\subseteq Y$ is a cone, then $K\neq Y\iff 0\notin\operatorname{cor}\,K\iff 0\notin\operatorname{int}\,K,$ and if further $K$ is convex, then (see also [48, Lem.
 2.10]) $K\neq Y\iff\ell(K)\cap\operatorname{int}\,K=\emptyset.$
@@
 Moreover, one has
-$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\text{ and }(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
-but the inclusion $(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}$ can be strict (see Göpfert et al. [23, p. 55]).
-In the following definition, we recall a general base concept for cones in real topological-linear spaces (cf. Göpfert et al. [23, Def. 2.1.42]).
-Definition 2.2
+$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\text{ and }(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2)
+but the inclusion $(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}$ can be strict (see Göpfert et al. [21, p. 55]).
+In the following definition, we recall a general base concept for cones in real topological-linear spaces (cf. Göpfert et al. [21, Def. 2.1.42] for the case of a convex cone).
+Definition 1 (Base).
 Consider a nontrivial cone $K\subseteq Y$.
@@
 Moreover, $K$ is said to be well-based if there exists a bounded, convex base of $K$.
-Remark 2.3
-Taking into account (2.1), for any nontrivial cone $K\subseteq Y$ in a real normed space $Y$, it is well-known (see [23, Prop. 2.2.23 and 2.2.32]) that
-$\operatorname{conv}\,K\text{ has a convex base}\iff K^{\#}\neq\emptyset,$ (2.2)
+Remark 1.
+Taking into account (2), for any nontrivial cone $K\subseteq Y$ in a real normed space $Y$, it is well-known (see [21, Prop. 2.2.23 and 2.2.32]) that
+$\operatorname{conv}\,K\text{ has a convex base}\iff K^{\#}\neq\emptyset,$ (3)
 and
-$\operatorname{conv}\,K\text{ is well-based}\iff\operatorname{cl}(\operatorname{conv}\,K)\text{ is well-based}\iff\operatorname{int}\,K^{+}\neq\emptyset.$ (2.3)
-Moreover, if $Y$ has finite dimension and $K$ is closed, then $K^{\#}=\operatorname{int}\,K^{+}$ (see [29, Th. 2.1(4), Rem. 2.6]), hence all the conditions involved in (2.2) and (2.3) are equivalent.
-Example 2.4
+$\operatorname{conv}\,K\text{ is well-based}\iff\operatorname{cl}(\operatorname{conv}\,K)\text{ is well-based}\iff\operatorname{int}\,K^{+}\neq\emptyset.$ (4)
+Moreover, if $Y$ has finite dimension and $K$ is closed, then $K^{\#}=\operatorname{int}\,K^{+}$ (see [15, Th. 2.1(4), Rem. 2.6]), hence all the conditions involved in (3) and (4) are equivalent.
+Example 1 (Bishop-Phelps cone).
 Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$.
-For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by $C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}$.
-It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [36, 37], Phelps [55]).
+For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [49]) is defined by
+$C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}.$
+It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [16, 17], Kasimbeyli [8], Phelps [50]).
 Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone.
@@
 Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$.
-It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 4.11).
-Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 2.3), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [49, Th.
+It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 23).
+Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 1), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [42, Th.
 2.4.7]).
-Moreover, we recall another (algebraic) base concept for cones in linear spaces (cf. Göpfert et al. [23, Def. 2.1.14], Jahn [42, Def. 1.10 (d)]).
-Definition 2.5
+Further properties of Bishop-Phelps cones are studied in [51, Th.
+4.5, Lem.
+4.2], particularly another characterization for cones which are representable as a Bishop-Phelps cone (a so-called representation theorem).
+See also Lemma 20 and Remark 16.
+Moreover, we recall another (algebraic) base concept for cones in linear spaces (cf. Göpfert et al. [21, Def. 2.1.14], Jahn [32, Def. 1.10 (d)] for the case of a convex cone).
+Definition 2 (Algebraic base).
 Consider a nontrivial cone $K\subseteq Y$.
 A set $B\subseteq K\setminus\{0\}$ is called an algebraic base for $K$, if $B$ is a nonempty set, and every $x\in K\setminus\{0\}$ has a unique representation of the form $x=\lambda b$ for some $\lambda>0$ and some $b\in B$.
-Remark 2.6
-If $B$ is base in the sense of Definition 2.2 or Definition 2.5 for the nontrivial cone $K\subseteq Y$, then $K=\operatorname{\mathbb{R}}_{+}\cdot B$ and $K\setminus\{0\}=\operatorname{\mathbb{P}}\cdot B$.
+Remark 2.
+If $B$ is base in the sense of Definition 1 or Definition 2 for the nontrivial cone $K\subseteq Y$, then $K=\operatorname{\mathbb{R}}_{+}\cdot B$ and $K\setminus\{0\}=\operatorname{\mathbb{P}}\cdot B$.
 Our focus will be on a specific type of base in real topological-linear spaces, which we will recall in the following definition.
-Definition 2.7
+Definition 3 (Normlike-base).
 Consider a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ and a cone $K\subseteq Y$. Define
@@
 If $\psi$ is a norm $||\cdot||:Y\to\operatorname{\mathbb{R}}$, then $B_{K}(||\cdot||)$ is called a norm-base for $K$.
-Note that $B_{K}(\psi)$ is a base in the sense of Definition 2.5, and if $\psi$ is continuous and $K$ is closed, then it is a base also in the sense of Definition 2.2.
-2.3 Augmented dual cones and Bishop-Phelps type cones
+Note that $B_{K}(\psi)$ is a base in the sense of Definition 2, and if $\psi$ is continuous, then it is a base also in the sense of Definition 1.
+Example 2 (Normlike-bases).
+  • a)
+    Consider a nontrivial cone $K\subseteq Y$ and some $x^{\ast}\in K^{\#}$.
+Define a seminorm $\psi$ by $\psi(x):=|x^{\ast}(x)|$ for all $x\in Y$.
+Notice that for dimension two or higher $\psi$ is not a norm.
+Since $x^{\ast}\in K^{\#}$, we have $\psi(x)>0$ for all $x\in K\setminus\{0\}$, hence $B_{K}(\psi)$ is a normlike-base for $K$.
+Notice that for $x^{\ast}\in K^{\#}$, we have
+    $B_{K}(\psi)=K\cap\mathbb{S}_{\psi}=\{x\in K\mid|x^{\ast}(x)|=1\}=\{x\in K\mid x^{\ast}(x)=1\},$
+    which is a usual type of bases for convex cones.
+  • b)
+    From [14, Lem. 2.16] we have the following results:
+    Consider a function $\varphi:Y\to\mathbb{R}$, a function $\psi_{\max}:Y\to\mathbb{R}$ defined by
+    $\psi_{\max}(y):=\max\{\varphi(y),\varphi(-y)\}\quad\text{ for all }y\in Y,$ (5)
+    and assume that $K,A\subseteq Y$ are nontrivial cones. Then, the following assertions hold:
+    • $1^{\circ}$
+      If $\varphi$ is sublinear (i.e., $\varphi$ is positive homogeneous and subadditive), then $\psi_{\max}$ is a seminorm.
+    • $2^{\circ}$
+      Assume that $\varphi$ is $-K$ representing. Then, $\ell(K)=\{y\in Y\mid\psi_{\max}(y)\leq 0\}.$ Moreover, we have:
+      $\psi_{\max}$ is nonnegative on $Y$ $\iff$ $\ell(K)=\{y\in Y\mid\psi_{\max}(y)=0\}$.
+      $\psi_{\max}$ is positive on $Y\setminus\{0\}$ $\iff$ $K$ is pointed.
+    • $3^{\circ}$
+      Assume that $\varphi$ is sublinear and $-K$ representing. Then, $\psi_{\max}$ is a seminorm. Moreover, we have:
+      $\ell(K)\cap A=\{0\}\iff$ $B_{A}(\psi_{\max})$ is a normlike-base for $A$.
+    • $4^{\circ}$
+      If $K$ is pointed, $\varphi$ is sublinear and $-K$ representing, then $\psi_{\max}$ is a norm, hence $B_{K}(\psi_{\max})$ and $B_{A}(\psi_{\max})$ are norm-bases for $K$ and $A$, respectively.
+  • c)
+    Assume that $A=\bigcup_{i=1}^{l}A_{i}$ for some nontrivial cones $A_{1},\ldots,A_{l}\subseteq Y$, and consider $y_{i}^{\ast}\in(A_{i})^{\#}$ for all $i=1,\ldots,l$.
+Define a seminorm $\psi$ (which is not necessarily a norm) by
+    $\psi(x):=\max\{|x^{\ast}(x)|,|y_{1}^{\ast}(x)|,\ldots,|y_{l}^{\ast}(x)|\}\quad\text{for all }x\in Y,$
+    or by
+    $\psi(x):=|x^{\ast}(x)|+\sum_{i=1}^{l}\,|y_{i}^{\ast}(x)|\quad\text{for all }x\in Y.$
+    Then, $\psi(x)>0$ for all $x\in(K\cup A)\setminus\{0\}$, hence $B_{K}(\psi)$ and $B_{A}(\psi)$ are normlike-bases for $K$ and $A$.
 Assume that $Y$ is a topological-linear space, $K\subseteq Y$ is a nontrivial cone, and $\psi:Y\to\mathbb{R}$ is a seminorm.
 For any $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, we consider the Bishop-Phelps type (seminorm-linear) function $\varphi_{x^{*},\alpha}:Y\to\mathbb{R}$ defined by
-$\varphi_{x^{*},\alpha}(y):=x^{*}(y)+\alpha\psi(y)\quad\text{for all }y\in Y,$
+$\varphi_{x^{*},\alpha}(y):=x^{*}(y)+\alpha\psi(y)\quad\text{for all }y\in Y,$ (6)
 as used in [14].
-Note that Bishop-Phelps type (norm-linear) functions in a normed setting were studied for instance by García-Castaño, Melguizo-Padial and Parzanese [16], Gasimov (Kasimbeyli) [17], Günther, Khazayel and Tammer [29], Guo and Hang [31], Ha [35], Jahn [43], Kasimbeyli [45, 46], and Zaffaroni [59].
+Note that the seminorm $\psi$ must not have anything to do with the underlying topology of the real (locally convex) topological-linear space.
+Several examples of seminorms $\psi$, which are not necessarily norms, that can be used in the definition of the function $\varphi_{x^{*},\alpha}$ in (6) as well as in the subsequent study presented below, are discussed in [14, Sec.
+2].
+Bishop-Phelps type (norm-linear) functions in a normed setting (i.e., $\psi$ is given by the underlying norm of the space) were studied for instance by García-Castaño, Melguizo-Padial and Parzanese [1], Gasimov (Kasimbeyli) [6], Günther, Khazayel and Tammer [15], Guo and Hang [44], Ha [18], Jahn [20], Kasimbeyli [8, 9], and Zaffaroni [43].
 Following the definitions of so-called augmented dual cones by Gasimov [6], Kasimbeyli [8] (see also [14] for some generalizations), we consider a (generalized) augmented dual cone by
@@
 If no danger of confusion arises, we will use the notations $K^{a+}$, $K^{a\circ}$ and $K^{a\#}$.
-Remark 2.8
+Remark 3.
 Note that $(\operatorname{conv}\,K)^{a+}=K^{a+}$ and $(\operatorname{conv}\,K)^{a\#}=K^{a\#}$, and if $\psi$ is lower semicontinuous, then $(\operatorname{cl}(\operatorname{conv}\,K))^{a+}=K^{a+}$.
@@
 $K^{a\#}$ =$=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid\forall\,y\in B_{K}:\;x^{*}(y)>\alpha\}.$
-Furthermore, it is easy to observe that $K^{a\#}\subseteq K^{a\circ}\cup K^{a+}$, and if $K$ is solid and $\psi$ is lower semicontinuous, then $K^{a\circ}\subseteq K^{a+}$.
-In the following definition, we introduce two classes of functions that satisfy cone (interior) representation properties (cf. Ha [35, Sec. 3.1], Jahn [43, Th. 3.1]).
-Definition 2.9
+Furthermore, it is easy to observe that $K^{a\#}\subseteq K^{a\circ}\cup K^{a+}$, and if $K$ is solid and convex, and $\psi$ is lower semicontinuous, then $K^{a\circ}\subseteq K^{a+}$.
+In the following definition, we introduce two classes of functions that satisfy cone (interior) representation properties (cf.
+Ha [18, Sec.
+3.1], Jahn [20, Th.
+3.1], see also Kasimbeyli [8, Th.
+3.8 and 3.9]).
+Definition 4 (Representing functionals).
 The functional $\varphi:Y\to\mathbb{R}$ is
@@
   • strictly $-K$ representing if $\varphi$ is $-K$ representing and $-\operatorname{int}\,K=\{y\in Y\mid\varphi(y)<0\}.$
-Remark 2.10
-It is known that several scalarizing functions in vector optimization (e.g., in the sense of Gerstewitz [20, 21]; Hiriart-Urruty [39] / Zaffaroni [58, Prop.
-3.2]) admit cone representation properties.
-Moreover, for certain topological-linear spaces $Y$, which are preordered by a (solid) convex cone $K\subseteq Y$, one can state $-K$ representing (strictly $-K$ representing) functionals (see, e.g., Günther et al.
-[30, Sec.
-4.1], Ha [35], and Jahn [43]).
 For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, the sublevel set of the Bishop-Phelps type function $\varphi_{x^{*},\alpha}$ at level zero is defined by
@@
 $C_{\psi}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha\psi(y)\}.$
-Remark that, in case $\psi=||\cdot||$ and $\alpha>0$, this reduces to the usual Bishop-Phelps cone (see Example 2.4). Moreover, the sets
+Remark that, in case $\psi=||\cdot||$ and $\alpha>0$, this reduces to the usual Bishop-Phelps cone (see Example 1). Moreover, the sets
 $S^{<}_{\psi}(x^{*},\alpha)$ :=$:=\{y\in Y\mid\varphi_{x^{*},\alpha}(y)=x^{*}(y)+\alpha\psi(y)<0\},$
@@
 $K^{a\#}$ =$=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)\}.$
-In the following lemmata, we collect some important properties of Bishop-Phelps type cones and functions. For the first lemma, the proof is straightforward and we omit it.
-Lemma 2.11
+Example 3 (Bishop-Phelps type cones).
+  • a)
+    Assume that $Y$ has dimension two or higher, and consider the linear continuous functional $y^{*}\in Y^{*}\setminus\{0\}$ and the seminorm $\psi:=|y^{*}|$.
+Then, for the closed, convex cone $C:=\{y\in Y\mid y^{*}(y)\geq 0\}$ which is not pointed (and therefore not well-based if $Y$ is normed; see Example 1), we have
+    $C=\{y\in Y\mid y^{*}(y)\geq|y^{*}(y)|\}=C_{\psi}(y^{*},1).$
+    This means that the cone $C$ can be written as a Bishop-Phelps type cone (with $(y^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and seminorm $\psi=|y^{*}|$).
+Note that $C$ cannot be written as a Bishop-Phelps cone (with $(y^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and norm |||||$\cdot$||| | ⋅ | |), since any Bishop-Phelps cone is pointed, but $C$ is not.
+  • b)
+    Consider the function $\psi_{\max}:Y\to\mathbb{R}$ defined by (5) in Example 2(b), where $\varphi:Y\to\mathbb{R}$ is a sublinear function.
+For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, the Bishop-Phelps type cone with respect to $\psi_{\max}$ is given by:
+    $C_{\psi_{\max}}(x^{*},\alpha)$ =$=\{y\in Y\mid x^{*}(y)\geq\alpha\max\{\varphi(y),\varphi(-y)\}\}$
+    =$=\{y\in Y\mid x^{*}(y)\geq\alpha\varphi(y)\text{ and }x^{*}(y)\geq\alpha\varphi(-y)\}\}.$
+    Moreover, note that, for any function $\psi:Y\to\mathbb{R}$, the following statements are equivalent:
+    • $1^{\circ}$
+      $\psi$ is a seminorm.
+    • $2^{\circ}$
+      There is a sublinear function $\varphi:Y\to\mathbb{R}$ such that $\psi(y)=\max\{\varphi(y),\varphi(-y)\}$ for all $y\in Y$.
+    Hence, the family of Bishop-Phelps type cones (for a fixed pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$) given by $\{C_{\psi_{\max}}(x^{*},\alpha)\mid\varphi:Y\to\mathbb{R}\text{ is sublinear}\}$ is equal to the family of Bishop-Phelps type cones given by $\{C_{\psi}(x^{*},\alpha)\mid\psi:Y\to\mathbb{R}\text{ is a seminorm}\}$.
+  • c)
+    Let $\mathcal{F}$ be the family of seminorms that generates the topology of the real locally convex space $Y$.
+Convex cones from the family of Bishop-Phelps type cones (for a fixed pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$) given by $\{C_{\psi}(x^{*},\alpha)\mid\psi\in\mathcal{F}\}$ could be of interest in order to define preorders in $Y$.
+In the case that the given space $Y$ is preordered but not partially ordered by a cone $K$, we get new ideas with considering Bishop-Phelps type type cones (since they are not necessarily pointed).
+We collect next some important properties of Bishop-Phelps type cones and functions. For the first lemma, the proof is straightforward and we omit it.
+Lemma 1.
 Consider $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. The following assertions hold:
-  • $1^{\circ}$.
+  • $1^{\circ}$
     Cone property: $0\in C_{\psi}(x^{*},\alpha),\operatorname{\mathbb{P}}\cdot C_{\psi}(x^{*},\alpha)=C_{\psi}(x^{*},\alpha)$ and $\operatorname{\mathbb{P}}\cdot C^{>}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)$.
-  • $2^{\circ}$.
+  • $2^{\circ}$
     Convexity: $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$ are convex.
-  • $3^{\circ}$.
+  • $3^{\circ}$
     Closedness: If $\varphi_{x^{*},\alpha}$ (respectively, $\psi$) is lower semicontinuous, then $C_{\psi}(x^{*},\alpha)$ is closed.
-  • $4^{\circ}$.
+  • $4^{\circ}$
     Pointedness: $\ell(C_{\psi}(x^{*},\alpha))=\{y\in Y\mid x^{*}(y)=0,\alpha\psi(y)=0\}$.
@@
 If $(x^{*},\alpha)\in K^{a\#}$, then $K\cap\ell(C_{\psi}(x^{*},\alpha))=\{0\}$.
-  • $5^{\circ}$.
+  • $5^{\circ}$
     Nontrivialness: If $(x^{*},\alpha)\in K^{a+}$, then $C_{\psi}(x^{*},\alpha)\neq\{0\}$. If $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)\neq Y$.
-  • $6^{\circ}$.
+  • $6^{\circ}$
     Solidness: Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ (e.g.
@@
 Then, $\operatorname{cor}\,C_{\psi}(x^{*},\alpha)\neq\emptyset,$ and if either $(x^{*},\alpha)\in K^{a\circ}$ and $K$ is solid or $\varphi_{x^{*},\alpha}$ (respectively, $\psi$) is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset.$
-  • $7^{\circ}$.
+  • $7^{\circ}$
     Lineality: If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ (e.g.
 if $(x^{*},\alpha)\in K^{a\#}$), and either $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)$ is not a linear subspace (or equivalently, $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$).
-  • $8^{\circ}$.
+  • $8^{\circ}$
     Dilating cone for $K$: If $(x^{*},\alpha)\in K^{a\#}$, then $C_{\psi}(x^{*},\alpha)$ is a dilating cone for $K$, i.e., $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$, and if further $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $K\setminus\{0\}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
-Lemma 2.12
+Lemma 2.
 For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, the following assertions are valid:
-  • $1^{\circ}$.
+  • $1^{\circ}$
     $-C_{\psi}(x^{*},\alpha)=S^{\leq}_{\psi}(x^{*},\alpha)$ and $-C^{>}_{\psi}(x^{*},\alpha)=S^{<}_{\psi}(x^{*},\alpha)$.
-  • $2^{\circ}$.
+  • $2^{\circ}$
     $\varphi_{x^{*},\alpha}$ is sublinear (convex) and $-C_{\psi}(x^{*},\alpha)$ representing.
-  • $3^{\circ}$.
+  • $3^{\circ}$
     $C^{>}_{\psi}(x^{*},\alpha)\subseteq\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$, and if one has $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $C^{>}_{\psi}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$.
-  • $4^{\circ}$.
+  • $4^{\circ}$
     If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
-Proof 2.13.
+Proof.
 $1^{\circ}$ Both equalities follow from the definitions.
@@
 In view of $1^{\circ}$, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing.
-$3^{\circ}$ This assertion follows from Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
+$3^{\circ}$ This assertion follows from Jahn [20, Prop. 3.6] taking into account $1^{\circ}$.
 $4^{\circ}$ Of course, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing by $2^{\circ}$.
 Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
-By $1^{\circ}$ and $3^{\circ}$ we get $-S^{<}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)=$.
-Since $C_{\psi}(x^{*},\alpha)$ is a convex set with $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ it follows that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=$.
-Thus, we conclude $-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=-$ which shows that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
-Remark 2.14.
+By $1^{\circ}$ and $3^{\circ}$ we get $-S^{<}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$.
+Since $C_{\psi}(x^{*},\alpha)$ is a convex set with $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ it follows that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$.
+Thus, we conclude $-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=-\operatorname{cor}\,C_{\psi}(x^{*},\alpha)=-C^{>}_{\psi}(x^{*},\alpha)=S^{<}_{\psi}(x^{*},\alpha),$ which shows that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
+∎
 Note that the properties of $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$, and the relationships between $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$ follow immediately from the one related to $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$.
-Furthermore, note that the conditions $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ (as given in the assumptions of Lemma 2.12 ($4^{\circ}$)) do not imply each other, as can be seen below:
+Furthermore, note that the conditions $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ (as given in the assumptions of Lemma 2 ($4^{\circ}$)) do not imply each other, as can be seen below:
 Let us first give an example with $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
-According to [37, Ex.
+According to [17, Ex.
 2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$ and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
 Next, we give an example with $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$.
-Recall that, for a set $\Omega\subseteq Y$ in a real normed space $(Y,||\cdot||)$, one denotes the weak interior and weak closure of $\Omega$ by ${\mathrm{i}nt}_{w}\,\Omega$ and ${\mathrm{c}l}_{w}\,\Omega$, respectively.
-We have the next result:
-Proposition 2.15.
+We endow a real infinite dimensional (normed) space $Y$ with the weak topology, we put $\psi(\cdot):=||\cdot||$ and we assume $||x^{*}||^{*}>\alpha>0$.
+In this case, it is known that $C^{>}_{||\cdot||}(x^{*},\alpha)=\operatorname{int}_{||\cdot||}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$.
+Since $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed cone in $(Y,||\cdot||)$ by Lemma 1 ($1^{\circ}$, $3^{\circ}$, $4^{\circ}$), we conclude that $\operatorname{int}_{w}\,C_{||\cdot||}(x^{*},\alpha)=\emptyset$, as the following proposition shows.
+Recall that, for a set $\Omega\subseteq Y$ in a real normed space $(Y,||\cdot||)$, one denotes the weak interior and weak closure of $\Omega$ by ${\rm int}_{w}\,\Omega$ and ${\rm cl}_{w}\,\Omega$, respectively.
+Proposition 3.
 Consider a closed, pointed cone $C\subseteq Y$ in a real infinite dimensional normed space $(Y,||\cdot||$). Then, we have $\operatorname{int}_{w}\,C=\emptyset$.
-Proof 2.16.
+Proof.
 On the contrary, assume that $\operatorname{int}_{w}\,C\neq\emptyset$.
@@
 Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction.
-As a desired second example, we endow a real infinite dimensional (normed) space $Y$ with the weak topology, we put $\psi(\cdot):=||\cdot||$ and we assume $||x^{*}||^{*}>\alpha>0$.
-In this case, it is known that $C^{>}_{||\cdot||}(x^{*},\alpha)=\operatorname{int}_{||\cdot||}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$.
-Since $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed cone in $(Y,||\cdot||)$ by Lemma 2.11 ($1^{\circ}$, $3^{\circ}$, $4^{\circ}$), we conclude that $\operatorname{int}_{w}\,C_{||\cdot||}(x^{*},\alpha)=\emptyset$ by Proposition 2.15.
-Remark also that, if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\psi$ is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ by Lemma 2.11 ($6^{\circ}$).
+∎
+Remark also that, if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\psi$ is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ by Lemma 1 ($6^{\circ}$).
 In the second example from above, the weak upper semicontinuity of |||||$\cdot$||| | ⋅ | | is not satisfied, since $Y$ has infinite dimension (cf.
-[29, Rem.
+[15, Rem.
 2.2]).
-2.4 Solution concepts in vector optimization
-Given two real topological-linear spaces $X$ and $Y$, a seminorm $\psi:Y\to\mathbb{R}$, a nonempty feasible set $\Omega\subseteq X$, and a vector-valued objective function $f:X\rightarrow Y$, we consider the following vector optimization problem:
+In what follows, we will briefly introduce different solution concepts in vector optimization.
+Given two real topological-linear spaces $X$ and $Y$, a nonempty feasible set $\Omega\subseteq X$, and a vector-valued objective function $f:X\rightarrow Y$, we consider the following vector optimization problem:
 $\begin{cases}f(x)\to\min\mbox{ with respect to }K\\ x\in\Omega,\end{cases}$ ($P$)
@@
 Two well-known types of solutions to the vector problem ($P$) can be defined according to the next two definitions (see, e.g., [25], [21], [32], [42]).
-Definition 2.17 (Efficiency).
+Definition 5 (Efficiency).
 A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$.
-The set of all Pareto efficient solutions of ($P$) is denoted by ${\mathrm{E}ff}(\Omega\mid f,K)$.
-Definition 2.18 (Weak efficiency).
+The set of all efficient solutions of ($P$) is denoted by ${\rm Eff}(\Omega\mid f,K)$.
+Definition 6 (Weak efficiency).
 A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$.
-The set of all weakly Pareto efficient solutions of ($P$) is denoted by ${\mathrm{W}Eff}(\Omega\mid f,K)$.
+The set of all weakly efficient solutions of ($P$) is denoted by ${\rm WEff}(\Omega\mid f,K)$.
 In addition to the efficiency concepts above, we also want to consider a proper efficiency with respect to a cone-valued map.
-Definition 2.19 (Proper efficiency with respect to a cone-valued map).
+Definition 7 (Proper efficiency with respect to a cone-valued map).
 Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$.
-A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and $\mathbb{A}(\bar{x})\cap(-K)=\{0\}$.
-The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
-Remark 2.20.
-In Definition 2.19 one can remove the assumption $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ if $\mathbb{A}(\bar{x})$ contains $f[\Omega]-f(\bar{x})$.
-Notice that ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)={\mathrm{E}ff}(\Omega\mid f,K)$ and ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,\{0\}\cup\operatorname{int}\,K)={\mathrm{W}Eff}(\Omega\mid f,K)$ with $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]-f(\cdot))$ or $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot)).$
-Remark 2.21.
-For the particular choice of the cone-valued map $\operatorname{\mathbb{A}}:\Omega\to 2^{Y}$, one can see that several well-known concepts of proper efficiency are special cases of Definition 2.19: Hurwicz proper efficiency ([40], 1958), Borwein proper efficiency ([3], 1977), Benson proper efficiency ([1], 1979), extended Borwein proper efficiency ([4], 1980).
-For more details see, e.g., [5, Sec.
-2.4.3], [25, Ch.
-3], [26], [49, Sec.
+A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$ and $\mathbb{A}(\bar{x})\cap(-K)=\{0\}$.
+The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$.
+Remark 4.
+In Definition 7 one can remove the assumption $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$ if $\mathbb{A}(\bar{x})$ contains $f[\Omega]-f(\bar{x})$.
+Notice that ${\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)={\rm Eff}(\Omega\mid f,K)$ and ${\rm PEff}_{\mathbb{A}}(\Omega\mid f,\{0\}\cup\operatorname{int}\,K)={\rm WEff}(\Omega\mid f,K)$ with $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]-f(\cdot))$ or $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot)).$
+Remark 5.
+For the particular choice of the cone-valued map $\operatorname{\mathbb{A}}:\Omega\to 2^{Y}$, one can see that several well-known concepts of proper efficiency are special cases of Definition 7: Hurwicz proper efficiency ([37], 1958), Borwein proper efficiency ([34], 1977), Benson proper efficiency ([33], 1979), extended Borwein proper efficiency ([35], 1980).
+For more details see, e.g., [25, Sec.
+2.4.3], [31, Ch.
+3], [41], [42, Sec.
 2.4] and references therein.
-Beside the proper efficiency concept given in Definition 2.19, we like also to consider a proper efficiency concept in the sense of Henig ([38], 1982).
+Beside the proper efficiency concept given in Definition 7, we like also to consider a proper efficiency concept in the sense of Henig ([36], 1982).
 To formulate this concept in the next definition, consider the following family of dilating cones:
 $\mathcal{D}(K):=\{C\subseteq Y\mid C\text{ is a nontrivial, convex cone with }K\setminus\{0\}\subseteq\operatorname{int}\,C\}.$
-Definition 2.22 (Proper efficiency in the sense of Henig).
+Definition 8 (Proper efficiency in the sense of Henig).
 A point $\bar{x}\in\Omega$ is said to be a Henig properly efficient solution of ($P$) if there is $C\in\mathcal{D}(K)$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C)$.
 The set of all Henig properly efficient solutions of ($P$) is denoted by $\operatorname{PEff}_{He}(\Omega\mid f,K)$.
-The following relationships between the different solution concepts are valid. The proof is straightforward.
-Proposition 2.23.
-Consider a cone $K\subseteq Y$ with $0\notin\operatorname{int}\,K$ (or equivalently, $K\neq Y$). Then, the following assertions hold:
-  • $1^{\circ}$.
-    $\operatorname{PEff}_{He}(\Omega\mid f,K)\subseteq\operatorname{Eff}(\Omega\mid f,K)\subseteq\operatorname{WEff}(\Omega\mid f,K)$.
-  • $2^{\circ}$.
-    $\operatorname{Eff}(\Omega\mid f,C)=\operatorname{WEff}(\Omega\mid f,K)$ for $C:=(\operatorname{int}\,K)\cup\{0\}$.
-  • $3^{\circ}$.
-    $\operatorname{PEff}_{He}(\Omega\mid f,K)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{Eff}(\Omega\mid f,C)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{WEff}(\Omega\mid f,C)$.
 3 Scalarization in Vector Optimization
-Under the assumptions of Section 2.4, consider the following scalarization of the vector optimization problem ($P$):
+Under the above assumptions, consider the following scalarization of the vector optimization problem ($P$):
 $\begin{cases}(\varphi\circ f)(x)\to\min\\ x\in\Omega,\end{cases}$ ($P_{\varphi}$)
-where $\varphi:Y\to\mathbb{R}$ is a (frequently parameter-dependent) function.
+where $\varphi:Y\to\mathbb{R}$.
 Clearly, to get useful relationships between the problems ($P$) and ($P_{\varphi}$) one needs to impose certain properties on $\varphi$.
-By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get some knowledge about the original vector problem ($P$).
+By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get certain characterizations for the solutions of the original vector problem ($P$).
 The application of such a strategy is called scalarization method, while the function $\varphi$ is called scalarizing function (see e.g.
@@
 For doing this, we recall monotonicity concepts for the function $\varphi$ (c.f.
-Jahn [42, Def.
+Jahn [32, Def.
 5.1]).
-Definition 3.1.
+Definition 9 ($K$- increasing functional).
 A scalarizing function $\varphi:Y\to\mathbb{R}$ is said to be
@@
 As an immediate consequence, any strongly $K$-increasing function is $K$-increasing, and if $K\neq Y$, strictly $K$-increasing as well.
-Lemma 3.2.
-Consider the subadditive functional $\varphi:Y\to\mathbb{R}$ and let $K,C$ be two cones in $Y$. The following assertions hold:
-  • $1^{\circ}$.
-    If $\varphi$ is (strictly) $-K$ representing, then $\varphi$ is (strictly) $K$-increasing.
-  • $2^{\circ}$.
+The following lemma collects several important relationships between cone monotone functions (based on different cones) as well as cone representing functions.
+The result will be useful to prove some basic scalarization results based on Bishop-Phelps type cone monotone scalarizing functions in Section 4.1.
+Lemma 4.
+Consider a function $\varphi:Y\to\mathbb{R}$ and two cones $K,C$ in $Y$. The following assertions hold:
+  • $1^{\circ}$
+    If $\varphi$ is (strictly) $-K$ representing and subadditive, then $\varphi$ is (strictly) $K$-increasing.
+  • $2^{\circ}$
     If $\operatorname{int}K\subseteq\operatorname{int}C$ and $\varphi$ is strictly $C$-increasing, then $\varphi$ is strictly $K$-increasing.
-  • $3^{\circ}$.
+  • $3^{\circ}$
     If $K\subseteq C$ and $\varphi$ is (strongly) $C$-increasing, then $\varphi$ is (strongly) $K$-increasing.
-Proof 3.3.
+Proof.
 $1^{\circ}$ We only prove the strict case, since the other case is similar.
@@
 The proofs of $2^{\circ}$ and $3^{\circ}$ are straightforward.
-The next scalarization results involving cone-monotone functions are well-known.
-Lemma 3.4 (c.f. [50, Lem. 5.6], [27, Lem. 5.2]).
-Consider a real-valued function $\varphi:Y\to\mathbb{R}$ and a cone $K\subseteq Y$. Then, the following assertions hold:
+∎
+The next scalarization results involving cone monotone functions (c.f. [48, Lem. 5.6], [52, Lem. 5.2]) are important for deriving the result in Theorem 12.
+Lemma 5.
+Consider a function $\varphi:Y\to\mathbb{R}$ and a cone $K\subseteq Y$. Then, the following assertions hold:
   • $1^{\circ}$
-    If $\varphi$ is strictly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
-  • $2^{\circ}$.
-    If $\varphi$ is strongly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
-  • $3^{\circ}$.
-    If $\varphi$ is $K$-increasing, and ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)=\{\overline{x}\}$ for some $\overline{x}\in\Omega$, then $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
-  • $4^{\circ}$.
-    If $\varphi$ is strictly $C$-increasing or strongly $C$-increasing for some $C\in\mathcal{D}(K)$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
-For a (strictly) $-K$ representing function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$ we have the following result.
-Lemma 3.5 (cf. [43, Th. 4.1]).
+    If $\varphi$ is strictly $K$-increasing, then ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\rm WEff}(\Omega\mid f,K)$.
+  • $2^{\circ}$
+    If $\varphi$ is strongly $K$-increasing, then ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\rm Eff}(\Omega\mid f,K)$.
+  • $3^{\circ}$
+    If $\varphi$ is $K$-increasing, and ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)=\{\overline{x}\}$ for some $\overline{x}\in\Omega$, then $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$.
+  • $4^{\circ}$
+    If $\varphi$ is strictly $C$-increasing or strongly $C$-increasing for some $C\in\mathcal{D}(K)$, then ${\rm argmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\rm PEff}_{He}(\Omega\mid f,K)$.
+We omit the proof because it is similar to the proofs for [48, Lem. 5.6] and [52, Lem. 5.2], which were established for convex cones in real linear spaces in an algebraic framework.
+For a (strictly) $-K$ representing function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$ we have the following scalarization result which will be used to prove some important characterizations in Propositions 17 and 18.
+Lemma 6.
 Suppose that $K\subseteq Y$ is a cone. Assume that $\varphi:Y\to\mathbb{R}$ is a function with $\varphi(0)=0$. Then, for any $\bar{x}\in\Omega$, we have:
-  • $1^{\circ}$.
+  • $1^{\circ}$
     If $\varphi$ is $-K$ representing, then
-    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)\iff\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
-  • $2^{\circ}$.
+    $\bar{x}\in{\rm Eff}(\Omega\mid f,K)\iff\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
+  • $2^{\circ}$
     If $\varphi$ is strictly $-K$ representing, then
-    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)\iff\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
+    $\bar{x}\in{\rm WEff}(\Omega\mid f,K)\iff\bar{x}\in{\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
+Remark 6.
+Note that Jahn [20, Th.
+4.1] stated a similar result as in $1^{\circ}$ of Lemma 6 for the case of a pointed convex cone $K$ in a real linear space $Y$.
+The proof of Lemma 6 is similar to [20, Th.
+4.1].
 Next, we recall two well-known nonlinear scalarization methods in vector optimization which are due to Pascoletti and Serafini [10], and Gerstewitz [26, 27].
 For a-priori given parameters $a\in Y$ and $k\in Y\setminus\{0\}$, in [10] the following scalar optimization problem is used
-$\begin{cases}\lambda\to\min\\ f(x)\in\lambda k+(a-K),\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ (3.1)
+$\begin{cases}\lambda\to\min\\ f(x)\in\lambda k+(a-K),\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ (7)
 while in [26, 27] the scalar optimization problem is given as
-$\begin{cases}\varphi^{G}_{a-K,k}(f(x))\to\min\\ x\in\Omega,\end{cases}$ (3.2)
-where the nonlinear scalarizing function $\varphi^{G}_{a-K,k}:Y\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty,+\infty\}$ is defined (for $k\in K\setminus\ell(K)$ and $K$ closed) by
-$\varphi^{G}_{a-K,k}(y):=\inf\{t\in\mathbb{R}\mid y\in tk+(a-K)\}\quad\text{forall }y\in Y.$ (3.3)
-Many facts about the problems (3.1) and (3.2), as well as about the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$, are known (see, e.g., [20, 21], [53], [22], [12, 13], [49], [57], [23]).
-For the sake of completeness, we recall some important facts below.
-Lemma 3.6.
+$\begin{cases}\varphi^{G}_{a-K,k}(f(x))\to\min\\ x\in\Omega,\end{cases}$ (8)
+where the nonlinear scalarizing function $\varphi^{G}_{a-K,k}:Y\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty,+\infty\}$ is defined (for $k\in K\setminus\ell(K)$, and $K$ closed and convex) by
+$\varphi^{G}_{a-K,k}(y):=\inf\{t\in\mathbb{R}\mid y\in tk+(a-K)\}\quad\text{forall }y\in Y.$ (9)
+Many facts about the problems (7) and (8), as well as about the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$, are known (see, e.g., [26, 27], [10], [28], [4, 5], [42], [12], [21]).
+Lemma 7.
 Suppose that $K\subseteq Y$ is a closed, convex cone and take $k\in K\setminus\ell(K)$.
-Then, the scalar problems (3.1) and (3.2) are equivalent (i.e., each problem has an optimal solution if and only if the other one has an optimal solution, and the optimal solutions $x$ as well as the optimal value of both problems coincide).
-Proof 3.7.
-Under the given assumptions, the function $\varphi^{G}_{a-K,k}$ has the level set property $f(x)\in\lambda k+(a-K)\iff\varphi^{G}_{a-K,k}(f(x))\leq\lambda$ (see [23, Th.
-2.4.1]).
-Thus, (3.1) is equivalent to
-$\begin{cases}\lambda\to\min\\ \varphi^{G}_{a-K,k}(f(x))\leq\lambda,\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$
-respectively, to (3.2).
-In the next lemma, we state classical relationships between the vector problem ($P$) and the scalar problem (3.1) (see, e.g., Eichfelder [12, Th. 2.1 and Cor. 2.9]).
-Lemma 3.8.
+Then, the scalar problems (7) and (8) are equivalent (i.e., each problem has an optimal solution if and only if the other one has an optimal solution, and the optimal solutions $x$ as well as the optimal value of both problems coincide).
+Proof.
+This fact is well-known (see [12, Th. 5.1.3]). ∎
+In the next lemma, we state classical relationships between the vector problem ($P$) and the scalar problem (7) (see, e.g., Eichfelder [4, Th. 2.1 and Cor. 2.9] for the finite dimensional case).
+Lemma 8.
 Suppose that $K\subseteq Y$ is a convex cone. Then, we have:
-  • $1^{\circ}$.
-    If $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in K\setminus\ell(K)$.
-More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1).
-  • $2^{\circ}$.
-    If $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,K$.
-  • $3^{\circ}$.
+  • $1^{\circ}$
+    If $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (7) for $a=f(\bar{x})$ and any $k\in K\setminus\ell(K)$.
+More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (7).
+  • $2^{\circ}$
+    If $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (7) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,K$.
+  • $3^{\circ}$
     Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$.
-If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
-  • $4^{\circ}$.
-    If $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$ solves the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$ and $f(\bar{x})\in\bar{\lambda}k+(a-\operatorname{bd}\,K)$.
-Alternatively, one can immediately obtain scalarization results by using the relationships between (3.1) and (3.2), and properties of the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$ defined in (3.3) (see [57, Ch.
-6] for an extended overview on this topic).
-Lemma 3.9.
-Suppose that $K\subseteq Y$ is a closed, convex cone, and take $k\in K\setminus\ell(K)$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
-  • $1^{\circ}$.
-    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$
-  • $2^{\circ}$.
-    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
-  • $3^{\circ}$.
-    $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the solution set of the problem (3.1) for $a=f(\bar{x})$.
-Lemma 3.10.
-Suppose that $K\subseteq Y$ is a nontrivial, closed, solid, convex cone, and take $k\in\operatorname{int}\,K$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
-  • $1^{\circ}$.
-    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$
-  • $2^{\circ}$.
-    $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
-  • $3^{\circ}$.
-    $(\bar{x},0)$ is a solution of the problem (3.1) for $a=f(\bar{x})$.
-The scalar optimization problems of type (3.1) and (3.2) involved in the classical scalarization results given in Lemmata 3.8, 3.9 and 3.10 have the drawback that the parameter $a$ depends on the solution $\bar{x}$ (more precisely, we have $a=f(\bar{x})$).
-In order to avoid this disadvantage, Jahn [43, Th.
+If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (7) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$.
+  • $4^{\circ}$
+    If $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$ solves the problem (7) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$ and $f(\bar{x})\in\bar{\lambda}k+(a-\operatorname{bd}\,K)$.
+The scalar optimization problems of type (7) and (8) involved in the classical scalarization results (see, e.g., Eichfelder [4, Th.
+2.1 and Cor.
+2.9]) have the drawback that the parameter $a$ depends on the solution $\bar{x}$ (more precisely, we have $a=f(\bar{x})$).
+In order to avoid this disadvantage, Jahn [20, Th.
 4.2] stated a scalarization result where the parameter $a$ does not depend on the solution $\bar{x}$.
-In the following proposition, we present a new scalarization result that contains the result of Jahn [43, Th. 4.2] as a special case (see Remark 3.13 for details).
-Proposition 3.11.
-Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$.
+In the following proposition, we present a new scalarization result that contains the result of Jahn [20, Th. 4.2] as a special case (see Remark 7 for details).
+Proposition 9.
+Suppose that $K\subseteq Y$ is a convex cone, and take sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$.
 Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
-  • $1^{\circ}$.
-    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
-  • $2^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of (3.1).
+  • $1^{\circ}$
+    $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$.
+  • $2^{\circ}$
+    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of (7).
 If, in addition, $K$ is closed, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
-  • $3^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
-Proof 3.12.
-The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is a direct consequence of Lemma 3.8 ($3^{\circ}$).
+  • $3^{\circ}$
+    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
+Proof.
+The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is a direct consequence of Lemma 8 ($3^{\circ}$).
 Now, we show the implication $1^{\circ}\Longrightarrow 2^{\circ}$.
-Let $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
-Take any $k\in\mathcal{P}_{k}\,(\subseteq K\setminus\ell(K))$.
-By Lemma 3.8 ($1^{\circ}$), the set of solutions of the problem (3.1) (with $f(\bar{x})$ in the role of $a$) is given by $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$.
-Note that, for any $x\in\Omega$, the pair ($x,0$) solves (3.1) (with $f(\bar{x})$ in the role of $a$) if and only if ($x,t$) solves (3.1) with $f(\bar{x})-tk$ in the role of $a$ for all $t\in\mathbb{R}$.
-By our assumption $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, there are $\bar{a}\in\mathcal{P}_{a}$, $\bar{k}\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ such that $f(\bar{x})=\bar{a}+s\bar{k}$.
-Thus, $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) with $\bar{a}=f(\bar{x})-s\bar{k}\in\mathcal{P}_{a}$ and $\bar{k}\in\mathcal{P}_{k}$ in the role of $a$ and $k$.
-If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
-Remark 3.13.
+Let $\bar{x}\in{\rm Eff}(\Omega\mid f,K)$.
+Take any $k\in\mathcal{T}\,(\subseteq K\setminus\ell(K))$.
+By Lemma 8 ($1^{\circ}$), the set of solutions of the problem (7) (with $f(\bar{x})$ in the role of $a$) is given by $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$.
+Note that, for any $x\in\Omega$, the pair ($x,0$) solves (7) (with $f(\bar{x})$ in the role of $a$) if and only if ($x,t$) solves (7) with $f(\bar{x})-tk$ in the role of $a$ for all $t\in\mathbb{R}$.
+By our assumption $f[{\rm Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$, there are $\bar{a}\in\mathcal{P}$, $\bar{k}\in\mathcal{T}$ and $s\in\mathcal{R}$ such that $f(\bar{x})=\bar{a}+s\bar{k}$.
+Thus, $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (7) with $\bar{a}=f(\bar{x})-s\bar{k}\in\mathcal{P}$ and $\bar{k}\in\mathcal{T}$ in the role of $a$ and $k$.
+If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 7. ∎
+Remark 7.
 […]
-For the concept of weak efficiency we have the following result.
-Proposition 3.14.
-Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq\operatorname{int}\,K$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$.
+For the concept of weak efficiency we have the following new result.
+Proposition 10.
+Suppose that $K\subseteq Y$ is a convex cone, and take sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq\operatorname{int}\,K$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm WEff}(\Omega\mid f,K)]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$.
 Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
-  • $1^{\circ}$.
-    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
-  • $2^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of (3.1).
+  • $1^{\circ}$
+    $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$.
+  • $2^{\circ}$
+    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $(\bar{x},s)$ is a solution of (7).
 If, in addition, $K$ is closed and nontrivial, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
-  • $3^{\circ}$.
-    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
-Proof 3.15.
-The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 3.8 ($4^{\circ}$) taking into account that $\operatorname{int}\,K\subseteq K\setminus\{0\}$, or equivalently $K\neq Y$ (since (3.1) has a solution), while the proof of the implication $1^{\circ}\Longrightarrow 2^{\circ}$ is similar to the proof of the corresponding implication in Proposition 3.11.
-If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
-Remark 3.16.
-Assuming that $K$ has a normlike-base $B_{K}(\psi)$ (which means that the seminorm $\psi:Y\to\mathbb{R}$ is positive on $K\setminus\{0\}$) and $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq a+K\setminus\ell(K)$ (respectively, $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq a+\operatorname{int}\,K$) for some $a\in Y$, in Proposition 3.11 (respectively, Proposition 3.14), one may take $\mathcal{P}_{a}:=\{a\}$, $\mathcal{P}_{k}:=B_{\{0\}\cup(K\setminus\ell(K))}(\psi)$ (respectively, $\mathcal{P}_{k}:=B_{\{0\}\cup\operatorname{int}\,K}(\psi)$ if $0\notin\operatorname{int}\,K$, or equivalently, $K\neq Y$) and $\mathcal{R}:=\mathbb{P}$.
-The following proposition provides possible choices for the sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ to ensure $f[D]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ with $D={\mathrm{E}ff}(\Omega\mid f,K)$, $D={\mathrm{W}Eff}(\Omega\mid f,K)$ and $D=\Omega$, respectively, in Propositions 3.11 and 3.14.
-Proposition 3.17.
+  • $3^{\circ}$
+    There are some $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\rm min}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
+Proof.
+The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 8 ($4^{\circ}$) taking into account that $\operatorname{int}\,K\subseteq K\setminus\{0\}$, or equivalently $K\neq Y$ (since (7) has a solution), while the proof of the implication $1^{\circ}\Longrightarrow 2^{\circ}$ is similar to the proof of the corresponding implication in Proposition 9 by using Lemma 8 ($2^{\circ}$).
+If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 7. ∎
+Remark 8.
+Assuming that $K$ has a normlike-base $B_{K}(\psi)$ (which means that the seminorm $\psi:Y\to\mathbb{R}$ is positive on $K\setminus\{0\}$) and $f[{\rm Eff}(\Omega\mid f,K)]\subseteq a+K\setminus\ell(K)$ (respectively, $f[{\rm WEff}(\Omega\mid f,K)]\subseteq a+\operatorname{int}\,K$) for some $a\in Y$, in Proposition 9 (respectively, Proposition 10), one may take $\mathcal{P}:=\{a\}$, $\mathcal{T}:=B_{\{0\}\cup(K\setminus\ell(K))}(\psi)$ (respectively, $\mathcal{T}:=B_{\{0\}\cup\operatorname{int}\,K}(\psi)$ if $0\notin\operatorname{int}\,K$, or equivalently, $K\neq Y$) and $\mathcal{R}:=\mathbb{P}$.
+The following proposition provides possible choices for the sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ to ensure $f[D]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$ with $D={\rm Eff}(\Omega\mid f,K)$, $D={\rm WEff}(\Omega\mid f,K)$ and $D=\Omega$, respectively, in Propositions 9 and 10.
+Proposition 11.
 Let $K\subseteq Y$ be a nontrivial, convex cone and let $y^{*}\in Y^{*}\setminus\{0\}$, $a\in Y$ and $k\in K\setminus\{0\}$ be given. Suppose that $y^{*}(k)>0$ (e.g. if $y^{*}\in K^{\#}$). Define
@@
 Then, the following assertions are valid:
-  • $1^{\circ}$.
+  • $1^{\circ}$
     $H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $H_{y^{*},a}^{<}=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
-  • $2^{\circ}$.
+  • $2^{\circ}$
     If $y^{*}\in K^{\#}$, then $a+K\setminus\{0\}\subseteq H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $a-K\setminus\{0\}\subseteq H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
-  • $3^{\circ}$.
+  • $3^{\circ}$
     If $y^{*}\in K^{+}$ and $k\in\operatorname{int}\,K$, then $a+\operatorname{int}\,K\subseteq H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $a-\operatorname{int}\,K\subseteq H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
-  • $4^{\circ}$.
+  • $4^{\circ}$
     $Y=\bigcup_{\sim\in\{<,=,>\}}\,H_{y^{*},a}^{\sim}=H_{y^{*},a}^{=}+\mathbb{R}\cdot k$.
-Proof 3.18.
-$1^{\circ}$.
-Let us verify the equality $H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$.
-Since $a\in H_{y^{*},a}^{=}$, $y^{*}(k)>0$ and $\mathbb{P}\cdot k\subseteq K\setminus\{0\}$, we conclude $H_{y^{*},a}^{=}+\mathbb{P}\cdot k\subseteq H_{y^{*},a}^{>}$.
-Take some $y\in H_{y^{*},a}^{>}$.
-Noting that $y-tk\in H_{y^{*},a}^{=}$ for some $t>0$, we derive $y=y-tk+tk\in H_{y^{*},a}^{=}+\mathbb{P}\cdot k$, which shows the reverse inclusion.
-Applying the result for the cone $-K$, the functional $-y^{*}$ and the direction $-k$ (in the role of $K$, $y^{*}$ and $k$), we get immediately $H_{y^{*},a}^{<}=H_{(-y^{*}),a}^{>}=H_{(-y^{*}),a}^{=}+\mathbb{P}\cdot(-k)=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
-$2^{\circ}$.
-Assume that $y^{*}\in K^{\#}$, which means that $y^{*}(x)>0$ for all $x\in K\setminus\{0\}$.
-Then, using $1^{\circ}$ we easily derive $a+K\setminus\{0\}\subseteq H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$.
-Applying the result for $-K$, $-y^{*}$ and $-k$ (in the role of $K$, $y^{*}$ and $k$), we obtain $a-K\setminus\{0\}\subseteq H_{-y^{*},a}^{=}-\mathbb{P}\cdot k=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
-$3^{\circ}$.
-Define $C:=({\mathrm{i}nt}\,K)\cup\{0\}$.
-Note that ${\mathrm{i}nt}\,K=C\setminus\{0\}$ and $K^{+}\setminus\{0\}=\{x^{*}\in Y^{*}\mid\forall x\in{\mathrm{i}nt}\,K:x^{*}(x)>0\}=C^{\#}$ by [50, Th.
+Proof.
+The proofs for $1^{\circ}$ and $2^{\circ}$ are straightforward, while $3^{\circ}$ and $4^{\circ}$ follow easily by using $1^{\circ}$ and $2^{\circ}$, taking into account that, for a nontrivial, solid, convex cone $K$ and $C:=({\rm int}\,K)\cup\{0\}$ we have ${\rm int}\,K=C\setminus\{0\}$ and $K^{+}\setminus\{0\}=\{x^{*}\in Y^{*}\mid\forall x\in{\rm int}\,K:x^{*}(x)>0\}=C^{\#}$ (see also [48, Th.
 4.8, Cor.
-4.9] (if $K$ is a solid, convex cone).
-Applying assertion $2^{\circ}$ for $C$ in the role of $K$ we get the result.
-$4^{\circ}$. The first equality is obvious, while the second equality follows by $1^{\circ}$.
-Remark 3.19.
-Let the assumptions of Proposition 3.17 be fulfilled.
-Eichfelder’s adaptive scalarization method in multiobjective optimization proposed in [12, 13] is based on parameterized scalar optimization problems of type (3.1) and (3.2), respectively, with different choices of the reference point $a\in H_{y^{*},a}^{=}$ and an a-priori fixed direction $k\in K\setminus\{0\}$.
-Note that Proposition 3.11 generalizes a result by Eichfelder [13, Th.
-3.2] taking into account that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq Y=\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ for $\mathcal{P}_{a}:=H_{y^{*},a}^{=}$, $\mathcal{P}_{k}:=\{k\}$ and $\mathcal{R}:=\mathbb{R}$ by Proposition 3.17 ($4^{\circ}$).
-4 Nonlinear Cone Separation Conditions
-The key tool to prove our main Bishop-Phelps type scalarization results in vector optimization in Section 5.3 will be the separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions.
+4.9]).
+∎
+Remark 9.
+Let the assumptions of Proposition 11 be fulfilled.
+Eichfelder’s adaptive scalarization method in multiobjective optimization proposed in [4, 5] is based on parameterized scalar optimization problems of type (7) and (8), respectively, with different choices of the reference point $a\in H_{y^{*},a}^{=}$ and an a-priori fixed direction $k\in K\setminus\{0\}$.
+Note that Proposition 9 generalizes a result by Eichfelder [5, Th.
+3.2] taking into account that $f[{\rm Eff}(\Omega\mid f,K)]\subseteq Y=\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$ for $\mathcal{P}:=H_{y^{*},a}^{=}$, $\mathcal{T}:=\{k\}$ and $\mathcal{R}:=\mathbb{R}$ by Proposition 11 ($4^{\circ}$).
+4 Bishop-Phelps Type Scalarization in Vector Optimization
+Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$.
+Consider a sublinear functional $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$.
+As highlighted in Section 1, for any given point $a\in Y$ and any direction $k\in Y\setminus\{0\}$, the following scalarizations of the vector problem ($P$) will play a key role in our studies:
+$\begin{cases}\varphi(f(x)-a)\to\min\\ x\in\Omega\end{cases}$ ($P_{\varphi}^{a}$)
+and
+$\begin{cases}\lambda\to\min\\ \varphi(f(x)-a-\lambda k)\leq 0,\\ (x,\lambda)\in\Omega\times\mathbb{R}.\end{cases}$ ($P_{\varphi}^{a,k}$)
+When dealing with the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$), we will impose that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing for some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$.
+Thus, the Bishop-Phelps type cone $C_{\psi}(x^{*},\alpha)$ plays an important role in our scalarization approach due to the cone representing property of $\varphi$.
+This is why our method is called Bishop-Phelps type scalarization in vector optimization.
+Remark 10.
+It is known that several scalarizing functions in vector optimization (e.g., in the sense of Gerstewitz [26, 27]; Hiriart-Urruty [53] / Zaffaroni [11, Prop.
+3.2]) admit cone representation properties.
+Such functions can therefore be used well in our approach.
+Let us describe two types of scalarizing functions in more detail:
+  • The nonlinear scalarizing function $\varphi^{G}_{-C_{\psi}(x^{*},\alpha),k}$ (i.e., the function defined in (9) with $a=0$ and with $C_{\psi}(x^{*},\alpha)$ in the role of $K$) is sublinear and $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) with $\varphi^{G}_{-C_{\psi}(x^{*},\alpha),k}(0)=0$, if $C_{\psi}(x^{*},\alpha)$ is nontrivial and closed, and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha)$ (respectively, $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
+  • According to Lemma 2 ($2^{\circ}$, $4^{\circ}$), the Bishop-Phelps type (seminorm-linear) function $\varphi_{x^{*},\alpha}$ introduced in (6) is sublinear and $-C_{\psi}(x^{*},\alpha)$ representing with $\varphi_{x^{*},\alpha}(0)=0$, and if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
+4.1 Basic scalarization results
+Theorem 12.
+For any $a\in Y$, the following assertions hold:
+  • $1^{\circ}$
+    If $(x^{*},\alpha)\in K^{a\circ}$, and $\varphi$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing (e.g., if $\varphi$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing), then ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)\subseteq{\rm WEff}(\Omega\mid f,K)$.
+  • $2^{\circ}$
+    If $(x^{*},\alpha)\in K^{a\#}$ and $\varphi$ is strongly $(C_{\psi}^{>}(x^{*},\alpha)\cup\{0\})$-increasing, then ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)\subseteq{\rm Eff}(\Omega\mid f,K)$.
+  • $3^{\circ}$
+    Assume that $\overline{x}\in\Omega$.
+If $(x^{*},\alpha)\in K^{a+}$, $\varphi$ is $C_{\psi}(x^{*},\alpha)$-increasing, and ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)=\{\overline{x}\}$, then $\overline{x}\in{\rm Eff}(\Omega\mid f,K)$.
+  • $4^{\circ}$
+    If $(x^{*},\alpha)\in K^{a\#}$, $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, and $\varphi$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing or strongly $C_{\psi}(x^{*},\alpha)$-increasing, then ${\rm argmin}_{x\in\Omega}\;\varphi(f(x)-a)\subseteq{\rm PEff}_{He}(\Omega\mid f,K)$.
+Proof.
+$1^{\circ}$ For $\operatorname{int}\,K=\emptyset$ the result is clear.
+Assume that $\operatorname{int}\,K\neq\emptyset$.
+Consider $(x^{*},\alpha)\in K^{a\circ}$.
+Obviously, if $\varphi$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing, then $\varphi(\cdot-a)$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing as well.
+Because $(x^{*},\alpha)\in K^{a\circ}$ we have $\emptyset\neq\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
+Finally, Lemma 4 ($2^{\circ}$) and Lemma 5 ($1^{\circ}$) yield the conclusion.
+$2^{\circ}$ and $3^{\circ}$ The proof of the assertions is similar to the proof of assertion $1^{\circ}$ by using Lemma 4 ($3^{\circ}$) and Lemma 5 ($2^{\circ},3^{\circ}$).
+$4^{\circ}$ Take some $(x^{*},\alpha)\in K^{a\#}$.
+Since $\varphi$ is strictly (respectively, strongly) $C_{\psi}(x^{*},\alpha)$-increasing we have that $\varphi(\cdot-a)$ is strictly (respectively, strongly) $C_{\psi}(x^{*},\alpha)$-increasing as well.
+Because $(x^{*},\alpha)\in K^{a\#}$ we get $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
+Taking into account that $C_{\psi}(x^{*},\alpha)$ is a nontrivial convex cone, we conclude that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$.
+Finally, with the aid of Lemma 5 ($4^{\circ}$) we get the desired inclusion.
+∎
+In particular, we like to point out monotonicity properties of the seminorm-linear function $\varphi_{x^{*},\alpha}$ with respect to a Bishop-Phelps type cone $C_{\psi}(x^{*},\alpha)$.
+Lemma 13.
+Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. Then, the following assertions hold:
+  • $1^{\circ}$
+    $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing.
+  • $2^{\circ}$
+    If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
+  • $3^{\circ}$
+    If $K$ is solid and $(x^{*},\alpha)\in K^{a\circ}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
+  • $4^{\circ}$
+    If $(x^{*},\alpha)\in K^{a\#}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
+  • $5^{\circ}$
+    $\varphi_{x^{*},\alpha}$ is strongly $(C_{\psi}^{>}(x^{*},\alpha)\cup\{0\})$-increasing.
+Proof.
+$1^{\circ}$ Noting that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing, the assertion follows from Lemma 4 ($1^{\circ}$).
+$2^{\circ}$ If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$, there is nothing to prove.
+If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, using the assumption $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ we have that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing, and the result follows from Lemma 4 ($1^{\circ}$).
+$3^{\circ}$ and $4^{\circ}$ Using the assumptions it follows that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, so by assertion $2^{\circ}$ we get the conclusion.
+$5^{\circ}$ It follows from the definitions. ∎
+The next lemma, which studies the monotonicity properties of $\varphi_{x^{*},\alpha}$ with respect to the original cone $K$, extends a corresponding result derived by Kasimbeyli [8, Th.
+3.5] in a normed setting.
+Lemma 14.
+Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
+  • $1^{\circ}$
+    $(x^{*},\alpha)\in K^{a+}$ $\iff$ $\varphi_{x^{*},\alpha}$ is $K$-increasing.
+  • $2^{\circ}$
+    $(x^{*},\alpha)\in K^{a\circ}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing.
+  • $3^{\circ}$
+    $(x^{*},\alpha)\in K^{a\#}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing.
+Proof.
+$1^{\circ}$ By Lemma 13 ($1^{\circ}$), the function $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing.
+If $(x^{*},\alpha)\in K^{a+}$, then $K\subseteq C_{\psi}(x^{*},\alpha)$, hence $\varphi_{x^{*},\alpha}$ is $K$-increasing by Lemma 4 ($3^{\circ}$).
+Conversely, assuming $\varphi_{x^{*},\alpha}$ is $K$-increasing, for $k\in K$ we have $-k\leq_{K}0$, hence $\varphi_{x^{*},\alpha}(-k)\leq\varphi_{x^{*},\alpha}(0)=0$, which means that $x^{*}(k)-\alpha\psi(k)\geq 0$.
+Noting that $x^{*}\in K^{+}$, we conclude $(x^{*},\alpha)\in K^{a+}$.
+$2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid.
+Assume that $\operatorname{int}\,K\neq\emptyset$.
+If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 13 ($2^{\circ}$) and 4 ($2^{\circ}$).
+The proof of the reverse implication is similar to $1^{\circ}$.
+$3^{\circ}$ If $(x^{*},\alpha)\in K^{a\#}$, then $K\subseteq C_{\psi}^{>}(x^{*},\alpha)\cup\{0\}$, so $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing by Lemmata 13 ($5^{\circ}$) and 4 ($3^{\circ}$).
+Also here the proof of the reverse implication is similar to $1^{\circ}$.
+∎
+Let us shift our focus to relationships between the parameterized scalar problem ($P_{\varphi}^{a,k}$) and the vector problem ($P$) involving a Bishop-Phelps type cone.
+Lemma 15.
+Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ and assume that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing. Then, the following assertions are valid:
+  • $1^{\circ}$
+    If $\bar{x}\in{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
+More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$).
+  • $2^{\circ}$
+    If $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
+  • $3^{\circ}$
+    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$.
+If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+  • $4^{\circ}$
+    If $(\bar{x},\bar{\lambda})$ solves the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+Proof.
+Follows immediately from Lemma 8 taking into account that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing. ∎
+Theorem 16.
+Consider $\bar{x}\in\Omega$.
+Take some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, $a\in Y,k\in Y\setminus\{0\}$ and assume that $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing and that $(\bar{x},\bar{\lambda})$ is a solution of the optimization problem ($P_{\varphi}^{a,k}$).
+Then:
+  • $1^{\circ}$
+    If $(x^{*},\alpha)\in K^{a\#}$ and $C_{\psi}(x^{*},\alpha)$ is solid, then $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$.
+  • $2^{\circ}$
+    If $(x^{*},\alpha)\in K^{a\circ}$, then $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$.
+Proof.
+$1^{\circ}$ Remark that $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in view of Lemma 15 ($4^{\circ}$).
+The latter set is contained in ${\rm PEff}_{He}(\Omega\mid f,K)$ (since $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$).
+$2^{\circ}$ For the case $\operatorname{int}\,K=\emptyset$ the statement is clear.
+Assume that $\operatorname{int}\,K\neq\emptyset$.
+Again we have $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+Here the latter set is contained in ${\rm WEff}(\Omega\mid f,K)$ (since $\emptyset\neq\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
+∎
+Let us summarize our findings related to vector problems involving a Bishop-Phelps type cone in the next two propositions.
+Proposition 17.
+Let $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ and $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing.
+Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
+  • $1^{\circ}$
+    $\bar{x}\in{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+  • $2^{\circ}$
+    For any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$, the set $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
+  • $3^{\circ}$
+    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x}))$.
+If there are sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$, then $1^{\circ}-3^{\circ}$ are equivalent to:
+If further $C_{\psi}(x^{*},\alpha)$ is closed, then $1^{\circ}-4^{\circ}$ are equivalent to:
+  • $5^{\circ}$
+    There are $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
+Proof.
+Note that $C_{\psi}(x^{*},\alpha)$ is a convex cone, $\varphi$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi(0)=0$ and $C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))\neq\emptyset$.
+$1^{\circ}\iff 2^{\circ}$ The equivalence follows immediately from Lemma 15 ($1^{\circ}$, $3^{\circ}$).
+$1^{\circ}\iff 3^{\circ}$ The result follows directly from Lemma 6 ($1^{\circ}$).
+$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$ The result follows from Proposition 9 taking into account that ($P_{\varphi}^{a,k}$) is the same problem as (7) with $C_{\psi}(x^{*},\alpha)$ in the role of $K$.
+∎
+Proposition 18.
+Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing.
+Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
+  • $1^{\circ}$
+    $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+  • $2^{\circ}$
+    For any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, the point $(\bar{x},0)$ is a solution of the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
+If $\varphi$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing (e.g., if $\varphi:=\varphi_{x^{*},\alpha}$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$), then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
+  • $3^{\circ}$
+    $\bar{x}\in{\rm argmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x}))$.
+If there are sets $\mathcal{P}\subseteq Y$, $\mathcal{T}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}+\mathcal{R}\cdot\mathcal{T}$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
+If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
+  • $5^{\circ}$
+    There are $a\in\mathcal{P}$, $k\in\mathcal{T}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\rm min}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
+Proof.
+First, note that $C_{\psi}(x^{*},\alpha)$ is a solid, convex cone and $\varphi$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi(0)=0$.
+$1^{\circ}\iff 2^{\circ}$ The equivalence follows immediately from Lemma 15 ($2^{\circ}$, $4^{\circ}$).
+$1^{\circ}\iff 3^{\circ}$ The result follows from Lemma 6 ($2^{\circ}$).
+Note, since $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, the function $\varphi$ is strictly $-C_{\psi}(x^{*},\alpha)$-representing by Lemma 2 ($4^{\circ}$) with $\varphi(0)=0$.
+$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$ The result follows directly from the Proposition 10. ∎
+Remark 11.
+The results derived in Theorem 12 (combined with Lemma 13) and Propositions 17 and 18 for the case $\varphi=\varphi_{x^{*},\alpha}$ extend some known results in the literature, which are given in a normed space $(Y,\psi)$, $\psi(\cdot):=||\cdot||$, preordered by a convex cone $K\subseteq Y$:
+  • Gasimov [6, Th. 1] proved a corresponding statement as in Theorem 12 ($4^{\circ}$) for Benson proper efficiency (instead of Henig proper efficiency).
+  • Kasimbeyli [9, Th. 5.4] stated similar assertions as in Theorem 12 for a closed, pointed, convex cone $K$ and Benson proper efficiency (instead of Henig proper efficiency).
+  • Gasimov [6, Th.
+2] (see also Guo and Zhang [44, Th.
+4.5]) stated that $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$ (with $\mathbb{A}$ considered in Benson proper efficiency) implies $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\rm argmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
+Remark 12.
+Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$.
+Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\rm Eff}(\Omega\mid f,K)$.
+If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g.
+if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\rm PEff}_{He}(\Omega\mid f,K)$.
+If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence $\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\rm WEff}(\Omega\mid f,K)$.
+4.2 Nonlinear Cone Separation Conditions
+The key tool to prove our main Bishop-Phelps type scalarization results in vector optimization will be the separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions.
 In this section we present all the necessary basics for nonlinear cone separation.
@@
 The following conditions will be of interest in our upcoming auxiliary results:
-$(\operatorname{cl}\,S_{A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (4.1)
-$A\cap(\operatorname{cl}(\operatorname{conv}(-K))=\{0\},$ (4.2)
-$0\notin\operatorname{cl}\,S_{-K}(\psi).$ (4.3)
-The following proposition will later play an important role in proving a Bishop-Phelps type scalarization result (namely Corollary 5.28) for vector optimization problems in real locally convex spaces.
-Proposition 4.1.
-Assume that (4.2) and (4.3) are valid, and one of the following statements is satisfied:
+$(\operatorname{cl}\,S_{A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (10)
+$A\cap(\operatorname{cl}(\operatorname{conv}(-K))=\{0\},$ (11)
+$0\notin\operatorname{cl}\,S_{-K}(\psi).$ (12)
+Remark 13.
+We would like to point out that a separation condition for two (not necessarily convex) cones,
+$(\operatorname{cl}\,S_{{\rm bd}\,A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (13)
+which is a similar condition to the one in (10) (more precisely, $A$ is replaced by ${\rm bd}\,A$), was initially studied by Kasimbeyli [8, Def.
+4.1, Th.
+4.3], [9] and more recently by García-Castaño, Melguizo-Padial and Parzanese [1] in the context of cone separation based on Bishop-Phelps separating cones / functions in real normed spaces (with $\psi$ given by the underlying norm of the space).
+Moreover, in the same framework, Günther, Khazayel and Tammer [15] studied the separation condition (10) and compared their results with those results (from [8] and [1]) obtained for the separation condition (13).
+In [14], the study of the separation condition (10) and cone separation based on Bishop–Phelps type separating cones / functions is conducted within the more general framework of real topological-linear spaces.
+[…]
+Proposition 19.
+Assume that (11) and (12) are valid, and one of the following statements is satisfied:
   1. a)
@@
     There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that
-    $\operatorname{cl}\,S_{-K}(\psi)$ =$=\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ (4.4)
+    $\operatorname{cl}\,S_{-K}(\psi)$ =$=\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ (14)
     is valid, and one of the following conditions is satisfied:
-    $B_{A}\text{ is compact},$ (4.5)
-    $S_{A}^{0}\text{ is closed},$ (4.6)
-    $A\text{ is closed and }\operatorname{cl}\,B_{A}\text{ is compact},$ (4.7)
-    $A\text{ is closed and }\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\text{ is closed}.$ (4.8)
-Then, (4.1) is valid.
-Proof 4.2.
-a) Assuming $A$ is closed and convex, by (4.2) we get $\operatorname{cl}\,S_{A}^{0}\cap\operatorname{cl}\,S_{-K}\subseteq A\cap\operatorname{cl}(\operatorname{conv}(-K))=\{0\}$, and so by (4.3) we conclude (4.1).
-b) Obviously, (4.2) and (4.3) imply $A\cap\operatorname{cl}\,S_{-K}=\emptyset$.
-Since there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with (4.4), we get that $x^{*}(x)<\alpha$ for all $x\in A\cap(\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi})$.
+    $B_{A}\text{ is compact},$ (15)
+    $S_{A}^{0}\text{ is closed},$ (16)
+    $A\text{ is closed and }\operatorname{cl}\,B_{A}\text{ is compact},$ (17)
+    $A\text{ is closed and }\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\text{ is closed}.$ (18)
+Then, (10) is valid.
+Proof.
+a) Assuming $A$ is closed and convex, by (11) we get $\operatorname{cl}\,S_{A}^{0}\cap\operatorname{cl}\,S_{-K}\subseteq A\cap\operatorname{cl}(\operatorname{conv}(-K))=\{0\}$, and so by (12) we conclude (10).
+b) Obviously, (11) and (12) imply $A\cap\operatorname{cl}\,S_{-K}=\emptyset$.
+Since there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with (14), we get that $x^{*}(x)<\alpha$ for all $x\in A\cap(\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi})$.
 Note that $B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$, and if $A$ is closed, then $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$.
 Consider cases:
-Case 1: (4.5) is valid.
+Case 1: (15) is valid.
 By the continuity of $x^{*}$ and the compactness of $B_{A}\cup\{0\}$, we get some $\beta\in(0,\alpha)$ such that $B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
 Hence, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
-Case 2: (4.6) is valid. Then, $\operatorname{cl}\,S_{A}^{0}=S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
-Case 3: (4.7) is valid.
+Case 2: (16) is valid. Then, $\operatorname{cl}\,S_{A}^{0}=S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
+Case 3: (17) is valid.
 By the continuity of $x^{*}$ and the compactness of $\operatorname{cl}\,B_{A}\cup\{0\}$ we get some $\beta\in(0,\alpha)$ such that $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
 Thus, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
-Case 4: (4.8) is valid.
+Case 4: (18) is valid.
 Then, $\operatorname{cl}\,S_{A}^{0}\subseteq\operatorname{cl}(\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\}))=\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
-Recalling that (4.4) is valid, we conclude (4.1).
-Remark 4.3.
-The result in Proposition 4.1 (regarding assumption b)) is a generalization of a result by Kasimbeyli [45, Lem.
+Recalling that (14) is valid, we conclude (10). ∎
+Remark 14.
+Note that the equality “$=$” in (14) of Proposition 19 is essential and can not be replaced by $\subseteq$”.
+In the left-hand side of Figure 1, conditions (11), (12), (14), (15) are valid and the expected conclusion given by (10) holds.
+In the right-hand side of the Figure 1, the conditions (11), (12), (15) are valid, there is no $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that condition (14) is valid, but there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that
+$\operatorname{cl}\,S_{-K}(\psi)\subsetneq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$
+holds.
+In this case, the conclusion given by (10) fails.
+This example encourages studying equivalent norms in the contexts of cone separation (see the results in this section) and scalarization (see Section 4.4).
+${\rm cl}\,S_{A}(||\cdot||_{2})$$A$$A$$-K$${\rm cl}\,S_{-K}(||\cdot||_{2})$$A$$A$$-K$$\{x\in\mathbb{R}^{2}\mid||x||_{2}=1\}$$\{x\in\mathbb{R}^{2}\mid||x||_{1}=1\}$$0$$\{x\in\mathbb{R}^{2}\mid x^{*}(x)=\alpha\}$${\rm cl}\,S_{A}(||\cdot||_{1})$${\rm cl}\,S_{-K}(||\cdot||_{1})$$\{x\in\mathbb{R}^{2}\mid x^{*}(x)=\alpha\}$
+Figure 1: Examples given in Remark 14.
+Remark 15.
+The result in Proposition 19 (regarding assumption b)) is a generalization of a result by Kasimbeyli [9, Lem.
 3], where the author studied the Euclidean space $Y=\operatorname{\mathbb{R}}^{n}$, a closed cone $A\subseteq Y$ and a closed convex cone $K\subseteq Y$ with $A\cap(-K)=\{0\}$.
-Moreover, it is assumed that there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid (hence $0\notin\operatorname{cl}\,S_{-K}$ is valid).
-The conclusion in [45, Lem.
-3] is given as the condition (4.1) with $\operatorname{bd}\,A$ in the role of $A$.
-The condition (4.4) involved in b) of Proposition 4.1 can be fulfilled by certain Bishop-Phelps type cones, as the following lemma shows.
-Lemma 4.4.
+Moreover, it is assumed that there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (14) is valid (hence $0\notin\operatorname{cl}\,S_{-K}$ is valid).
+The conclusion in [9, Lem.
+3] is given as the separation condition (13) (i.e., (10) with $\operatorname{bd}\,A$ in the role of $A$).
+The condition (14) involved in b) of Proposition 19 can be fulfilled by certain Bishop-Phelps type cones, as the following lemma shows.
+Lemma 20.
 Consider a real topological-linear space $Y$, a continuous seminorm $\psi:Y\to\mathbb{R}$, and a nontrivial Bishop-Phelps type cone $C=C_{\psi}(x^{*},\alpha)$ for some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$.
@@
 and if
-$\{e\in Y\setminus\{0\}\mid\psi(e)=1,x^{*}(e)=0\}\neq\emptyset$ (4.9)
+$\{e\in Y\setminus\{0\}\mid\psi(e)=1,x^{*}(e)=0\}\neq\emptyset$ (19)
 (e.g., if $Y$ has dimension two or higher, and $\psi$ is a norm), then
 $\operatorname{cl}\,S_{C}(\psi)=S_{C}(\psi)=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}.$
-Proof 4.5.
+Proof.
 The equality $B_{C}=\{x\in\operatorname{\mathbb{S}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ is obvious.
@@
 Put $\gamma:=x^{*}(d)\,(\geq\alpha)$.
-By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$.
+By our assumption (19), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$.
 Then, $x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma$ for all $t\geq 0$, and $\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty$ for $t\to\infty$.
@@
 Since $x^{*}(d+t_{+}e)=x^{*}(d-t_{-}e)=\gamma\geq\alpha$ we infer $d-t_{-}e\in B_{C}$ and $d+t_{+}e\in B_{C}$, hence $d\in\operatorname{conv}(\{d-t_{-}e,\,d+t_{+}e\})\subseteq S_{C}.$
-Of course, $S_{C}=D$ and the closedness of $D$ imply $\operatorname{cl}\,S_{C}=S_{C}$.
-Remark 4.6.
-Kasimbeyli and Kasimbeyli [44, Th.
-4.5] stated a similar result as in Lemma 4.4 in a reflexive Banach space setting (with $\psi(\cdot):=||\cdot||$), but note that the assumption given in (4.9), which in this setting means that $Y$ has dimension two or higher, was not considered in [44, Th.
-4.5].
-In the one dimensional space ($Y:=\mathbb{R}$, ||$\cdot$|| ⋅ |), the condition (4.9) fails and the conclusion of Lemma 4.4 may fail.
+Of course, $S_{C}=D$ and the closedness of $D$ imply $\operatorname{cl}\,S_{C}=S_{C}$. ∎
+Remark 16.
+Kasimbeyli and Kasimbeyli [51, Th.
+4.5] stated a similar result as in Lemma 20 in a reflexive Banach space setting (with $\psi(\cdot):=||\cdot||$).
+Note that the assumption given in (19), which in this setting means that $Y$ has dimension two or higher, is essential.
+In the one dimensional space ($Y:=\mathbb{R}$, ||$\cdot$|| ⋅ |), the condition (19) fails and the conclusion of Lemma 20 may fail.
 To see this, consider $Y:=\mathbb{R}$, $x^{*}:=1$ and $\alpha\in(0,1]$.
 Then, $C=C(x^{*},\alpha)=\mathbb{R}_{+}$ and $B_{C}(|\cdot|)=\{1\}$ as well as $\operatorname{cl}\,S_{C}(|\cdot|)=\{1\}$ while $\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}=[\alpha,1]$.
-To prove some more specific results in the normed space setting in the upcoming Proposition 4.11 and Corollary 5.30 (see Section 5.3), we need additional lemmata.
-Lemma 4.7.
+To prove some more specific results in the normed space setting in the upcoming Proposition 23 and Corollary 33 (see Section 4.4), we need additional lemmata.
+Lemma 21.
 Consider a real normed space $(Y,||\cdot||)$, a (weakly) compact set $\Omega_{1}\subseteq Y$ with $0\notin\Omega_{1}$, and a (weakly) closed, bounded set $\Omega_{2}\subseteq Y$ with $\Omega_{2}\subseteq\operatorname{\mathbb{R}}_{+}\cdot\Omega_{1}$.
 Then, $\Omega_{2}$ is (weakly) compact.
-Proof 4.8.
+Proof.
 Note that a set $\Omega_{2}$ in a normed space is weakly bounded if and only if it is (norm) bounded. First, we prove that there exists $t>0$ such that
-$\Omega_{2}\subseteq[0,t]\cdot\Omega_{1}.$ (4.10)
+$\Omega_{2}\subseteq[0,t]\cdot\Omega_{1}.$ (20)
 On the contrary, assume that there are sequences $(x_{n}^{1})$, $(x_{n}^{2})$ and $(t_{n})$ with $x_{n}^{1}\in\Omega_{1}$, $x_{n}^{2}\in\Omega_{2}$, $t_{n}\in\mathbb{P}$ such that $x_{n}^{2}=t_{n}\cdot x_{n}^{1}$ with $t_{n}\to\infty$ for $n\to\infty$.
@@
 Since $[0,t]\times\Omega_{1}$ is (weakly) compact in ×$\operatorname{\mathbb{R}}\times Y$, and the function $h$ is (weakly) continuous by the topological vector space properties, we get that $[0,t]\cdot\Omega_{1}$ is (weakly) compact.
-From (4.10) we derive that the (weakly) closed subset $\Omega_{2}$ of $[0,t]\cdot\Omega_{1}$ is (weakly) compact as well.
-Lemma 4.9.
+From (20) we derive that the (weakly) closed subset $\Omega_{2}$ of $[0,t]\cdot\Omega_{1}$ is (weakly) compact as well.
+∎
+Lemma 22.
 Consider a real normed space $(Y,||\cdot||)$ and a nontrivial cone $C\subseteq Y$. Then, the following assertions are valid:
-  • $1^{\circ}$.
+  • $1^{\circ}$
     If $B_{C}(||\cdot||)$ is compact, then $B_{C}(||\cdot||^{\prime})$ is compact for all $||\cdot||^{\prime}\sim||\cdot||$.
-  • $2^{\circ}$.
+  • $2^{\circ}$
     If $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, then $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is weakly compact for all $||\cdot||^{\prime}\sim||\cdot||$.
-  • $3^{\circ}$.
+  • $3^{\circ}$
     If $\operatorname{cl}\,S_{C}(||\cdot||)$ is (weakly) compact with $0\notin\operatorname{cl}\,S_{C}(||\cdot||)$, then $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is (weakly) compact for all $||\cdot||^{\prime}\sim||\cdot||$.
-  • $4^{\circ}$.
+  • $4^{\circ}$
     If $Y$ is reflexive (or equivalently, $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact), then $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact for all $||\cdot||^{\prime}\sim||\cdot||$.
-Proof 4.10.
+Proof.
 Consider any $||\cdot||^{\prime}\sim||\cdot||$.
-$1^{\circ}$.
-Obviously, $B_{C}(||\cdot||^{\prime})\subseteq C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)$.
+$1^{\circ}$ Obviously, $B_{C}(||\cdot||^{\prime})\subseteq C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)$.
 Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $B_{C}(||\cdot||^{\prime})$ is bounded.
-Moreover, according to [23, Lemma 2.1.43 (iii)], for the bounded and closed norm-base $B_{C}(||\cdot||)$ with $0\notin B_{C}(||\cdot||)$ we have $C=\mathbb{R}_{+}\cdot B_{C}(||\cdot||)$ is closed, hence $B_{C}(||\cdot||^{\prime})$ is closed as well.
-Then, the compactness of $B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the compact set $\Omega_{1}:=B_{C}(||\cdot||)$ and the closed, bounded set $\Omega_{2}:=B_{C}(||\cdot||^{\prime})$.
+Moreover, according to [21, Lemma 2.1.43 (iii)], for the bounded and closed norm-base $B_{C}(||\cdot||)$ with $0\notin B_{C}(||\cdot||)$ we have $C=\mathbb{R}_{+}\cdot B_{C}(||\cdot||)$ is closed, hence $B_{C}(||\cdot||^{\prime})$ is closed as well.
+Then, the compactness of $B_{C}(||\cdot||^{\prime})$ follows from Lemma 21 applied for the compact set $\Omega_{1}:=B_{C}(||\cdot||)$ and the closed, bounded set $\Omega_{2}:=B_{C}(||\cdot||^{\prime})$.
 Note that $0\notin B_{C}(||\cdot||)=\Omega_{1}$.
-$2^{\circ}$.
-Since $C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)\subseteq\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)\subseteq\operatorname{cl}_{w}\,C$, we have $\operatorname{cl}_{w}\,C=\operatorname{cl}_{w}(\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||))$.
-Taking into account that $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is a weakly closed, bounded set with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, applying [23, Lemma 2.1.43 (iii)] we conclude that $\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly closed (remark that the weak topology of the normed space $Y$ is Hausdorff).
+$2^{\circ}$ Since $C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)\subseteq\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)\subseteq\operatorname{cl}_{w}\,C$, we have $\operatorname{cl}_{w}\,C=\operatorname{cl}_{w}(\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||))$.
+Taking into account that $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is a weakly closed, bounded set with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, applying [21, Lemma 2.1.43 (iii)] we conclude that $\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly closed (remark that the weak topology of the normed space $Y$ is Hausdorff).
 Consequently, $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}_{w}\,C=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$.
 Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is bounded.
-Then, the weak compactness of $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the weakly compact set $\Omega_{1}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ and the weakly closed, bounded set $\Omega_{2}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$.
-$3^{\circ}$.
-We have $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}(\operatorname{conv}\,C)=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}\,S_{C}(||\cdot||)$, where the latter equality is discussed in [29, Rem.
+Then, the weak compactness of $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ follows from Lemma 21 applied for the weakly compact set $\Omega_{1}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ and the weakly closed, bounded set $\Omega_{2}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$.
+$3^{\circ}$ We have $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}(\operatorname{conv}\,C)=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}\,S_{C}(||\cdot||)$, where the latter equality is discussed in [15, Rem.
 2.5].
 Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence the set $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is bounded.
-Finally, the (weak) compactness of $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the (weakly) compact set $\Omega_{1}:=\operatorname{cl}\,S_{C}(||\cdot||)$ and the (weakly) closed, bounded set $\Omega_{2}:=\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$.
-$4^{\circ}$.
-Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence both weakly closed sets $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact if $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact.
-To later prove a Bishop-Phelps type scalarization result (Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
-Proposition 4.11.
+Finally, the (weak) compactness of $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ follows from Lemma 21 applied for the (weakly) compact set $\Omega_{1}:=\operatorname{cl}\,S_{C}(||\cdot||)$ and the (weakly) closed, bounded set $\Omega_{2}:=\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$.
+$4^{\circ}$ Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence both weakly closed sets $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact if $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact.
+∎
+To later prove a Bishop-Phelps type scalarization result (Corollary 33) for vector optimization problems in real normed spaces, we need the following proposition.
+Proposition 23.
 Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$.
-Suppose that (4.2) and (4.3) with $\psi(\cdot):=||\cdot||$ are valid.
+Suppose that (11) and (12) with $\psi(\cdot):=||\cdot||$ are valid.
 Moreover, assume that one of the following statements is valid:
@@
     $A$ is weakly closed, and either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
-Then, there is $||\cdot||^{\prime}\sim||\cdot||$ such that (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
-Proof 4.12.
-Suppose that (4.2) and (4.3) (with $\psi(\cdot):=||\cdot||$) are valid.
-First, note that the following assertions are known to be equivalent (see [14, Lem.
-1.25 and 1.26], [29, Lem.
-3.4] and [54, Th.
+Then, there is $||\cdot||^{\prime}\sim||\cdot||$ such that (10) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
+Proof.
+Suppose that (11) and (12) (with $\psi(\cdot):=||\cdot||$) are valid.
+First, note that the following assertions are known to be equivalent (see [54, Lem.
+1.25 and 1.26], [15, Lem.
+3.4] and [55, Th.
 3.2]):
-  • $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ (i.e., (4.3) with $\psi(\cdot):=||\cdot||$ is valid);
+  • $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ (i.e., (12) with $\psi(\cdot):=||\cdot||$ is valid);
   • $0\notin\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||)$;
@@
   • $-\operatorname{cl}(\operatorname{conv}\,K)$ is well-based;
-  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is representable as a Bishop-Phelps cone (cf. Example 2.4).
-Thus, under assumption (4.3) (with $\psi(\cdot):=||\cdot||$), there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $-\operatorname{cl}(\operatorname{conv}\,K)=C_{||\cdot||^{\prime}}(x^{*},\alpha)$.
-By Lemma 4.4, in the real normed space $(Y,||\cdot||^{\prime})$ with dimension two or higher we have $\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\{x\in\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\mid x^{*}(x)\geq\alpha\}$, hence (4.3) and (4.4) with $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ are valid.
+  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is representable as a Bishop-Phelps cone (cf. Example 1).
+Thus, under assumption (12) (with $\psi(\cdot):=||\cdot||$), there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $-\operatorname{cl}(\operatorname{conv}\,K)=C_{||\cdot||^{\prime}}(x^{*},\alpha)$.
+By Lemma 20, in the real normed space $(Y,||\cdot||^{\prime})$ with dimension two or higher we have $\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\{x\in\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\mid x^{*}(x)\geq\alpha\}$, hence (12) and (14) with $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ are valid.
 Let us finish the proof by assuming a) and b), respectively:
-a) In view of Lemma 4.9 ($1^{\circ}$), $B_{A}(||\cdot||^{\prime})$ is compact if $B_{A}(||\cdot||)$ is compact.
-Proposition 4.1 b) applied for the space $(Y,||\cdot||^{\prime})$, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures $(\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap(\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime}))=\emptyset$, hence (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid, taking into account that $\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
-b) In view of Lemma 4.9 ($2^{\circ}$, $4^{\circ}$), $\operatorname{cl}_{w}\,B_{A}(||\cdot||^{\prime})$ is weakly compact if either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
-Proposition 4.1 b) applied for the space $Y$ endowed with the weak topology, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$, taking into account that $\operatorname{cl}_{w}\,S_{A}^{0}(||\cdot||^{\prime})=\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime})$ and $\operatorname{cl}_{w}\,S_{-\operatorname{cl}_{w}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
+a) In view of Lemma 22 ($1^{\circ}$), $B_{A}(||\cdot||^{\prime})$ is compact if $B_{A}(||\cdot||)$ is compact.
+Proposition 19 b) applied for the space $(Y,||\cdot||^{\prime})$, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures $(\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap(\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime}))=\emptyset$, hence (10) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid, taking into account that $\operatorname{cl}\,S_{{-\rm cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
+b) In view of Lemma 22 ($2^{\circ}$, $4^{\circ}$), $\operatorname{cl}_{w}\,B_{A}(||\cdot||^{\prime})$ is weakly compact if either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
+Proposition 19 b) applied for the space $Y$ endowed with the weak topology, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures (10) with $\psi(\cdot):=||\cdot||^{\prime}$, taking into account that $\operatorname{cl}_{w}\,S_{A}^{0}(||\cdot||^{\prime})=\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime})$ and $\operatorname{cl}_{w}\,S_{-\operatorname{cl}_{w}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\operatorname{cl}\,S_{{-\rm cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
+∎
 Furthermore, we recall some recent results on cone separation in real topological-linear (locally convex) spaces established in [14].
-Proposition 4.13 ([28, Th. 5.4]).
+Proposition 24 ([14, Th. 5.4]).
 Assume that $Y$ is a real topological-linear space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$, and $K$ is solid.
@@
 or equivalently,
-$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad K\subseteq C_{\psi}(x^{*},\alpha)\quad$ (4.11)
+$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad K\subseteq C_{\psi}(x^{*},\alpha)\quad$ (21)
 $\text{and}\quad\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha).$
-Proposition 4.14 ([28, Th. 5.8]).
+Proposition 25 ([14, Th. 5.8]).
 Assume that $Y$ is a real topological-linear space, $K,A\subseteq Y$ are nontrivial, convex cones with normlike-bases $B_{K}$ and $B_{A}$, and $K$ is solid.
@@
 Then, the following assertions are equivalent:
-  • $1^{\circ}$.
+  • $1^{\circ}$
     $S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset$.
-  • $2^{\circ}$.
+  • $2^{\circ}$
     $A\cap\operatorname{int}(-K)=\emptyset$.
-  • $3^{\circ}$.
-    There exists $(x^{*},\alpha)\in K^{a\circ}$ such that (4.11) is valid.
-Proposition 4.15 ([28, Th. 5.9]).
+  • $3^{\circ}$
+    There exists $(x^{*},\alpha)\in K^{a\circ}$ such that (21) is valid.
+Proposition 26 ([14, Th. 5.9]).
 Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$.
-Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.1) is valid.
+Suppose that one of the sets ${\rm cl}\,S_{-K}$ and ${\rm cl}\,S_{A}^{0}$ is compact, and (10) is valid.
 Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
@@
 or equivalently,
-$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$ (4.12)
-From the general strict cone separation result in Proposition 4.15 and our Propositions 4.1 and 4.11 we immediately derive some new variants of strict cone separation results involving the classical separation condition (4.2).
-Corollary 4.16.
+$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha).$ (22)
+From the general strict cone separation result in Proposition 26 and our Propositions 19 and 23 we immediately derive some new variants of strict cone separation results involving the classical separation condition (11).
+Corollary 27.
 Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$.
-Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.2) and (4.3) are valid.
-Moreover, assume that the assumptions in a) or in b) of Proposition 4.1 are satisfied.
-Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that (LABEL:sep_conditions_strict) is valid.
-Corollary 4.17.
+Suppose that one of the sets ${\rm cl}\,S_{-K}$ and ${\rm cl}\,S_{A}^{0}$ is compact, and (11) and (12) are valid.
+Moreover, assume that the assumptions in a) or in b) of Proposition 19 are satisfied.
+Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that (22) is valid.
+Corollary 28.
 Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$.
-Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ and ${\mathrm{c}l}\,S_{A}^{0}(||\cdot||)$ is weakly compact (e.g., if $Y$ is reflexive), and (4.2) and (4.3) with $\psi(\cdot):=||\cdot||$ are valid.
-Moreover, assume that the assumptions in a) or in b) of Proposition 4.11 are satisfied.
-Then, there are $||\cdot||^{\prime}\sim||\cdot||$ and $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that (LABEL:sep_conditions_strict) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
-5 Bishop-Phelps Type Scalarization in Vector Optimization
-Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$.
-Then, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, and any given reference point $a\in Y$, one defines $\varphi_{x^{*},\alpha}^{a}:Y\to\operatorname{\mathbb{R}}$ as $\varphi_{x^{*},\alpha}^{a}(y):=\varphi_{x^{*},\alpha}(y-a)$ for all $y\in Y$.
-The following parameterized scalar optimization problem of the vector problem ($P$) (in the sense of Kasimbeyli [45]) will be of interest:
-$\begin{cases}(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\varphi_{x^{*},\alpha}(f(x)-a)=x^{*}(f(x)-a)+\alpha\psi(f(x)-a)\to\min\\ x\in\Omega.\end{cases}$ ($P_{\varphi}^{a}$)
-In addition, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$ and any given reference point $a\in Y$ and direction $k\in Y\setminus\{0\}$, we are interested in the parameterized scalar problem (in the sense of Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]):
-$\begin{cases}\lambda\to\min\\ \varphi_{x^{*},\alpha}(f(x)-a-\lambda k)\leq 0,\\ (x,\lambda)\in\Omega\times\mathbb{R}.\end{cases}$ ($P_{\varphi}^{a,k}$)
-5.1 Basic scalarization results
-First, we present monotonicity properties of the seminorm-linear function $\varphi_{x^{*},\alpha}$ with respect to a Bishop-Phelps type cone $C_{\psi}(x^{*},\alpha)$.
-Lemma 5.1.
-The following assertions hold:
-  • $1^{\circ}$.
-    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, then $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing.
-  • $2^{\circ}$.
-    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
-  • $3^{\circ}$.
-    If $K$ is solid and $(x^{*},\alpha)\in K^{a\circ}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
-  • $4^{\circ}$.
-    If $(x^{*},\alpha)\in K^{a\#}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
-  • $5^{\circ}$.
-    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, then $\varphi_{x^{*},\alpha}$ is strongly $(C_{\psi}^{>}(x^{*},\alpha)\cup\{0\})$-increasing.
-Proof 5.2.
-$1^{\circ}$ Noting that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing, the assertion follows from Lemma 3.2 ($1^{\circ}$).
-$2^{\circ}$ If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$, there is nothing to prove.
-If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, using the assumption $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ we have that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing, and the result follows from Lemma 3.2 ($1^{\circ}$).
-$3^{\circ}$ and $4^{\circ}$ Using the assumptions it follows that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, so by assertion $2^{\circ}$ we get the conclusion.
-$5^{\circ}$ It follows from the definitions.
-The next lemma, which studies the monotonicity properties of $\varphi_{x^{*},\alpha}$ with respect to the original cone $K$, extends a corresponding result derived by Kasimbeyli [45, Th.
-3.5] in a normed setting.
-Lemma 5.3.
-Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
-  • $1^{\circ}$.
-    $(x^{*},\alpha)\in K^{a+}$ $\iff$ $\varphi_{x^{*},\alpha}$ is $K$-increasing.
-  • $2^{\circ}$.
-    $(x^{*},\alpha)\in K^{a\circ}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing.
-  • $3^{\circ}$.
-    $(x^{*},\alpha)\in K^{a\#}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing.
-Proof 5.4.
-$1^{\circ}$ By Lemma 5.1 ($1^{\circ}$), the function $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing.
-If $(x^{*},\alpha)\in K^{a+}$, then $K\subseteq C_{\psi}(x^{*},\alpha)$, hence $\varphi_{x^{*},\alpha}$ is $K$-increasing by Lemma 3.2 ($3^{\circ}$).
-Conversely, assuming $\varphi_{x^{*},\alpha}$ is $K$-increasing, for $k\in K$ we have $-k\leq_{K}0$, hence $\varphi_{x^{*},\alpha}(-k)\leq\varphi_{x^{*},\alpha}(0)=0$, which means that $x^{*}(k)-\alpha\psi(k)\geq 0$.
-Noting that $x^{*}\in K^{+}$, we conclude $(x^{*},\alpha)\in K^{a+}$.
-$2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid.
-Assume that $\operatorname{int}\,K\neq\emptyset$.
-If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 5.1 ($2^{\circ}$) and 3.2 ($2^{\circ}$).
-The proof of the reverse implication is similar to $1^{\circ}$.
-$3^{\circ}$ If $(x^{*},\alpha)\in K^{a\#}$, then $K\subseteq C_{\psi}^{>}(x^{*},\alpha)\cup\{0\}$, so $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing by Lemmata 5.1 ($5^{\circ}$) and 3.2 ($3^{\circ}$).
-Also here the proof of the reverse implication is similar to $1^{\circ}$.
-Now, we are able to present first relationships between the parameterized scalar problem ($P_{\varphi}^{a}$) and the vector problem ($P$).
-Theorem 5.5.
-For any $a\in Y$, the following assertions hold:
-  • $1^{\circ}$.
-    If $(x^{*},\alpha)\in K^{a\circ}$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
-  • $2^{\circ}$.
-    If $(x^{*},\alpha)\in K^{a\#}$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
-  • $3^{\circ}$.
-    Assume that $\overline{x}\in\Omega$.
-If $(x^{*},\alpha)\in K^{a+}$ and ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\{\overline{x}\}$, then $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
-  • $4^{\circ}$.
-    If $(x^{*},\alpha)\in K^{a\#}$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
-Proof 5.6.
-$1^{\circ}$ Consider $(x^{*},\alpha)\in K^{a\circ}$.
-By Lemma 5.3 ($2^{\circ}$) $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing.
-It is easy to check that $\varphi_{x^{*},\alpha}^{a}$ is strictly $K$-increasing for any $a\in Y$ as well.
-Finally, Lemma 3.4 ($1^{\circ}$) yields the conclusion in assertion $1^{\circ}$.
-$2^{\circ}$ and $3^{\circ}$ The proof of the assertions is similar to the proof of assertion $1^{\circ}$ by using Lemmata 5.3 and 3.4.
-$4^{\circ}$ Take some $(x^{*},\alpha)\in K^{a\#}$.
-By Lemma 5.1 ($4^{\circ}$) $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing, and so $\varphi_{x^{*},\alpha}^{a}$ as well.
-Because $(x^{*},\alpha)\in K^{a\#}$ we get $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
-Taking into account that $C_{\psi}(x^{*},\alpha)$ is a nontrivial convex cone, we conclude that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$.
-Finally, with the aid of Lemma 3.4 ($4^{\circ}$) we get the desired inclusion.
-Remark 5.7.
-Consider $\bar{x}\in\Omega$.
-In some upcoming results, the specific choice $a:=f(\bar{x})$ will be of interest.
-In this case, the scalarized objective function $\varphi_{x^{*},\alpha}^{a}\circ f$ is given by $(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$ for all $x\in\Omega$.
-According to Theorem 5.5, for a specific choice of the pair $(x^{*},\alpha)$ one can ensure the (weak, proper) efficiency of $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
-Let us shift our focus to relationships between the parameterized scalar problem ($P_{\varphi}^{a,k}$) and the vector problem ($P$) involving a Bishop-Phelps type cone.
-Lemma 5.8.
-Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions are valid:
-  • $1^{\circ}$.
-    If $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
-More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$).
-  • $2^{\circ}$.
-    If $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
-  • $3^{\circ}$.
-    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$.
-If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-  • $4^{\circ}$.
-    If $(\bar{x},\bar{\lambda})$ solves the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Proof 5.9.
-Follows immediately from Lemma 3.8 taking into account that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing.
-Theorem 5.10.
-Consider $\bar{x}\in\Omega$.
-Take some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, $a\in Y,k\in Y\setminus\{0\}$ and assume that $(\bar{x},\bar{\lambda})$ is a solution of the optimization problem ($P_{\varphi}^{a,k}$).
-Then:
-  • $1^{\circ}$.
-    If $(x^{*},\alpha)\in K^{a\#}$ and $C_{\psi}(x^{*},\alpha)$ is solid, then $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
-  • $2^{\circ}$.
-    If $(x^{*},\alpha)\in K^{a\circ}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
-Proof 5.11.
-$1^{\circ}$.
-Remark that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in view of Lemma 5.8 ($4^{\circ}$).
-The latter set is contained in ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)$ (since $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$).
-$2^{\circ}$.
-For the case $\operatorname{int}\,K=\emptyset$ the statement is clear.
-Assume that $\operatorname{int}\,K\neq\emptyset$.
-Again we have $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Here the latter set is contained in ${\mathrm{W}Eff}(\Omega\mid f,K)$ (since $\emptyset\neq\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
-Let us summarize our findings related to vector problems involving a Bishop-Phelps type cone in the next two propositions.
-Proposition 5.12.
-Let $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
-  • $1^{\circ}$.
-    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-  • $2^{\circ}$.
-    For any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$, the set $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
-  • $3^{\circ}$.
-    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
-If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}-3^{\circ}$ are equivalent to:
-If further $C_{\psi}(x^{*},\alpha)$ is closed, then $1^{\circ}-4^{\circ}$ are equivalent to:
-  • $5^{\circ}$.
-    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
-Proof 5.13.
-Note that $C_{\psi}(x^{*},\alpha)$ is a convex cone, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi_{x^{*},\alpha}(0)=0$ and $C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))\neq\emptyset$.
-$1^{\circ}\iff 2^{\circ}$. The equivalence follows immediately from Lemma 5.8 ($1^{\circ}$, $3^{\circ}$).
-$1^{\circ}\iff 3^{\circ}$. The result follows directly from Lemma 3.5 ($1^{\circ}$).
-$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$.
-The result follows from Proposition 3.11 taking into account that ($P_{\varphi}^{a,k}$) is the same problem as (3.1) with $C_{\psi}(x^{*},\alpha)$ in the role of $K$.
-Proposition 5.14.
-Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
-  • $1^{\circ}$.
-    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-  • $2^{\circ}$.
-    For any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, the point $(\bar{x},0)$ is a solution of the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
-If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
-  • $3^{\circ}$.
-    $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
-If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
-If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
-  • $5^{\circ}$.
-    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
-Proof 5.15.
-First, note that $C_{\psi}(x^{*},\alpha)$ is a solid, convex cone and $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi_{x^{*},\alpha}(0)=0$.
-$1^{\circ}\iff 2^{\circ}$. The equivalence follows immediately from Lemma 5.8 ($2^{\circ}$, $4^{\circ}$).
-$1^{\circ}\iff 3^{\circ}$.
-Since $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, the function $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$-representing by Lemma 2.12 ($4^{\circ}$) with $\varphi_{x^{*},\alpha}(0)=0$.
-Then, the result follows from Lemma 3.5 ($2^{\circ}$).
-$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$. The result follows directly from the Proposition 3.14.
-Remark 5.16.
-The results derived in Theorem 5.5 and Propositions 5.12 and 5.14 extend some known results in the literature, which are given in a normed space $(Y,\psi)$, $\psi(\cdot):=||\cdot||$, preordered by a convex cone $K\subseteq Y$:
-  • Gasimov [17, Th. 1] proved a corresponding statement as in Theorem 5.5 ($4^{\circ}$) for Benson proper efficiency (instead of Henig proper efficiency).
-  • Kasimbeyli [46, Th. 5.4] stated similar assertions as in Theorem 5.5 for a closed, pointed, convex cone $K$ and Benson proper efficiency (instead of Henig proper efficiency).
-  • Gasimov [17, Th.
-2] (see also Guo and Zhang [31, Th.
-4.5]) stated that $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$ (with $\mathbb{A}$ considered in Benson proper efficiency) implies $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
-Remark 5.17.
-It is worth noting that in Proposition 5.12 (respectively, in Proposition 5.14) the function $\varphi_{x^{*},\alpha}$, which appears in the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$), can be replaced by any $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$.
-For instance, the nonlinear scalarizing function $\varphi^{G}_{-C_{\psi}(x^{*},\alpha),k}$ (i.e., the function defined in (3.3) with $a=0$ and with $C_{\psi}(x^{*},\alpha)$ in the role of $K$) is $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) if $C_{\psi}(x^{*},\alpha)$ is nontrivial and closed, and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha)$ (respectively, $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
-Remark 5.18.
-Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$.
-Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
-If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g.
-if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so (see Proposition 2.23) $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
-If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence $\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
+Suppose that one of the sets ${\rm cl}\,S_{-K}(||\cdot||)$ and ${\rm cl}\,S_{A}^{0}(||\cdot||)$ is weakly compact (e.g., if $Y$ is reflexive), and (11) and (12) with $\psi(\cdot):=||\cdot||$ are valid.
+Moreover, assume that the assumptions in a) or in b) of Proposition 23 are satisfied.
+Then, there are $||\cdot||^{\prime}\sim||\cdot||$ and $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that (22) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
 4.3 Scalarization results for the concept of weak efficiency
@@
 More precisely, we will present conditions that ensure that a weakly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
-Theorem 5.19.
-Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid cone, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial.
+Theorem 29.
+Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid cone, $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot))$, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})$ is nontrivial.
 Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively.
 Suppose $S_{-K}$ is solid and
-$S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ (5.1)
+$S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ (23)
 Then, there is $(x^{*},\alpha)\in K^{a\circ}$ such that $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
-Proof 5.20.
-Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that $A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
+Moreover, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 18 is valid.
+Proof.
+Under our assumptions, by Proposition 24 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that $A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
 Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$ hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 5.14 yields the rest of the conclusion.
-Remark 5.21.
-In view of Remark 5.18, the point $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.19 belongs to ${\mathrm{W}Eff}(\Omega\mid f,K)$.
-In the convex case (i.e., $K$ and $\mathbb{A}(\bar{x})$ are convex cones), using Proposition 4.14 we are able to replace the separation condition (5.1) by $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, as shown in the following result.
-Corollary 5.22 (convex case).
-Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid, convex cone, $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial and convex.
+Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 18 yields the rest of the conclusion.
+∎
+In the convex case (i.e., $K$ and $\mathbb{A}(\bar{x})$ are convex cones), using Proposition 25 we are able to replace the separation condition (23) by $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$, as shown in the following result.
+Corollary 30 (convex case).
+Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid, convex cone, $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot))$, $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is nontrivial and convex.
 Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively, and $S_{K}$ is solid.
-Then, the conclusion of Theorem 5.19 holds.
-Proof 5.23.
-Follows from Theorem 5.19 and Proposition 4.14.
-Remark 5.24.
-Recall that, for any set $\Omega\subseteq Y,$ the quasi-interior of $\Omega$ is the set ${\mathrm{q}i}\,\Omega:=\{x\in\Omega\mid\operatorname{cl}(\mathbb{R}_{+}\cdot(\Omega-x))=Y\}$.
-In order to ensure that the assumption $A$ is nontrivial and convex from Corollary 5.22 is satisfied it is sufficient to impose that $Y$ is a real Hausdorff locally convex space and $f$ is $K$-convexlike (i.e., $f[\Omega]+K$ is convex).
-Indeed, assume that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
+Then, the conclusion of Theorem 29 holds.
+Proof.
+Follows from Theorem 29 and Proposition 25. ∎
+Remark 17.
+In view of Remark 12, since the point $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in the conclusion of Theorem 29 belongs to ${\rm WEff}(\Omega\mid f,K)$, the result provides sufficient conditions for weak efficiency.
+Under the assumptions of Corollary 30 (including convexity of both cones) and taking into account Remark 12, we derive the equivalence of the following assertions:
+  • $1^{\circ}$
+    $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$.
+  • $2^{\circ}$
+    There is $(x^{*},\alpha)\in K^{a\circ}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 17 is valid, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing).
+From this perspective, $2^{\circ}$ is a necessary and sufficient condition for the weak efficiency of $\bar{x}\in\Omega$.
+Remark 18.
+Recall that, for any set $\Omega\subseteq Y,$ the quasi-interior of $\Omega$ is the set ${\rm qi}\,\Omega:=\{x\in\Omega\mid\operatorname{cl}(\mathbb{R}_{+}\cdot(\Omega-x))=Y\}$.
+In order to ensure that the assumption $A$ is nontrivial and convex from Corollary 30 is satisfied it is sufficient to impose that $Y$ is a real Hausdorff locally convex space and $f$ is $K$-convexlike (i.e., $f[\Omega]+K$ is convex).
+Indeed, assume that $\bar{x}\in{\rm WEff}(\Omega\mid f,K)$.
 Since $f$ is $K$-convexlike, we infer that $A=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is a convex cone in $Y$.
-By [27, Lem.
+By [52, Lem.
 3.3 (4)] we know that $f(\bar{x})\notin\operatorname{int}(f[\Omega]+K)$.
-Since the interior and the quasi interior of a solid, convex set $f[\Omega]+K$ in a real Hausdorff locally convex space coincide (see [23, p.
-66-67]), we have $A\subseteq{\mathrm{c}l}(\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x})))\neq Y$.
+Since the interior and the quasi interior of a solid, convex set $f[\Omega]+K$ in a real Hausdorff locally convex space coincide (see [21, p.
+66-67]), we have $A\subseteq{\rm cl}(\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x})))\neq Y$.
 Moreover, since $K\neq\{0\}$ we have $A\neq\{0\}$ as well, hence $A$ is nontrivial.
+Remark 19.
+In this and upcoming sections, we consider a seminorm $\psi$ which is positive on $K\setminus\{0\}$ and on $\mathbb{A}(\bar{x})\setminus\{0\}$.
+In Example 2(a,b) we discussed ways to construct such seminorms.
+Of course, changing the $\bar{x}$ may alter the cone $\mathbb{A}(\bar{x})$, so one may require another seminorm $\psi$ (at least if $\psi$ is not a norm).
+In general, one must be careful when choosing a seminorm.
+Ideally, one would want a seminorm $\psi$ that is positive on $K\setminus\{0\}$ and on $\mathbb{A}(\bar{x})\setminus\{0\}$ for all points $\bar{x}\in\Omega$ which are relevant in the results (e.g.
+$\bar{x}\in{\rm WEff}(\Omega\mid f,K)$ in Corollary 30).
+This is because the scalarized problems appearing in the conclusions of the results are then defined with respect to the same seminorm $\psi$.
+Note that any norm has the desired positivity property since it is positive on $Y\setminus\{0\}$.
 4.4 Scalarization results for the concept of proper efficiency with respect to a cone-valued map
 We are ready to present some further main scalarization results which ensure that an $\mathbb{A}$-properly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
-Theorem 5.25.
+Remember that $\mathbb{A}:\Omega\to 2^{Y}$ is a cone-valued map.
+Theorem 31.
 Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})$ is a nontrivial cone in $Y$ with $f[\Omega]-f(\bar{x})\subseteq A$.
@@
 Suppose that one of the sets $\operatorname{cl}\,S_{-K}$ and $\operatorname{cl}\,S_{A}^{0}$ is compact, and
-$({\mathrm{c}l}\,S_{A}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$ (5.2)
+$({\rm cl}\,S_{A}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$ (24)
 Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Moreover, any of the equivalent assertions given in Proposition 5.12 is valid.
-Proof 5.26.
-By Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $A\cap-C_{\psi}(x^{*},\alpha)=\{0\}$ and $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$.
+Moreover, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 17 is valid.
+Proof.
+By Proposition 26 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $A\cap-C_{\psi}(x^{*},\alpha)=\{0\}$ and $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$.
 Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$ hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Finally, observing that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
-Remark 5.27.
-In view of Remark 5.18, if $K$ is solid or $\psi$ is continuous, then the point $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.25 belongs to ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
-Using Proposition 4.1 we are able to replace the separation condition (5.2) by $0\notin\operatorname{cl}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ as well as some more specific conditions that we formalize below.
-Corollary 5.28.
-Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
-Suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, and that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact.
+Finally, observing that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 1 ($7^{\circ}$), Proposition 17 yields the rest of the conclusion.
+∎
+Remark 20.
+In view of Remark 12, if $K$ is solid or $\psi$ is continuous, then the point $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in the conclusion of Theorem 31 belongs to ${\rm PEff}_{He}(\Omega\mid f,K)$, hence the result provides sufficient conditions for proper efficiency in the sense of Henig.
+Using Proposition 19 we are able to replace the separation condition (24) by $0\notin\operatorname{cl}\,S_{-K}$ and $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$ as well as some more specific conditions that we formalize below.
+Corollary 32.
+Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}$, $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
+Suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, and that one of the sets ${\rm cl}\,S_{-K}$ and ${\rm cl}\,S_{A}^{0}$ is compact.
 Moreover, assume that one of the following statements is valid:
@@
   2. b)
-    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid, and one of the conditions (4.5)–(4.8) holds.
-Then, the conclusion of Theorem 5.25 is valid.
-Proof 5.29.
-Using Proposition 4.1 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (5.2).
+    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (14) is valid, and one of the conditions (15)–(18) holds.
+Then, the conclusion of Theorem 31 is valid.
+Proof.
+Using Proposition 19 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (24).
 Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$.
-Then, the result follows from Theorem 5.25.
-In the remaining part of the section, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.
-Corollary 5.30.
-Assume that ($Y,||\cdot||$) is a real normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
+Then, the result follows from Theorem 31.
+∎
+In the remaining part of the section, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 23 and Lemma 22.
+Corollary 33.
+Assume that ($Y,||\cdot||$) is a real normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$, $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
 Moreover, assume that one of the following statements is valid:
   1. a)
-    $B_{A}(||\cdot||)$ and ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ are compact.
+    $B_{A}(||\cdot||)$ and ${\rm cl}\,S_{-K}(||\cdot||)$ are compact.
   2. b)
-    $A$ is weakly closed, and either $Y$ is reflexive or ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ and $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ are weakly compact and $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
+    $A$ is weakly closed, and either $Y$ is reflexive or ${\rm cl}\,S_{-K}(||\cdot||)$ and $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ are weakly compact and $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
 Then, there is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$.
-Moreover, any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid.
-Proof 5.31.
+Moreover, if $\varphi$ is $-C_{||\cdot||^{\prime}}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 17, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid.
+Proof.
 Note that $\varphi_{x^{*},\alpha}$ from the conclusion is constructed with respect to the equivalent norm $||\cdot||^{\prime}$.
-Since $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ and $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$, using Proposition 4.11 a) and b) respectively, under our standing assumptions a) and b), one gets that $({\mathrm{c}l}\,S_{A}^{0}(||\cdot||^{\prime}))\cap({\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime}))=\emptyset$ for some norm $||\cdot||^{\prime}\sim||\cdot||$.
-Applying Lemma 4.9 ($3^{\circ}$, $4^{\circ}$), it follows that ${\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime})$ is compact under a), respectively, weakly compact under b).
-Consequently, the conclusion follows by applying Theorem 5.25 for $Y$ endowed with the norm $\psi(\cdot):=||\cdot||^{\prime}$ under a), respectively, endowed with the weak topology and $\psi(\cdot):=||\cdot||^{\prime}$ under b).
-For the finite dimensional case, we obtain from Corollary 5.30 the following result.
-Corollary 5.32.
-Assume that ($Y,||\cdot||$) is a real finite dimensional normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, pointed, convex cone, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a closed cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
-Then, the conclusion of Corollary 5.30 holds.
-Proof 5.33.
-By the well-known Krein-Rutman theorem (see, e.g., [49, Th.
-2.4.7]) we know that $K^{\#}\neq\emptyset$, and by [29, Th.
+Since $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ and $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$, using Proposition 23 a) and b) respectively, under our standing assumptions a) and b), one gets that $({\rm cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap({\rm cl}\,S_{-K}(||\cdot||^{\prime}))=\emptyset$ for some norm $||\cdot||^{\prime}\sim||\cdot||$.
+Applying Lemma 22 ($3^{\circ}$, $4^{\circ}$), it follows that ${\rm cl}\,S_{-K}(||\cdot||^{\prime})$ is compact under a), respectively, weakly compact under b).
+Consequently, the conclusion follows by applying Theorem 31 for $Y$ endowed with the norm $\psi(\cdot):=||\cdot||^{\prime}$ under a), respectively, endowed with the weak topology and $\psi(\cdot):=||\cdot||^{\prime}$ under b).
+∎
+For the finite dimensional case, we obtain from Corollary 33 the following result.
+Corollary 34.
+Assume that ($Y,||\cdot||$) is a real finite dimensional normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, pointed, convex cone, $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a closed cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
+Then, the conclusion of Corollary 33 holds.
+Proof.
+By the well-known Krein-Rutman theorem (see, e.g., [42, Th.
+2.4.7]) we know that $K^{\#}\neq\emptyset$, and by [15, Th.
 3.1 ($4^{\circ}$)] it follows that $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$.
-So, the assumptions of Corollary 5.30 are satisfied.
-Remark 5.34.
-Besides the proper efficiency concepts in the senses of Benson, Borwein and Hurwicz, we gain also scalarization results for the proper efficiency concept in the sense of Henig taking into account the known relationships between proper efficiency concepts in vector optimization.
+So, the assumptions of Corollary 33 are satisfied.
+∎
+Remark 21.
+Note that Corollaries 32, 33 and 34 provide necessary conditions for $\mathbb{A}$-proper efficiency (in the sense of Definition 7), including the proper efficiency concepts in the senses of Benson, Borwein and Hurwicz (as mentioned in Remark 5).
+Moreover, we gain scalarization results for the proper efficiency concept in the sense of Henig taking into account the known relationships between proper efficiency concepts in vector optimization.
 In order to emphasize this fact, consider a nontrivial cone $K\subseteq Y$ with normlike-base $B_{K}(\psi)$ and suppose that either $K$ is solid or $\psi$ is continuous.
-Assume that ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)\subseteq{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
-Taking into account Remark 5.27, in this subsection, we have provided sufficient conditions for the equivalence of the following assertions:
-  • $1^{\circ}$.
-    $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
-  • $2^{\circ}$.
-    $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
-  • $3^{\circ}$.
-    There is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12 is valid),
+Assume that ${\rm PEff}_{He}(\Omega\mid f,K)\subseteq{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$.
+Taking into account Remark 20, in this subsection, we have provided sufficient conditions for the equivalence of the following assertions:
+  • $1^{\circ}$
+    $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$.
+  • $2^{\circ}$
+    $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$.
+  • $3^{\circ}$
+    There is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 17 is valid, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing),
 where in a normed setting the third assertion is replaced by
   • $3^{\circ}$’.
-    There is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid).
-Remark 5.35.
-Under the assumptions in Corollary 5.28 a) (in particular the convexity of $A$), we were able to replace the separation condition (5.2) by $0\notin{\mathrm{c}l}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ (which implies the condition $A\cap(-K)=\{0\}$) in our scalarization results.
-In contrast, the scalarization result (for Henig and Benson proper efficiency concepts) by García-Castaño, Melguizo-Padial and Parzanese [16, Cor.
+    There is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 17, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid, if $\varphi$ is $-C_{||\cdot||^{\prime}}(x^{*},\alpha)$ representing).
+From this perspective, $3^{\circ}$ (respectively, $3^{\circ}$’) is a necessary and sufficient condition for the proper efficiency of $\bar{x}\in\Omega$.
+Remark 22.
+Under the assumptions in Corollary 32 a) (in particular the convexity of $A$), we were able to replace the separation condition (24) by $0\notin{\rm cl}\,S_{-K}$ and $\bar{x}\in{\rm PEff}_{\mathbb{A}}(\Omega\mid f,K)$ (which implies the condition $A\cap(-K)=\{0\}$) in our scalarization results.
+In contrast, the scalarization result (for Henig and Benson proper efficiency concepts) by García-Castaño, Melguizo-Padial and Parzanese [1, Cor.
 3.10] in real normed spaces still contains a separation condition (SSP), also under convexity of $A$ (ensured by a convexlikness assumption on $f$).
+It is worth mentioning that a novel aspect of our work in comparison to other conic scalarization related works in the normed space framework (e.g.
+[8, 9], [1]) is the treatment of equivalent norms in the scalarization results (as done in Corollaries 33 and 34).
 4.5 Scalarization results for the concept of proper efficiency in the sense of Henig
 Let us end the paper with stating some further scalarization results for the concept of proper efficiency in the sense of Henig by making special use of the structure of the dilating cones $D\in\mathcal{D}(K)$.
-Lemma 5.36.
+Lemma 35.
 Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, and $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$ for some $D\in\mathcal{D}(K)$.
 Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$.
-Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
-$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
+Suppose that ${\rm cl}\,S_{-K}$ is compact, and
+$({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
 Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
 $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D.$
-Proof 5.37.
+Proof.
 First, note that $\emptyset\neq K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ for all $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ taking into account the continuity of $\psi$.
-By Proposition 4.15, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
-$x^{*}(d)+\alpha\psi(d)\geq 0>x^{*}(-k)+\alpha\psi(k)\quad\mbox{for all }d\in\bar{D}\mbox{ and all }k\in K\setminus\{0\}.$ (5.3)
-Any $x\in C^{>}_{\psi}(x^{*},\alpha)$ satisfies $0>x^{*}(-x)+\alpha\psi(x)$, hence $-x\notin\bar{D}$ in view of (5.3), and so $x\in\operatorname{int}\,D$.
-Theorem 5.38.
+By Proposition 26, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
+$x^{*}(d)+\alpha\psi(d)\geq 0>x^{*}(-k)+\alpha\psi(k)\quad\mbox{for all }d\in\bar{D}\mbox{ and all }k\in K\setminus\{0\}.$ (25)
+Any $x\in C^{>}_{\psi}(x^{*},\alpha)$ satisfies $0>x^{*}(-x)+\alpha\psi(x)$, hence $-x\notin\bar{D}$ in view of (25), and so $x\in\operatorname{int}\,D$. ∎
+Theorem 36.
 Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone.
-Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)$.
+Let $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\rm Eff}(\Omega\mid f,D)$.
 Suppose that $D$ is closed.
@@
 Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$.
-Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
-$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
-Then, the conclusion of Theorem 5.25 holds.
-Proof 5.39.
-By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
+Suppose that ${\rm cl}\,S_{-K}$ is compact, and
+$({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
+Then, the conclusion of Theorem 31 holds.
+Proof.
+By Lemma 35 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
 $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D.$
 Since $D$ is closed, we have $C_{\psi}(x^{*},\alpha)\subseteq D$, hence $C_{\psi}(x^{*},\alpha)\setminus\{0\}\subseteq D\setminus\{0\}$. We conclude that
-$\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)\subseteq{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha)).$
-Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
-Theorem 5.40.
+$\bar{x}\in{\rm Eff}(\Omega\mid f,D)\subseteq{\rm Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha)).$
+Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 1 ($7^{\circ}$), Proposition 17 yields the rest of the conclusion. ∎
+Theorem 37.
 Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone.
-Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)$.
+Let $\bar{x}\in{\rm PEff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\rm WEff}(\Omega\mid f,D)$.
 Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$.
 Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$.
-Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
-$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
-Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
-Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
-Proof 5.41.
-By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
+Suppose that ${\rm cl}\,S_{-K}$ is compact, and
+$({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
+Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
+Moreover, if $\varphi$ is $-C_{\psi}(x^{*},\alpha)$ representing, any of the equivalent assertions given in Proposition 18 is valid.
+Proof.
+By Lemma 35 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
 $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D,$
-hence $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)\subseteq{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Proposition 5.14 finishes the proof.
-Remark 5.42.
-Proposition 4.1 gives sufficient conditions for the equality $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset$, which is given as an assumption in Theorems 5.38 and 5.40.
-Suppose that there is $(x^{*},\alpha)\in Y^{*}\times\mathbb{P}$ such that (4.4) is valid.
+hence $\bar{x}\in{\rm WEff}(\Omega\mid f,D)\subseteq{\rm WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Proposition 18 finishes the proof. ∎
+Remark 23.
+Taking Remark 12 into account, in Theorems 36 and 37, we derived necessary (and sufficient) conditions for the proper efficiency in the sense of Henig of $\bar{x}\in\Omega$.
+Proposition 19 gives sufficient conditions for the equality $({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset$, which is given as an assumption in Theorems 36 and 37.
+Suppose that there is $(x^{*},\alpha)\in Y^{*}\times\mathbb{P}$ such that (14) is valid.
 Notice that the set $\bar{D}$ is a nontrivial, closed cone in $Y$ with $\bar{D}\cap(-K)=\{0\}$.
-Assume that the nontrivial, convex cone $K$ is closed and that either ${\mathrm{c}onv}({\mathrm{c}l}(B_{\bar{D}})\cup\{0\})$ is closed or ${\mathrm{c}l}\,B_{\bar{D}}$ is compact.
-Then, Proposition 4.1 ensures $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
-Remark 5.43.
-The approach in this Section 5.4 follows basically some ideas by Kasimbeyli [45, Th.
+Assume that the nontrivial, convex cone $K$ is closed and that either ${\rm conv}({\rm cl}(B_{\bar{D}})\cup\{0\})$ is closed or ${\rm cl}\,B_{\bar{D}}$ is compact.
+Then, Proposition 19 ensures $({\rm cl}\,S_{\bar{D}}^{0})\cap({\rm cl}\,S_{-K})=\emptyset.$
+Remark 24.
+The approach in this Section 4.5 follows basically some ideas by Kasimbeyli [8, Th.
 4.4 and 5.7], who deals (in a normed setting) with $\varepsilon$-conic neighborhoods for $K$ (as dilating cones) and supposes a separation property for a whole family of dilating cones.
-In Section 5.3 one has a separation condition for $A=\mathbb{A}(\bar{x})$ and $-K$ (for a given $\bar{x}$), and one can further ensure that $A$ is convex.
+In Section 4.4, one has a separation condition for $A=\mathbb{A}(\bar{x})$ and $-K$ (for a given $\bar{x}$), and one can further ensure that $A$ is convex.
 Here in Section 4.5 we have a separation condition for $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ and $-K$.
@@
 In the general framework of real topological-linear spaces, our paper provides new insights into the relationships between a vector optimization problem based on an original (convex) cone, an updated vector optimization problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions.
-Thus, our new scalarization results with respect to the concepts of weak efficiency and different types of proper efficiency (e.g., in the sense of Benson; Borwein; Henig; Hurwicz) are also useful for several established scalarization methods in vector optimization (see also Remark 5.17).
+Thus, our new scalarization results with respect to the concepts of weak efficiency and different types of proper efficiency (e.g., in the sense of Benson; Borwein; Henig; Hurwicz) are useful for several established scalarization methods in vector optimization.
 In future research, we like to derive Bishop-Phelps type scalarization results for approximate solutions of vector optimization problems in real topological-linear spaces, which would extend recent scalarization results obtained by García-Castaño, Melguizo-Padial and Parzanese [1].